### v1

# Embodied Co-Design for Rapidly Evolving Agents: Taxonomy, Frontiers, and Challenges

Abstract
Brain-body co-evolution enables animals to develop complex behaviors in their environments. Inspired by this biological synergy, embodied co-design (ECD) has emerged as a transformative paradigm for creating intelligent agents—from virtual creatures to physical robots—by jointly optimizing their morphologies and controllers rather than treating control in isolation. This integrated approach facilitates richer environmental interactions and robust task performance. In this survey, we provide a systematic overview of recent advances in ECD. We first formalize the concept of ECD and position it within related fields. We then introduce a hierarchical taxonomy: a lower layer that breaks down agent design into three fundamental components—controlling brain, body morphology, and task environment—and an upper layer that integrates these components into four major ECD frameworks: bi-level, single-level, generative, and open-ended. This taxonomy allows us to synthesize insights from more than one hundred recent studies. We further review notable benchmarks, datasets, and applications in both simulated and real-world scenarios. Finally, we identify significant challenges and offer insights into promising future research directions. A project associated with this survey has been created at https://github.com/Yuxing-Wang-THU/SurveyBrainBody.
I Introduction
Embodied intelligence (EI) seeks to integrate cognitive capabilities into autonomous machines, enabling them to learn from and adapt to complex environments [howard2019evolving]. Situated at the intersection of artificial intelligence and robotics, EI has recently made significant advancements in fusing perception, decision-making, and actuation. This progress has empowered various robots to tackle previously intractable challenges, such as dynamic locomotion [lee2020learning, zhang2024whole], end-to-end navigation [zheng2024towards, schumann2024velma], and dexterous manipulation [chi2023diffusion, huang2023voxposer, wang2024dexcap], to name just a few.
In traditional EI, an agent’s morphology (body) is typically hand-designed and treated as a fixed part of the environment, while only its controller (brain) is optimized for task-oriented behaviors. However, this separation has clear limitations. Since an agent’s performance is intertwined with both its morphology and control, a pre-specified morphology can be suboptimal: it constrains sensing and actuation, thereby narrowing the space of achievable behaviors. In other words, when morphological design prioritizes mechanical optimality without considering how it will be controlled, the burden of achieving coordinated behavior shifts almost entirely to the controller. This imbalance often becomes a bottleneck for task performance and raises a fundamental question: How should we design the optimal form of an embodied agent to effectively perform tasks in its environment?
Figure 1: Differences between control design only and embodied co-design (ECD).
The concept of morphological computation [paul2006morphological, muller2017morphological, pfeifer2009morphological, pfeifer2006body, mertan2024no] posits that a well-designed body can offload or simplify computations typically assigned to the brain, thereby facilitating more efficient behaviors. This perspective introduces a brain-body dynamic: changes in morphology necessitate corresponding adaptations in control to exploit newly acquired capabilities, while environmental feedback can, in turn, guide further refinement of morphology. These synergies have motivated various efforts to improve the current design procedure for embodied agents. As a result, embodied co-design (ECD), an integrated design framework that co-optimizes an agent’s morphology with its control mechanism to enhance task performance, becomes a critical topic in EI.
ECD was originally studied in the Evolutionary Robotics (ER) community [harvey2005evolutionary, silva2016open], where both robotic behaviors and morphologies are concurrently optimized using Evolutionary Algorithms (EAs). However, this topic remains emergent for EI, particularly as agents and their tasks continue to grow in complexity and diversity. In contrast to controller-only optimization, ECD introduces a vast and often combinatorial design space for morphology-control configurations, accompanied by fragile interdependencies. Specifically, evaluating a candidate morphology typically requires re-training the controller from scratch, since even minor morphological changes can invalidate previously learned controllers. Moreover, variations in environmental conditions, design constraints, and task objectives produce heterogeneous feedback signals, further complicating the co-design process, as illustrated in Fig. 1. Consequently, ECD is highly non-convex and computationally expensive, highlighting the increasing necessity for automated ECD techniques over traditional manual engineering.
Figure 2: Co-designed embodied agents in both virtual and real environments. Example images numbered from (a) to (p) are adapted from [sims1994competition, hornby2001evolving, shim2003generating, Veenstra2023, ma2021diffaqua, ha2019reinforcement, luck2020data, zhao2020robogrammar, gupta2021embodied, xu2021accelerated, cheney2013unshackling, zardini2021seeking, medvet2021biodiversity, wang2023curriculum, wang2023softzoo, yue2025toward], respectively. Example images numbered from (q) to (y) are adapted from [nygaard2021environmental, revolve2, Strgar-RSS-24, xie2024map, schaff2022soft, liu_voxcraft_2020, kobayashi2024computational, matthews2023efficient, cheng2024structural], respectively.
Motivation: Recent breakthroughs in domains such as evolutionary algorithms [davis2023subtract, bergonti2024co, ferigo2025totipotent], reinforcement learning (RL) [yuan2022transformact, wang2023curriculum, dongleveraging2024, lu2025bodygen], differentiable simulation [xu2021accelerated, ma2021diffaqua, Strgar-RSS-24], and large models [gao2025vlmgineer, fang2025robomore, song2025laser] have led to a surge of novel ECD approaches, rapidly reshaping the capabilities of co-design systems. However, their assumptions, representations, computational pipelines, evaluation benchmarks, and limitations have not been systematically reviewed, which makes it difficult for researchers to develop a coherent understanding of the ECD landscape. A comprehensive survey is increasingly needed. Our goal is to consolidate these frontier developments, clarify conceptual foundations, highlight emerging trends, and identify open challenges that will shape the next generation of embodied intelligence research.
Taxonomy: To enhance readability for researchers across disciplines, we propose a two-layer hierarchical taxonomy in Section III. The lower layer structures ECD around three core dimensions: controlling brain (Section III-A), body morphology (Section III-B), and task environment (Section III-C). Building upon this foundation, the upper layer categorizes existing ECD methods into four major frameworks: bi-level (Section IV-A), single-level (Section IV-B), generative (Section IV-C), and open-ended (Section IV-D), with further subcategories identified within each. This taxonomy establishes a common vocabulary and provides a conceptual map for embodied co-design, facilitating cross-domain understanding. It also encourages researchers to explore beyond their primary areas. For instance, specialists in RL for control optimization may discover how their methodologies align with a specific ECD framework, in conjunction with previously unconsidered morphological evolution, thereby opening pathways for new interdisciplinary collaborations.
Scope: ECD bridges robotics, artificial intelligence, control theory, and cognitive science. Guided by our proposed taxonomy, we focus on analyzing the design spaces and representation methods of morphology, control, and task environments. More importantly, we investigate how these components are jointly coordinated within different co-design frameworks. Our aim is to provide an organizational perspective that clarifies the conceptual and technical interrelations among various ECD works, rather than exhaustively reviewing any particular algorithmic subfield (e.g., RL for control optimization) or engineering detail (e.g., 3D printing for morphological fabrication).
Related Works: Most prior surveys either provide broad overviews of embodied intelligence [long2025survey, duan2022survey, sun2024comprehensive, liu2025embodied] or focus on specific subtopics, such as embodied learning [liang2025large], manipulation [li2025survey], and navigation [xiong2025sensing]. These works primarily emphasize sensorimotor learning, while ECD has received limited attention. Partial reviews of ECD exist within specific domains, including soft robotics [zhao2024exploring, chen2020design, stroppa2024optimizing], humanoid robots [liu2025embracing, yue2025toward], evolutionary robotics [doncieux2014beyond, prabhu2018survey, alattas2019evolutionary, pandey2022accessible, strgar2024evolution], and artificial life [lai2021virtual, casadei2023artificial, ha2022collective]. However, they lack a structured analytical framework for systematically comparing co-design methodologies and do not incorporate recent advances, benchmarks, and emerging applications. In contrast, our survey provides a comprehensive and up-to-date analysis of ECD, covering a wide range of agent types and offering in-depth discussions on key challenges and future directions.
Contributions: To the best of our knowledge, this paper presents the first comprehensive survey of ECD. The main contributions are summarized below:
  1. 1.
    We conceptualize modeling approaches to ECD from three complementary perspectives: bi-level optimization, sequential decision-making, and model-based optimization (Section II). Together, this set of lenses provides a multifaceted view of the ECD landscape.
  2. 2.
    We contribute a hierarchical taxonomy (Section III) that organizes co-design spaces and identifies key design choices. The proposed taxonomy aims to deepen researchers’ understanding of the fundamental components involved in embodied co-design.
  3. 3.
    We present a systematic technical review (Section IV) of major advances in ECD over the past five years, analyzing representative methodologies and their core principles.
  4. 4.
    We survey simulation platforms and real-world testbeds commonly used in ECD research (Section V), highlighting their capabilities, representative demonstrations, and application domains.
  5. 5.
    We identify current bottlenecks and existing gaps in ECD, outlining promising research directions (Section VI) that may benefit a broad range of communities and help advance the state of the art.
The remainder of this survey is organized as follows. We begin with the background of ECD in Section II, which includes an introduction to its fundamental elements, mathematical definitions, and related topics. In Section III, we introduce our hierarchical taxonomy. In Section IV, we elaborate on four co-design frameworks based on the proposed taxonomy. We discuss notable benchmarks and contemporary applications of ECD in Section V. Finally, we outline the challenges of modern ECD and provide future perspectives in Section VI.
II Background: Formalizing Embodied Co-Design
II-A Problem Definition
II-A1 Embodied Agents
An embodied agent, such as a robot, can be characterized by its body morphology and controlling brain. The body morphology $\omega\in\Omega$ defines the agent’s physical interface with the outer surroundings, including its structural topology, material properties, and the spatial arrangement of sensors and actuators. The controlling brain $\pi\in\Pi$ maps observed states to action commands, driving the body to exert forces that interact with the task environment. These forces induce state transitions governed by the physical dynamics $F\in\mathcal{F}$, which, in turn, generate new sensory inputs. Through this sensorimotor loop, the agent learns to produce inherently goal-directed behaviors, with its performance evaluated using a fitness function $R\in\mathcal{R}$.
II-A2 Bi-Level Optimization View
ECD naturally has a bi-level structure, in which one optimization task is nested within another [sinha2017review, liu2021investigating]. The goal is to identify both the optimal morphology $\omega^{*}$ and the optimal control policy $\pi^{*}$, leading to the following formulation:
[…]
Here, the upper level searches for the optimal morphology, while the lower level trains a controller for the specified morphology to maximize the agent’s performance. Notably, this bi-level characteristic is also observable in other deep learning problems, such as Neural Architecture Search (NAS) [liu2021survey], where the objective is to jointly optimize both the network structure (analogous to morphology in ECD) and its parameters (analogous to control in ECD).
II-A3 Sequential Decision-Making View
While the bi-level formulation captures the hierarchical structure inherent in ECD, it fails to specify the causal connection between morphology and control. To bridge this gap, we formulate ECD as a single-level Markov Decision Process (MDP) [wang2022deep]. Following the sequential workflow commonly adopted in engineering practice, we divide the overall co-design process into two main stages: the design stage $\mathcal{D}$ and the control stage $\mathcal{C}$. As a result, the ECD problem can be represented as a $9$-tuple $\left\langle\mathcal{S_{D}},\mathcal{S_{C}},\mathcal{A_{D}},\mathcal{A_{C}},\mathcal{F_{D}},\mathcal{F_{C}},\mathcal{R_{D}},\mathcal{R_{C}},\gamma\right\rangle$, where $\gamma\in(0,1]$ denotes the reward discount factor, and the remaining elements correspond to the relevant design spaces. For clarity, explicit representations of constraints—such as morphological or task-specific constraints—are omitted and incorporated implicitly within their respective design spaces.
In this formulation, the overall state space comprises two components: the design state $s_{t}^{d}\in\mathcal{S_{D}}$, which governs morphology creation, and the control state $s_{t}^{c}\in\mathcal{S_{C}}$, which governs interactions with the environment. The corresponding action space includes design actions $a_{t}^{d}\in\mathcal{A_{D}}$ and control actions $a_{t}^{c}\in\mathcal{A_{C}}$. Under an episodic framework, a morphology is initially generated as $\omega=s_{T_{d}}^{d}$ by a design policy $\pi_{d}(a_{t}^{d}|s_{t}^{d})$, which executes a sequence of design actions guided by the design transition dynamics $F_{d}(s^{d}_{t+1}|s^{d}_{t},a^{d}_{t})$ from the initial design state $\omega_{0}=s^{d}_{0}$ to the final design step $T_{d}$. In the subsequent control stage, a policy $\pi_{c}(a_{t}^{c}|s_{t}^{c},\omega)$ conditioned on $\omega$ directs the generated morphology to perform tasks, following the environmental transition dynamics $F_{c}(s^{c}_{t+1}|s^{c}_{t},a^{c}_{t},\omega)$. Based on the above notations, the single-level objective $R$ of ECD is now expressed as:
$\begin{split}R=\mathbb{E}_{\pi_{d},\pi_{c}}\left[\sum_{t=0}^{T_{d}}\gamma^{t}r^{d}_{t}+\sum_{t=T_{d}+1}^{\infty}\gamma^{t}r^{c}_{t}\right]\end{split}$ (2)
where $r^{d}_{t}\in\mathcal{R_{D}}$ and $r^{c}_{t}\in\mathcal{R_{C}}$ denote the rewards obtained during the design and control stages, respectively. Leveraging effective gradient approximation techniques in RL [schulman2017proximal, silver2014deterministic, ciosek2020expected], both the design policy and the control policy can be updated iteratively through trial-and-error interactions.
II-A4 Model-Based Optimization View
It is noteworthy that, in the perspectives discussed above, the environmental transition model $F_{c}(s^{c}_{t+1}|s^{c}_{t},a^{c}_{t},\omega)$ is typically treated as a black box. However, as discussed in Section IV-B3, many ECD approaches construct $F_{c}$ as either a differentiable simulator [hu2019chainqueen, du2021diffpd, hu2019difftaichi] or a learned dynamics model [navez2025modeling]. Conceptually, the transition can be written as an abstract function $s^{c}_{t+1}=F_{c}(s^{c}_{t},a^{c}_{t},\omega)$, where $a^{c}_{t}$ denotes a vector of control signals, such as the joint torques applied during that time-step. Thus, given an initially parameterized morphology $\omega$ and a sequence of actions potentially derived from a control policy, the simulator unrolls a trajectory to a terminal control state $s^{c}_{T_{c}}$ and computes a task loss $\mathcal{L}$ that depends on this final state (e.g., distance traveled or final velocity [ma2021diffaqua, xu2021accelerated]). Since the simulator or learned model is differentiable with respect to environmental dynamics, morphological parameters, and control variables, gradients can propagate through the entire trajectory via the chain rule:
$\partialderivative{\mathcal{L}}{s^{c}_{t}}=\left(\partialderivative{\mathcal{L}}{s^{c}_{t+1}}\right)\left(\partialderivative{F_{c}}{s^{c}_{t}}\right)$ (3)
$\partialderivative{\mathcal{L}}{a^{c}_{t}}=\left(\partialderivative{\mathcal{L}}{s^{c}_{t+1}}\right)\left(\partialderivative{F_{c}}{a^{c}_{t}}\right)$ (4)
$\partialderivative{\mathcal{L}}{\omega}=\left(\partialderivative{\mathcal{L}}{s^{c}_{t+1}}\right)\left(\partialderivative{F_{c}}{\omega}\right)$ (5)
This model-based formulation facilitates end-to-end, gradient-based optimization for inverse design in ECD and provides substantial advantages in sample efficiency.
II-B Philosophical Foundations
Embodied Cognition has emerged from a sustained dialog between philosophical reflection and empirical science. Early philosophical discussions often treated cognition as an abstract, disembodied process largely independent of physical embodiment [descartes2016meditations]. This perspective was increasingly challenged in the mid-twentieth century, notably by Turing [turing2021computing], who argued that intelligence should arise through learning and adaptation, envisioning machines that grow and learn like children. Subsequent theoretical and empirical advances have further strengthened the critique of brain-centric perspectives. Clark [clark1998being] emphasized the centrality of sensorimotor processes, arguing that bodily interactions with the environment both constrain and enable cognitive capacities. Building on this foundation, Pfeifer and Scheier proposed that the body is not merely an instrument for perception and action, but rather an essential component of cognitive processing [pfeifer2006body, pfeifer2001understanding].
Morphological Computation recognizes that an agent’s physical body can inherently perform computations that simplify control and perception, rather than serving solely as a passive structure governed by a controller [paul2006morphological, muller2017morphological, pfeifer2009morphological, wilson2013embodied, koehl1996does]. A classic illustration is the passive dynamic walker [collins2005efficient], which achieves stable locomotion solely through its morphology, without any software control. However, such examples also reveal critical limitations: passive solutions typically support only a narrow set of behaviors and are effective only under constrained environmental conditions. This introduces a fundamental brain-body trade-off [xie2025morphology, hoffmann2017simple, rosendo2017trade, zhang2025co]: simple morphologies allow lightweight controllers for basic tasks, whereas more expressive morphologies often require sophisticated control mechanisms to fully exploit their capabilities in complex environments. A central goal in ECD, therefore, is to achieve an effective balance between morphological computation (as quantified in Section III-F) and controller-driven (informational) computation.
Evolutionary Theory suggests that brains and bodies co-evolve through adaptive processes driven by environmental pressures. Central to this theory is Natural Selection, originally articulated by Darwin [bonner1988evolution], which posits that traits that enhance an organism’s survival and reproduction are more likely to be passed on to future generations. Over time, this process shapes increasingly specialized and efficient morphologies and behaviors, improving the organism’s fit to its environment. Complementing this view, the Baldwin Effect links evolution and learning by suggesting that an organism’s ability to learn can accelerate evolutionary adaptation [jelisavcic2019lamarckian, luo2023enhancing, simpson1953baldwin]. Another corollary of evolution is that more complex environments stimulate the development of more capable brain–body systems [mahon2008critical, smith2005development, smith2005cognition]. Recent empirical findings from [gupta2021embodied] have confirmed these theories, showing that agents evolved in progressively more challenging terrains develop morphologies that enhance downstream learning efficiency.
Our survey will position various ECD methods within these foundational principles, illustrating how ECD integrates these insights to design agents that are both structurally sound and behaviorally adept.
II-C Related Topics
Evolutionary Robotics (ER) originally focused on evolving robot behaviors [bongard2013evolutionary, doncieux2015evolutionary, medvet2023evolutionary]. For example, behavioral trees have been evolved using genetic programming [scheper2016behavior]. Over time, the scope expanded to include the evolutionary design of both robotic hardware and control policies. A seminal contribution is Sims’ work [sims1994evolving], which evolves 3D virtual creatures by jointly optimizing their graph-based morphologies and neural controllers. Contemporary ECD builds on these insights but extends beyond traditional evolutionary robotics, incorporating a wider range of optimization and learning techniques to improve scalability, generalization, and applicability.
Artificial Life (A-Life) investigates the co-evolution of morphology and control, aiming to recreate life-like processes in virtual environments [moreno2005agency, aguilar2014past]. While A-Life shares the central concept of brain–body–environment coupling with ECD, it prioritizes the exploration of biological phenomena such as self-organization [gershenson2020self], adaptation [bedau2003artificial], and evolutionary dynamics [ito2016population, pilat2012evolution] in nature, rather than engineering practical embodied agents with specific functionalities.
Computational Robot Design leverages algorithmic tools to automate the synthesis of robots, typically following a sequential design process. By formulating complex constrained optimization problems, this approach employs techniques such as topology optimization [sigmund2000topology, sigmund2013topology] and CAD integration [tian2022assemble, willis2022joinable] to address structural feasibility, material selection, electronics integration, fabrication constraints, and dynamic modeling. Control, however, is often simplified or treated as a secondary component. In contrast, ECD is a broader concept that tightly couples morphology and control within a unified optimization loop, aiming not only to generate feasible body designs but also to produce adaptive, task-capable behaviors.
Robot Learning aims to equip robots with the ability to autonomously acquire perceptual, decision-making, and action skills, rather than relying solely on hand-engineered solutions [kroemer2021review]. It is closely connected to ECD, as learning-based control optimization serves as a central mechanism for co-design in tasks such as locomotion, manipulation, and navigation. Recent advances in robot learning have integrated model-based approaches [grandia2023perceptive], reinforcement learning [matsuo2022deep, singh2022reinforcement], large-model-based reasoning [xiao2025robot], imitation learning [hussein2017imitation, fang2019survey], and self-supervised methods [nair2020contextual, jeong2020self, kahn2021badgr] to enhance high-dimensional sensorimotor loops. However, most robot learning frameworks assume a fixed morphology. ECD generalizes this setting by treating morphology and control as interdependent variables and optimizing them jointly.
III Taxonomy: Decomposing Embodied Co-Design
Figure 3: An overview of the embodied co-design landscape including foundational elements and co-design frameworks.
ECD operates within an expansive workspace shaped by the interplay of control, morphology, and task environment. As shown in Fig. 3, the lower layer of our taxonomy organizes this space by complexity, while the upper layer defines four ECD frameworks, each with sub-categories that guide the co-design process. In the following sections, we detail each component, highlight recent advances, discuss common constraints, summarize key evaluation metrics, and illustrate a representative ECD pipeline using a prototype.
III-A Controlling Brain
The controlling brain enables an agent to process sensory information and generate appropriate motor responses for body control, thereby closing the sensorimotor loop.
III-A1 Design Space
The design space encompasses both the ways controllers are parameterized and the modes in which they operate.
Controller Parameters can be predefined—such as action primitives [iii2021taskagnostic], fixed waveform generators [cheney2013unshackling], and hand-crafted gaits [kobayashi2024computational]—or be fully learnable, optimized through methods such as trajectory optimization [spielberg2017functional, bergonti2024co], reinforcement learning, and imitation learning. Modern ECD approaches typically employ learnable controllers, which enable adaptation to varying morphologies and support flexible, real-time improvements in behavior.
Operating Mode defines the fundamental strategy by which a control system executes its commands and manages its actions. The two primary modes of control are open-loop and closed-loop. Open-loop controllers execute predefined action sequences without incorporating real-time sensory feedback; they are efficient and effective for simple, predictable tasks but cannot compensate for disturbances or failures during execution. In contrast, closed-loop controllers continuously incorporate sensory feedback to adapt their actions, offering greater robustness and generality at the cost of increased computational demands and more expressive policy representations. In embodied co-design, the closed-loop operating mode is typically favored.
III-A2 Representation
The controller representation defines the mapping between the control policy and the agent’s physical structure. Classical approaches adopt either individual or universal representations, with recent research exploring hybrid variants that combine their strengths.
Individual representations directly bind the controller architecture or dimensionality to a specific morphology. Examples include PID controllers with per-joint gains that account for the inertia and stiffness of specific body segments [ang2005pid, borase2021review], as well as Model Predictive Control (MPC) when an accurate dynamic model is available [schwenzer2021review, carron2019data]. Another example is Central Pattern Generators (CPGs), which mirror a robot’s limb topology to generate coordinated rhythmic motion [ijspeert2008central, bliss2012central, deshpande2023deepcpg]. Higher-level neural controllers, such as Multi-Layer Perceptrons (MLPs), adopt input/output sizes tied to a robot’s sensors and actuators.
Individualized controllers are appropriate for topology-fixed agents that share the same morphological structures but differ solely in their kinematic or dynamic parameters. This specialization, however, reduces transferability: even minor morphological changes—such as the addition of a joint or modifications to limb geometry—typically require controller restructuring or retraining. Consequently, per-agent representations struggle to generalize across heterogeneous state and action spaces, restricting scalability.
Universal representations aim to generalize across diverse morphologies. Representing a body as a kinematic graph naturally invites Graph Neural Networks (GNNs) [munikoti2023challenges]: their message-passing mechanisms enable a single policy to generalize across graphs with varying node counts, thereby supporting scalable control across different robotic embodiments. Representative GNN-based controllers include Nervernet [wang2018nervenet], DGN [pathak2019learning], Whitman et al. [whitman2023learning], SMP [huang2020one], MAT [li2024mat], MG [pan2025morphology], and GCNT [ijcai2025p972]. Subsequent research in cognitive science has further shown that modularity can enhance control generalization [schilling2024modularity, nadizar2022collective, pigozzi2022evolving, pigozzi2022evolving2, dong2022low, zhang2024composer, tjandrasuwita2024memo]. To capture the long-range dependencies that GNNs inherently struggle with, transformer-based controllers incorporate global attention over all body parts. This idea has been explored in numerous embodied AI systems, such as Amorpheus [kurin2020my], MetaMorph [gupta2022metamorph], SWAT [hong2021structure], AnyMorph [trabucco2022anymorph], Xiong et al. [xiong2023universal], Body Transformer [sferrazza2024body], HeteroMorpheus [hao2024heteromorpheus], and URMA [bohlinger2024one]. Beyond ECD, universal controllers have also shown strong utility in transfer learning [chen2018hardware, peng2018sim, cho2024meta], meta-learning [nagabandi2018learning, ghadirzadeh2021bayesian], and cross-embodiment learning [yang2024pushing, shafiee2024manyquadrupeds, wang2024sparse, chen2024mirage, patel2025get, mertan2024towards, feng2023genloco, luo2024moral]. This underscores their potential as a unifying representation for embodied AI.
Hybrid representations seek to balance specialization and generality by decomposing the controller or control mechanism into a shared, morphology-agnostic module that encodes universal motor and perceptual priors [bjelonic2023learning], and a morphology-specific module that adapts outputs to specific embodiments [chen2023evolving, luck2020data, luck2021robot, sliacka2023co, sun2023co]. Hybrid representations can be seamlessly integrated into various pipelines, such as pretraining-finetuning [yue2025toward, jin2025evolutionary] and meta-learning [belmonte2022meta], to facilitate cross-morphology generalization and faster adaptation during the ECD process.
III-A3 Discussion
Training a dedicated controller for each morphology yields an accurate estimate of that body’s potential by fully exploiting its passive dynamics and affordances. However, this approach is computationally expensive and biases the search toward morphologies that are easier to learn. Empirical results from [kvalsund2022centralized, mertan2023modular, chiappa2022dmap, luo2023comparison] have demonstrated that universal controllers can promote better synchronization, greater morphological diversity, and reduced premature convergence. However, they also require more comprehensive evaluation metrics to “diagnose” whether performance gains arise from enhanced controller capabilities or advantageous morphological adaptations.
These observations raise an important question: must one optimize a controller to evaluate morphology? Recent work in ECD suggests that the answer may be no. For instance, TAME [iii2021taskagnostic] introduces an information-theoretic, controller-free fitness function that measures a morphology’s intrinsic ability to explore its environment and produce predictable state transitions using only random action primitives.
III-B Body Morphology
Body morphology refers to the physical structure of an embodied agent. It is a set of properties that define the agent’s mechanical form, material composition, and physical interaction capabilities.
III-B1 Design Space
The morphological design space specifies the allowable parametric and structural variations.
Morphological Parameters are commonly categorized as discrete, continuous, and hybrid. Discrete parameters typically arise in modular systems where agents are composed of a finite set of building blocks, topologies, or design configurations. Continuous parameters govern real-valued attributes such as link lengths, shape, and material properties. More general embodied systems adopt a hybrid scheme in which certain choices, such as whether to attach a limb, are discrete, whereas attributes like limb length are continuous. This formulation balances structural flexibility with optimization tractability, enabling scalable search over high-dimensional morphologies.
Topological Structure further defines the scope of morphological exploration. In fixed topology settings, such as those involving bipedal [ha2019reinforcement], quadrupedal [belmonte2022meta], or manipulator skeletons [cai2023task2morph], optimization is constrained to parametric variations within a predefined template. This approach ensures stability and facilitates hardware deployment, but it restricts the discovery of novel morphologies. In contrast, unfixed topology settings allow for arbitrary structural modifications, ranging from voxel-based [bhatia2021evolution] to graph-based agents [yuan2022transformact] capable of growing, pruning, or reconfiguring their skeletons. While enabling open-ended exploration and emergent specialization, these approaches also raise concerns regarding physical plausibility, degeneracy, and manufacturability.
III-B2 Representation
Morphological representation refers to the encoding of an agent’s body. Three paradigms are commonly employed: direct, indirect, and hybrid.
Direct encodings explicitly specify the agent’s morphological attributes—such as link lengths, joint locations, voxel occupancy, and actuator strengths—using vectors [luck2020data], matrices [bhatia2021evolution, sun2023co], densities [sato2023topology, kobayashi2024computational, sato2025computational], or trees [gupta2021embodied, jelisavcic2019lamarckian, harada2024lamarckian]. Although highly interpretable and precise, direct schemes often result in high-dimensional search spaces that hinder scalability for complex morphologies.
Indirect encodings compress the agent’s morphological information into compact descriptors or generative rules, which are then expanded into complete morphologies through developmental or decision-making processes. Examples include L-Systems [veenstra2022effects], generative grammars [zhao2020robogrammar, zhao2022automatic, zharkov2024synergizing], Neural Cellular Automata (NCA) [veenstra2020different, wang2023curriculum], Compositional Pattern-Producing Networks (CPPNs) [cheney2013unshackling], and reinforcement learning-based decision-making [yuan2022transformact, wang2023curriculum]. Indirect encodings typically yield smoother fitness landscapes [thomson2024understanding], facilitate the escape from morphological local optima, and produce modular, symmetric morphologies with enhanced transferability.
Hybrid encodings aim to combine the strengths of both approaches by explicitly parameterizing part of the morphology while generating the remainder procedurally or via learned models [dongleveraging2024, ringel2025text2robot, saito2024effective, le2024improving, wang2023softzoo, mertan2025controller]. This design facilitates both global exploration and local refinement. However, it also introduces the challenge of balancing the search across heterogeneous spaces while maintaining coherence between explicit and generated components.
III-B3 Discussion
The trade-offs between direct and indirect morphological encodings remain an open question in ECD. Prior work [pigozzi2023morphology] has reported that direct encodings often yield faster learning and more consistent convergence. In contrast, indirect encodings tend to generate more diverse, spatially coherent, and globally varying morphologies, although the corresponding learning improvements can be less predictable. Comparative studies [komosinski2001comparison, veenstra2017evolution] further demonstrated that indirect encodings are advantageous during early exploration due to their compactness and genetic reuse, whereas direct encodings may outperform them in later optimization stages. These insights underscore the potential of employing hybrid morphology representation strategies that begin with indirect encodings for broad exploration, followed by a transition to direct encodings for fine-grained refinement.
III-C Task Environment
The task environment defines the external conditions under which an agent operates.
III-C1 Design Space
The design space of the task environment specifies whether environmental parameters remain fixed or vary during co-design, and how feedback signals are delivered.
Environmental Parameters are typically categorized into predefined and modifiable types. In predefined settings, all environmental attributes, such as terrain roughness and object placement, are fixed, resulting in stable and repeatable dynamics. Such static environments are widely utilized to isolate environmental factors, thereby enabling the targeted specialization of morphology and control. However, this specialization often results in brittle solutions that degrade under unmodeled perturbations or deployment beyond training conditions. In contrast, treating the environment as modifiable—for example, by sampling from a distribution of scenario parameters, introducing curriculum-based difficulty [wang2019poet], or employing non-stationary task sequences—encourages the emergence of agents that generalize more effectively [bossens2020qed, auerbach2014environmental, urbina2024brain].
Observability is a fundamental property of the task environment, as it determines the fidelity and completeness of the information an agent can perceive about the underlying system state. In a fully observable setting, the agent’s observations constitute a sufficient statistic of the true environment state. This condition simplifies learning and planning, as seen in classical control and various reinforcement learning tasks. Conversely, partial observability arises when agents are provided with incomplete [schaff2022soft], noisy [bjelonic2023learning], or local per-body/per-joint information [wang2023curriculum, schaff2022n, yuan2022transformact], which decouples immediate observations from the global state. Under such conditions, the agent must maintain an internal belief state or memory to integrate temporal observations. The degree of observability, therefore, shapes an agent’s perceptual and control mechanisms, ranging from simple reactive policies to sophisticated memory architectures or active perception strategies. It is also a critical axis for benchmarking an agent’s robustness and generalization in real-world scenarios.
III-C2 Representation
This defines the objectives, constraints, and evaluative (reward) signals that guide both the learning dynamics and the emergent morphology–control solutions in ECD. We broadly categorize these representations into three types: task-driven, task-agnostic, and hybrid.
Task-driven representations specify explicit performance goals—such as locomotion speed, energy efficiency, manipulation accuracy, and task completion—providing direct reward signals that align with predefined objectives, whether single or multiple. These representations support efficient optimization but may bias the search toward narrow solutions. In contrast, task-agnostic representations employ objective functions that do not encode a specific downstream task. Instead, they promote properties such as behavioral diversity, exploration, novelty, and morphological richness (see Section III-F). By decoupling environment representation from explicit task rewards, these approaches encourage broader coverage of the design space and can reveal unanticipated yet beneficial behaviors. Hybrid approaches integrate both paradigms, combining explicit task specifications with auxiliary task-agnostic signals. This joint formulation balances directed optimization with exploratory pressure and may serve as a more effective driving force for embodied co-design.
III-C3 Discussion
The task environment shapes, and is shaped by, the agent’s morphology and control. Prior studies have demonstrated that robots co-designed on uneven terrain tend to develop more stable, balanced, and limb-rich morphologies, whereas those created on flat terrain favor simpler forms [miras2020environmental, miras2019effects, spanellis2021environment, corucci2018evolving, ito2013alife]. Similarly, morphologies that evolved in complex, multi-objective environments transfer more effectively to novel tasks than those from simpler settings [gupta2021embodied]. Moreover, multi-task scenarios in ECD introduce additional brain–body trade-offs, where improvements on one task may degrade performance on others [wang2023preco]. As a result, multi-task co-design aims to cultivate broadly capable brain–body pairs by leveraging shared structures across tasks rather than optimizing each task independently.
III-D Co-Design Constraints
Constraints are intrinsic to the design space of embodied agents. Unconstrained ECD often leads to overfitting: the morphological search may yield designs that are unnecessarily complex for a single task or, conversely, result in trivial solutions with poor transferability. Therefore, principled regularization is essential for guiding the search toward balanced brain-body pairs. Morphological constraints, such as structural stability, material availability, joint range-of-motion, and actuation limits, need to be considered during optimization. Common strategies include enforcing symmetry priors [dong2023symmetry, gupta2021embodied] and incorporating penalties for manufacturing time, cost, and complexity into objective functions, as demonstrated in monotone co-design [censi2015mathematical, zardini2021co, wilhelm2023constraint]. Behavioral constraints, such as safety considerations, actuator saturation, and collision avoidance, can further limit the feasible solution space and hinder the exploitation of simulators. Despite their significance, many ECD approaches still operate in largely unconstrained settings. Integrating practical engineering and fabrication constraints remains a challenge for the deployment of real-world applications.
III-E Co-Design Frameworks
Based on the fundamental elements mentioned above, we further categorize ECD approaches into four frameworks: (1) bi-level co-design, (2) single-level co-design, (3) generative co-design, and (4) open-ended co-design. These categories reflect how algorithms optimize morphology, control, and, in some cases, the task environment. In Section IV, we elaborate on these frameworks and analyze the existing literature through the lens of our new taxonomy.
III-F Evaluation Metrics
In this section, we summarize representative metrics for evaluating ECD approaches, each of which quantifies different aspects of co-designed agent performance.
III-F1 Task Performance
Task performance remains the primary application-level metric, directly measuring goal achievement. Evaluating this aspect is expensive during morphology search, because it typically requires complete policy optimization and extensive rollout-based testing.
III-F2 Morphological Computation
Morphological computation (MC) quantifies the contribution of body–environment dynamics to state transitions, independent of control mechanisms. A common formalization uses conditional mutual information:
$\mathrm{MC}\;=\;I\big(s^{c}_{t+1};s^{c}_{t}\mid a^{c}_{t}\big)$ (6)
which measures how informative the previous control state $s^{c}_{t}$ is about the next state $s^{c}_{t+1}$ when the action $a^{c}_{t}$ is known. A large MC indicates that morphology and environment strongly drive the state transition. While recent studies have directly quantified and optimized MC in physical robots [chandiramani2024improving, ghazi2016evaluating], practical applications often rely on simplified or surrogate estimators [zahedi2013quantifying, ghazi2015quantifying, fuchslin2013morphological, muller2017morphological].
III-F3 Empowerment
Similar to MC, empowerment measures the maximum influence an agent can exert on future states through its actions [klyubin2005empowerment, de2018unified, leibfried2019unified, jung2011empowerment], and is widely used in reinforcement learning exploration [houthooft2016vime, kumar2018empowerment, choi2021variational]. Formally, the empowerment value for a control state $s^{c}_{t}$ is defined as the channel capacity between the action $a^{c}_{t}$ and the next state $s^{c}_{t+1}$:
$\mathcal{E}(s)=\max_{\pi}I(s^{c}_{t+1};a^{c}_{t}|s^{c}_{t})$ (7)
where $\pi$ is the empowerment maximizing policy. In ECD, high empowerment indicates that a morphology can effectively translate diverse control signals into predictable outcomes, making downstream control easier [iii2021taskagnostic].
III-F4 Quality-Diversity (QD) Score
As discussed in Section IV-B4, diversity-based ECD approaches focus on discovering a wide repertoire of high-performing morphologies and behaviors. The QD score quantifies coverage and performance within a discretized Morphological Descriptor (MD) space, where each cell represents a morphology niche. For an archive with $N$ cells, the QD score is computed as the sum of the task returns $\mathcal{R}_{i}$ of the best individual in each occupied cell:
$QDScore={\textstyle\sum_{i=1}^{N}}\mathcal{R}_{i}$ (8)
III-F5 Complexity
Complexity measures the cost of both the controller and the morphology. Controller complexity can be quantified using architecture-agnostic metrics such as FLOPs, which estimate computational cost, latency, and energy consumption [xie2025morphology]. Morphological complexity can be quantified through component counts (e.g., links, joints, voxels, and degrees of freedom) or structural descriptors (e.g., connectivity, actuator distribution, material heterogeneity, and symmetry).
III-F6 Robustness
Robustness evaluates an agent’s resilience to perturbations, including morphological damage (e.g., locked joints, sensor faults) and environmental variations. Robust agents maintain high performance and enable rapid adaptation, reflecting their inherent fault tolerance [powers2021good].
In summary, these metrics provide a multidimensional evaluation framework for ECD approaches, helping researchers enhance their understanding of how physical structures and controllers computationally contribute to behavior.
III-G Case Analysis
At the bottom of Fig. 4, we use BipedalWalker agent  [^1] https://gymnasium.farama.org/environments/box2d/bipedal_walker/ in the OpenAI Gymnasium platform as a working example to summarize the common pipeline of ECD.
  • Controlling Brain: The controller is a learnable multilayer perceptron (MLP) that receives joint angles, velocities, and ground-contact signals as inputs and outputs joint torques (individual representation), forming a standard closed-loop controller architecture.
  • Body Morphology: The agent has two legs, each composed of two segments, yielding a continuous, topology-fixed design space. Morphology is directly encoded by an eight-dimensional vector specifying the width and height of each leg segment. The sizes are constrained to values within 75% of the leg sizes in the original bipedal walker environment [wang2019poet, stensby2021co].
  • Task Environment: The agent operates within a predefined, fully observable terrain. The goal is to maximize the forward distance traveled, which is task-driven and characterized as a single-objective task reward function without constraints.
  • Co-Design Framework: A bi-level co-design framework is employed. Morphologies are randomly initialized, and the outer-loop morphological optimization is performed by a genetic algorithm, while the inner-loop controller optimization is carried out using reinforcement learning.
IV Methods: Synthesizing Embodied Co-Design
Figure 4: An overview of four embodied co-design frameworks. Example images of co-designed agents from (a) to (e) are adapted from DERL [gupta2021embodied], Softzoo [wang2023softzoo], EvolutionGym [bhatia2021evolution], RoboGrammar [zhao2020robogrammar], and [ha2019reinforcement].
In this section, we elaborate on four ECD frameworks and organize them into sub-categories following the taxonomy outlined in Section III. For each sub-category, we analyze its core methodological components and highlight recent approaches. An overview of these frameworks is presented in Fig. 4.
IV-A Bi-Level Co-Design
We start by introducing the bi-level co-design framework, which is the most widely adopted ECD approach. In bi-level co-design, nested optimization is employed: the inner loop optimizes the control policy for each morphology, while the outer loop explores the morphological space based on the performance evaluations of the morphologies, as illustrated in Fig. 4 (a). We organize recent work along this line into three subtopics: 1) Evolutionary Reinforcement Learning, 2) Surrogate-Assisted Methods, and 3) Other Methods.
IV-A1 Evolutionary Reinforcement Learning
In this method, the inner loop for control policy optimization typically employs reinforcement learning (RL) [singh2022reinforcement, wang2022deep], whereas the outer loop for morphology evolution relies on various evolutionary algorithms (EAs).
Two major classes of RL methods are used for policy optimization: model-based and model-free approaches. Model-free RL learns a policy directly from interactions, without explicitly modeling transition dynamics or reward functions. Representative algorithms include proximal policy optimization (PPO) [schulman2017proximal], trust region policy optimization (TRPO) [schulman2015trust], deep deterministic policy gradient (DDPG) [silver2014deterministic, lillicrap2015continuous], twin delayed DDPG [fujimoto2018addressing], and soft actor-critic (SAC) [haarnoja2018soft]. Model-based RL, in contrast, leverages a provided or learned dynamics model to refine the control policy. Representative frameworks include Mb-Mf [nagabandi2018neural], PETS [chua2018deep], GPS [levine2013guided], and PILCO [deisenroth2011pilco]. Model-based approaches are more sample-efficient but can struggle in complex domains where accurate dynamics modeling is harder than direct policy learning.
EAs are effective for exploring high-dimensional morphology spaces. In the outer loop, evolutionary operators determine how new morphologies are generated, thereby shaping both the breadth and depth of exploration [carvalho2024role]. In discrete design spaces, the mutation operator typically adds or removes structural elements—e.g., joints, limbs, modules, kinematic branches—or modifies material properties. In continuous design spaces—including link lengths, joint stiffness, and mass distribution—Gaussian perturbations are commonly used. The crossover operator further expands the search space by recombining structural components or genotype encodings from multiple parents. Notably, generative models, such as Large Language Models (LLMs), have been used as high-level mutation operators to propose structured, human-interpretable edits based on prompts [song2025laser], further expanding the operator toolbox for morphology search.
By alternating between individual-level controller learning and population-level morphology evolution, ERL methods mirror Darwinian evolution augmented with a Baldwinian learning mechanism [luo2022effects, li2023evaluation]. A representative example is Neural Graph Evolution (NGE) [wang2019neural], which evolves graph-structured, rigid-body morphologies while training universal GNN-based controllers via PPO. NGE demonstrates that GNN controllers can generalize across related body structures, thereby reducing the cost of controller retraining during morphology search. The scalability of ERL is further illustrated by Deep Evolutionary Reinforcement Learning (DERL) [gupta2021embodied]. In contrast to NGE, DERL employs individual controllers that are trained from scratch for each morphology. Coupled with increasingly challenging environments—ranging from flat-ground locomotion to uneven terrain and object manipulation—this framework enables the emergence of morphological intelligence, resulting in agents that exhibit increased robustness, adaptability, and behavioral diversity. Similar strategies appear in later works [park2021computational, bhatia2021evolution, chen2023evolving, wang2024learn, yue2025toward].
As summarized in Table I, recent research has also expanded the classical ERL workflow along several dimensions. First, richer morphological representations have been introduced, including hyperbolic embeddings [dongleveraging2024] and text-to-mesh generation [ringel2025text2robot]. Second, advanced control-learning techniques, such as meta-RL [belmonte2022meta], transfer learning [harada2024lamarckian], imitation learning [hernandezfollowing], and policy-reuse [xie2025accelerating], have improved sample efficiency and generalization. A parallel line of work introduces novel morphological selection mechanisms, such as successive halving [katayama2023ga], multi-fidelity archiving [zhao2024morphological, nagiredla2024ecode], and multi-objective balancing [zhao2025multi], which can prevent premature convergence in morphology space [doncieux2014beyond, mertan2024investigating].
Notably, the combination of EA and RL has also been widely explored in control optimization. These hybrid methods maintain a population of policies that evolve across generations, enabling broad exploration of the policy space while refining individual controllers via RL’s gradient-based or value-based updates. Representative approaches include ERL [khadka2018evolution], CERL [khadka2019collaborative], PDERL [bodnar2020proximal], SERL [wang2022surrogate], and ERL-Re2 [hao2023erl]. Comprehensive reviews of this line of work can be found in [drugan2019reinforcement, sigaud2023combining, bai2023evolutionary, zhu2023survey, lin2025evolutionary, li2024bridging].
IV-A2 Surrogate-Assisted Methods
A central challenge in bi-level co-design arises from its hierarchical structure, where the inner-loop optimization is computationally intensive and often requires days of training for complex tasks. Surrogate-assisted methods address this issue by employing predictive models to approximate task performance, thereby reducing the need for repeated full-scale policy evaluations.
Bayesian Optimization (BO) [shahriari2015taking] has emerged as a prominent framework for modeling the morphology–performance relationship in ECD. As a black-box, zeroth-order optimizer, BO efficiently handles expensive objective evaluations. At each iteration, BO updates a surrogate model using previously evaluated morphologies and their performance, then selects new candidates through acquisition functions such as Probability of Improvement (PI), Expected Improvement (EI), Entropy Search (ES), or Upper Confidence Bound (UCB), thereby balancing exploration and exploitation [wilson2018maximizing]. The selected candidates are evaluated, and their results are used to update the surrogate for the subsequent iteration.
Gaussian Processes (GPs) are among the most commonly used surrogates in BO. For instance, HPC-BBO [liao2019data] introduced a hierarchical framework that facilitates parallel hardware validation. Similarly, [bjelonic2023learning] pre-trained a universal controller in simulation and used BO to optimize parallel–elastic spring parameters for a real quadruped, demonstrating effective sim-to-real transfer. Variants of BO have been further explored in [bhatia2021evolution, hu2021neural, saito2024effective, pan2021emergent]. To enhance generalization, [chen2024pretraining] combined domain randomization with a pretraining–finetuning pipeline. Extending beyond single-task BO, [dong2024cagecoopt] proposed a multi-task BO framework for robot manipulators. More recently, [schneider2025task] combined BO with Hyperband to generate novel manipulator designs that improve performance across both seen and unseen tasks.
In addition to traditional surrogates used in BO, recent approaches have explored alternative choices. For example, Q-value functions (Q-Surro) from RL have been used to evaluate candidate morphologies [luck2020data, luck2021robot, du2025efficient]. Multi-dynamics models can further enhance generalization across different agent designs [sliacka2023co]. Alternatively, when historical interaction data is available, near-optimal control actions (A-Surro) can be partially or fully transferred from previously optimized designs to accelerate evaluation [liu2023rapidly]. More recent studies by [zhao2025cross, strgar2025accelerated] developed universal policies that span multiple tasks and can be efficiently used as a zero-shot performance evaluator (A-Surro), thereby improving efficiency. In [bai2025learning], manipulator morphologies are evaluated and optimized using pre-collected trajectories and a learned reward model (R-Model), with the reliance on reward model evaluations gradually increasing throughout the training process.
Table I: Summary of bi-level co-design approaches.
Method Year Agent Type Application (Sim/Real) Controlling Brain Body Morphology Task Environment
Design Space Representation Design Space Representation Design Space Representation
Evolutionary Reinforcement Learning
NGE [wang2019neural]
2019
[link]
Rigid
Robot
Simulation
{MuJoCo}
Learnable (CL) Universal Discrete (UF) Direct Predefined (Partial) Task-Driven
RL (PPO) GNN EA (Mutation) Graph Locomotion Task Reward
POEM [banarse2019body]
2019
[link]
Rigid
Robot
Simulation
{MuJoCo}
Learnable (CL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
RL (SVG) DNN EA (Mutation) Vector Sumo-Wrestling Task Reward
DERL [gupta2021embodied]
2021
[link]
Rigid
Robot
Simulation
{MuJoCo}
Learnable (CL) Individual Discrete (UF) Direct Modifiable (Full) Task-Driven
RL (PPO) DNN EA (Mutation) Tree Loco.&Manip. Task Reward
Park et al. [park2021computational]
2021
Wheeled
Robot
Simulation
{Unity}
Learnable (CL) Individual Discrete (UF) Direct Predefined (Full) Task-Driven
RL (PPO) DNN EA (GA) Tree Locomotion Task Reward
GA [bhatia2021evolution]
2021
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Individual Discrete (UF) Direct Predefined (Full) Task-Driven
RL (PPO) DNN EA (Mutation) Matrix Loco.&Manip. Task Reward
CPPN-NEAT [bhatia2021evolution]
2021
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Individual Discrete (UF) Indirect Predefined (Full) Task-Driven
RL (PPO) DNN EA (NEAT) CPPN Loco.&Manip. Task Reward
Belmonte et al. [belmonte2022meta]
2022
Quadrupedal
Robot
Simulation
{RaiSim}
Learnable (CL) Hybrid Continuous (F) Direct Modifiable (Full) Task-Driven
Meta-RL (PPO) DNN EA (CMA-ES) Vector Locomotion Task Reward
Chen et al. [chen2023evolving]
2023
[link]
Legged
Robot
Simulation
{Isaac Gym}
Learnable (CL) Hybrid Continuous (F) Direct Predefined (Full) Task-Driven
RL (PPO) DNN&Instinct EA (PSO) Vector Locomotion Task Reward
GASH [katayama2023ga]
2023
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Individual Discrete (UF) Direct Predefined (Full) Task-Driven
RL (PPO) DNN EA (GA) Matrix Locomotion Task Reward
HERD [dongleveraging2024]
2024
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Universal Discrete (UF) Hybrid Predefined (Partial) Task-Driven
RL (PPO) Transformer EA (CEM) Tree Loco.&Manip. Task Reward
SERL [cheng2024structural]
2024
[link]
Legged
Robot
Sim&Real
{Isaac Gym}
Learnable (CL) Universal Continuous (F) Direct Predefined (Full) Task-Driven
RL (PPO) DNN EA (GA) Vector Locomotion Task Reward
Li et al. [li2024generating]
2024
[link]
Endoskeletal
Robot
Simulation
{Customized}
Learnable (CL) Universal Discrete (UF) Indirect Predefined (Partial) Task-Driven
RL (PPO) Transformer EA (CMA-ES) Latent Locomotion Task Reward
MTMFEA [zhao2024morphological]
2024
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Individual Discrete (UF) Direct Predefined (Full) Task-Driven
RL (PPO) DNN EA (GA) Matrix Loco.&Manip. Task Reward
Harada et al. [harada2024lamarckian]
2024
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Individual Discrete (UF) Direct Predefined (Full) Task-Driven
RL (PPO) DNN EA (GA) Matrix Loco.&Manip. Task Reward
Wang et al. [wang2024evolution]
2024
Rigid
Robot
Simulation
{Isaac Gym}
Learnable (CL) Universal Continuous (F) Direct Predefined (Full) Task-Driven
RL (PPO) DNN EA (GA) Graph Locomotion Task Reward
ECoDe [nagiredla2024ecode]
2024
[link]
Rigid
Robot
Simulation
{MuJoCo}
Learnable (CL) Universal Continuous (F) Direct Predefined (Full) Task-Driven
RL (PPO) DNN HyperBand Vector Loco.&Balancing Task Reward
CoSIL [hernandezfollowing]
2024
[link]
Rigid
Robot
Simulation
{MuJoCo}
Learnable (CL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
RL (SAC) DNN EA (PSO) Vector Locomotion Task Reward
TLE [wang2024learn]
2024
Rigid
Robot
Simulation
{MuJoCo}
Learnable (CL) Universal Hybrid (UF) Direct Predefined (Full) Task-Driven
RL (PPO) GNN EA (Mutation) Graph Loco.&Manip. Task Reward
Text2robot [ringel2025text2robot]
2024
[link]
Quadrupedal
Robot
Sim&Real
{Isaac Gym}
Learnable (CL) Individual Hybrid (F) Hybrid Modifiable (Full) Task-Driven
RL (PPO) DNN EA (GA) Mesh+URDF Locomotion Task Reward
OS-MOEA [zhao2025multi]
2025
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Universal Discrete (UF) Direct Predefined (Full) Task-Driven
RL (PPO) DNN EA (NSGA-II) Matrix Locomotion Task Reward
Xie et al. [xie2025accelerating]
2025
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Individual Discrete (UF) Direct Predefined (Full) Task-Driven
RL (PPO) DNN EA (GA) Matrix Locomotion Task Reward
EA-CoRL [jin2025evolutionary]
2025
Humanoid
Robot
Simulation
{Isaac Gym}
Learnable (CL) Hybrid Continuous (F) Direct Predefined (Full) Task-Driven
RL (PPO) DNN EA (CMA-ES) Vector Chin-Up Task Reward
RoboCraft [yue2025toward]
2025
Humanoid
Robot
Simulation
{MuJoCo}
Learnable (CL) Hybrid Continuous (F) Direct Modifiable (Full) Task-Driven
RL (SAC) DNN EA Toolbox Vector Fall Recovery Task Reward
Surrogate-Assisted Methods
HPC-BBO [liao2019data]
2019
[link]
Micro
Robot
Simulation
{V-REP}
Learnable (OL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
Bayesian-Opt CPG Bayesian-Opt Vector Locomotion Task Reward
Co-Adapt [luck2020data]
2020
[link]
Legged
Robot
Simulation
{Bullet}
Learnable (CL) Hybrid Continuous (F) Direct Predefined (Full) Task-Driven
RL (SAC) DNN EA (Q-Surro) Vector Locomotion Task Reward
SG-MOrph [luck2021robot]
2021
Legged
Robot
Simulation
{Bullet}
Learnable (CL) Hybrid Hybrid (UF) Direct Predefined (Full) Task-Driven
RL (SAC) GNN&DNN EA (Q-Surro) Graph&Vector Locomotion Task Reward
NFW [hu2021neural]
2021
[link]
Legged
Robot
Simulation
{Bullet}
Learnable (CL) Individual Continuous (F) Direct Modifiable (Full) Task-Driven
RL (PPO) DNN Bayesian-Opt Vector Locomotion Task Reward
BO [bhatia2021evolution]
2021
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Individual Discrete (UF) Direct Predefined (Full) Task-Driven
RL (PPO) DNN Bayesian-Opt Matrix Loco.&Manip. Task Reward
CoIm [sliacka2023co]
2023
Legged
Robot
Simulation
{Bull.&MuJo.}
Learnable (CL) Hybrid Continuous (F) Direct Predefined (Full) Task-Driven
RL (SAC) DNN EA (Q-Surro) Vector Locomotion Task Reward
Bjelonic et al. [bjelonic2023learning]
2023
Quadrupedal
Robot
Sim&Real
{ANYmal}
Learnable (CL) Hybrid Continuous (F) Direct Modifiable (Partial) Task-Driven
RL (PPO) DNN Bayesian-Opt Vector Locomotion Task Reward
CoIL [rajani2023co]
2023
[link]
Humanoid
Robot
Simulation
{MuJoCo}
Learnable (CL) Universal Continuous (F) Direct Predefined (Full) Task-Agnostic
RL (SAC) DNN Bayesian-Opt Vector Motion Imitation Imitation Loss
AIEA [liu2023rapidly]
2024
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Individual Discrete (UF) Direct Predefined (Full) Task-Driven
RL (PPO) DNN EA (A-Surro) Matrix Loco.&Manip. Task Reward
Chen et al. [chen2024pretraining]
2024
Quadrupedal
Robot
Simulation
{Isaac Gym}
Learnable (CL) Hybrid Continuous (F) Direct Modifiable (Full) Task-Driven
RL (PPO) DNN Bayesian-Opt Vector Locomotion Task Reward
MTBO [dong2024cagecoopt]
2024
[link]
End Effector
Tools
Sim&Real
{Bullet}
Learnable (CL) Universal Continuous (F) Direct Predefined (Full) Task-Driven
RL (PPO) DNN Bayesian-Opt Vector Manipulation Task Reward
Saito et al. [saito2024effective]
2024
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Individual Discrete (UF) Hybrid Predefined (Full) Task-Driven
RL (PPO) DNN Bayesian-Opt Vector Loco.&Manip. Task Reward
MoMa [schneider2025task]
2024
[link]
Robot
Manipulator
Simulation
{Customized}
Learnable (CL) Individual Continuous (F) Direct Modifiable (Full) Task-Driven
RL (N2M2) DNN BO&HyperBand Vector Manipulation Task Reward
Table II: Continued from previous page.
Method Year Agent Type Application (Sim/Real) Controlling Brain Body Morphology Task Environment
Design Space Representation Design Space Representation Design Space Representation
Du et al. [du2025efficient]
2025
Humanoid
Robot
Sim&Real
{Isaac&Bull}
Learnable (CL) Universal Hybrid (F) Direct Modifiable (Full) Task-Driven
RL (PPO) DNN EA (Q-surro) Vector Locomotion Task Reward
CTKTEA [zhao2025cross]
2025
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Hybrid Discrete (UF) Direct Predefined (Partial) Task-Driven
RL (PPO) Transformer EA (A-Surro) Matrix Loco.&Manip. Task Reward
Strgar et. al. [strgar2025accelerated]
2025
[link]
Modular
Soft Robot
Simulation
{Taichi}
Learnable (CL) Universal Discrete (UF) Direct Modifiable (Full) Task-Driven
Grad. Opt. DNN&CPG EA (A-Surro) Matrix Locomotion Task Reward
Bai et. al. [bai2025learning]
2025
Soft
Manipulator
Sim&Real
{Nvidia Warp}
Predefined (CL) Universal Continuous (F) Direct Modifiable (Full) Task-Driven
None Trajectories EA (R-Model) Vector Manipulation Task Reward
Ha et al. [ha2018computational2]
2018
Rigid
Robot
Sim&Real
{DART}
Learnable (OL) Individual Discrete (UF) Direct Predefined (Full) Task-Driven
Numerical Solver Torques Heuristic Search Graph Loco.&Manip. Task Reward
Other Methods
Jelisavcic et al. [jelisavcic2019lamarckian]
2019
[link]
Modular
Robot
Simulation
{Gazebo}
Learnable (OL) Hybrid Discrete (UF) Direct Predefined (Full) Task-Driven
EA (HyperNEAT) CPPN&CPG EA (GA) Tree Locomotion Task Reward
TAME [iii2021taskagnostic]
2021
[link]
Rigid
Robot
Simulation
{MuJoCo}
Predefined (OL) Universal Hybrid (UF) Direct Modifiable (Full) Task-Agnostic
None Random EA (Mutation) Graph Loco.&Manip. Mutual Info.
MELAI [le2022morpho]
2022
Modular
Robot
Simulation
{CoppeliaSim}
Learnable (CL) Individual Hybrid (UF) Indirect Predefined (Full) Task-Driven
EA (ES) Recurrent NN NEAT CPPN Locomotion Task Reward
Fadini et. al. [fadini2022simulation]
2022
Rigid
Robot
Simulation
{Bullet}
Learnable (CL) Individual Continuous (F) Direct Modifiable (Full) Task-Driven
DDP LQR EA (CMA-ES) Vector Loco.&Manip. Task Reward
Sartore et.al. [sartore2023codesign]
2023
[link]
Humanoid
Robot
Simulation
{Gazebo}
Learnable (OL) Individual Hybrid (F) Direct Predefined (Full) Task-Driven
Nonlinear-Opt Torques EA (GA) Vector Payload Lifting Task Reward
A-MEL [le2024improving]
2024
Modular
Robot
Simulation
{Box2D}
Learnable (OL) Individual Discrete (UF) Indirect Predefined (Full) Task-Driven
EA CPPN&CPG EA (Mutation) CPPN Locomotion Task Reward
Strgar et al. [Strgar-RSS-24]
2024
[link]
Soft
Robot
Sim&Real
{Taichi}
Learnable (CL) Individual Discrete (UF) Direct Predefined (Full) Task-Driven
Grad. Opt. DNN&CPG EA (GA) Vector Locomotion Task Reward
MEHK [le2025efficient]
2025
Modular
Robot
Simulation
{Customized}
Learnable (CL) Individual Hybrid (UF) Indirect Predefined (Partial) Hybrid
Grad. Opt. Linear EA (AME) CPPN Locomotion Task+Hom.
  • Note: The [link] in each row directs to the corresponding code or project website. The shaded section indicates the category of each item, while the white section below lists the specific methods and representations used. CL: Closed-Loop, OL: Open-Loop, F: Fixed, UF: Unfixed, Loco.: Locomotion, Manip.: Manipulation, Bayesian-Opt: Bayesian Optimization, Grad. Opt.: Gradient-Based Optimization, Nonlinear-Opt: Nonlinear Optimization, DDP: Differential Dynamic Programming, LQR: Linear Quadratic Regulators, AME: Asynchronous Morpho-Evolution,Hom.: Homeokinesis
IV-A3 Other Methods
The inner-loop controller optimization can be implemented using alternatives to RL, including trajectory optimization [fadini2022simulation], evolutionary algorithms [jelisavcic2019lamarckian, le2022morpho, le2025efficient], numerical solvers [sartore2023codesign, ha2018computational2], and information-theoretic classifiers [iii2021taskagnostic]. To accelerate bi-level co-design, Goff et al. proposed A-MEL [goff2024efficient], which integrates asynchronous learning and evolution to reduce both time and computational costs. Strgar et al. [Strgar-RSS-24] further leveraged massively parallel differentiable simulations to efficiently optimize neural controllers for large populations of candidate morphologies.
IV-B Single-Level Co-Design
As shown in Fig. 4 (b), single-level co-design treats both morphology and control as variables within a unified optimization process, unlike bi-level co-design, which employs nested loops for these aspects. We organize recent work along this line into four subcategories: 1) EA-Based Methods, 2) RL-Based Methods, 3) Physics-Based Methods, and 4) Diversity-Based Methods, which are summarized in Table III.
IV-B1 EA-Based Methods
EAs have a long-standing history of jointly evolving robot morphology and control. In these methods, each individual in the population encodes a complete agent, encompassing both the body structure and the controller parameters. The evolutionary loop evaluates fitness and applies operators such as selection, mutation, and crossover to explore the combined design space.
Early work by Sims [sims1994evolving, sims1994competition] evolved 3D virtual creatures with embedded neural controllers, demonstrating the emergence of locomotion and showing that graph-based encodings naturally yield modular and symmetric morphologies. Subsequent research expanded this paradigm through diverse genotype representations, including developmental encodings [bongard2001repeated, brodbeck2015morphological], CPPNs [auerbach2011evolving, tanaka2022co], and learned morphological representations [zhao2025concurrent]. Recent efforts have introduced novel evolutionary search mechanisms. Morphological innovation protection [cheney2018scalable] mitigates premature convergence by preserving newly evolved traits, while other studies incorporate heuristic search [ha2018computational], biologically inspired strategies [davis2023subtract], and neuroscience-inspired approaches [ferigo2025totipotent]. Multi-objective EAs, such as NSGA-II [deb2002fast], have been utilized to balance competing design objectives and constraints. For example, minimizing energy consumption and mission time jointly in morphing aerial vehicles [bergonti2024co].
EA-based methods excel at discovering creative, unexpected morphologies due to their global and relatively unconstrained search. In soft robotics, they have been widely used to automatically generate complex geometries and behaviors [cheney2013unshackling, hiller2010evolving, cheney2014electro, hiller2012automatic, kriegman2021fractals]. However, scalability remains a major limitation: population-based evaluation is computationally expensive—often requiring thousands of simulations per generation—and convergence is typically slow, mirroring the efficiency bottlenecks observed in bi-level co-design.
IV-B2 RL-Based Methods
RL has proven effective for solving sequential decision-making problems, making it a natural tool for single-level embodied co-design. Modern RL-based approaches often consider control and morphology as jointly learnable parameters or approximate gradients for both.
Early methods incorporated morphology directly into the policy representation. For example, [ha2019reinforcement] used a population-based policy-gradient method (REINFORCE) to update control and morphology simultaneously. Similarly, [chen2018hardware] embedded body parameters into the computational graph of an RL policy, enabling joint optimization of physical parameters and neural network weights. [schaff2019jointly] co-optimized a differentiable distribution over morphology parameters alongside a universal controller; this framework was later extended by [schaff2022n] for more complex embodiments. Other works, such as [whitman2020modular2], employed Q-learning to guide modular manipulator design, while [luck2020data, luck2021robot] combined RL with a learned Q-value surrogate to predict the agent’s task performance.
A complementary paradigm treats design modifications themselves as actions within the RL framework. Under this formulation, the agent concurrently learns to control and perform morphological “design actions”. A representative work is Transform2Act [yuan2022transformact], which introduces a transform-and-control paradigm that formulates design optimization as a conditional policy over both control and morphology, with policy-gradient updates and experience sharing across designs for improved sample efficiency. Subsequent extensions incorporated curriculum learning [wang2023curriculum], multi-task pretraining [wang2023preco], symmetry priors [dong2023symmetry], hierarchical modeling [sun2023co], competition mechanisms [huang2024competevo], and temporal credit assignment [lu2025bodygen], with applications ranging from embodied manipulation [liu2023learning, kulz2025design] to freeform modular robot design [li2024reinforcement]. Moreover, real-world considerations have also motivated sim-to-real approaches. [schaff2022soft] showed that co-designed soft robots can effectively navigate complex terrains. MORPH [he2024morph] co-optimized a neural proxy of hardware alongside the control policy, enabling gradient-based updates even with non-differentiable physics.
Notably, another related direction investigates the real-time adaptation of morphology and control using RL. Pathak et al. [pathak2019learning] demonstrated that modular, self-assembling agents can learn to reconfigure their bodies online, with individual modules attaching or detaching to form composite entities that enhance adaptability across diverse tasks. Similar approaches have been applied to soft robots with dynamically deformable morphologies [huang2024dittogym, pigozzi2024pressure], rigid robots with reconfigurable structures [chiappa2022dmap], and block-based robotic assemblies [ghasemipour2022blocks].
IV-B3 Physics-Based Methods
Another approach to single-level co-design leverages physics-based optimization. If the simulator, analytic model, or learned dynamics is differentiable, gradients can be computed with respect to both the controller and body parameters. Task losses can then be backpropagated through time using standard optimizers (Grad. Opt.), such as ADAM [adam] or L-BFGS [zhu1997algorithm]. Compared with black-box EA search or gradient approximations in RL, these approaches exhibit greater sample efficiency.
Early efforts relied on parametric analytic models [spielberg2017functional, ha2018computational] and adjoint methods [desai2018interactive] to model the complex relationships between morphology and control parameters. Subsequent research introduced stochastic programming combined with trajectory optimization [bravo2020one, bravo2024engineering], improving scalability across tasks and environments. More recent methods incorporate hardware and environmental constraints [ghansah2023humanoid, vaish2024co, sartore2022optimization]. Most analytic approaches assume fixed body topologies, primarily due to modeling challenges. Differentiable simulators address this limitation by providing end-to-end gradients through various physical processes, such as soft-body [du2021diffpd, ma2021diffaqua, qiao2021differentiable, hu2019chainqueen, hu2019difftaichi], rigid-body [freeman2021brax, xu2021end, degrave2019differentiable], and differentiable rendering [murthy2020gradsim, ma2021risp]. These processes enable gradients to propagate through more expressive design spaces. Built on this foundation, topology-agnostic parameterizations, such as deformation-based [xu2021end, cai2023task2morph, liu2023learning, lirobottool2023] and density-based schemes [yuhn20234d, matthews2023efficient, sato2023topology, sato2025computational], have been proposed for agents with complex geometries. Additional mechanisms include particle sets [spielberg2021co, matthews2023efficient], Wasserstein barycenters [ma2021diffaqua, wang2023softzoo], and CPPN-based encodings [cochevelou2023differentiable].
Physics-based co-design can scale to very high-dimensional bodies (e.g., thousands of masses or “voxels”) because a single gradient pass provides information about all design parameters simultaneously. However, accurate differentiable models for contact, fluid dynamics, and soft-material deformation remain limited. These models often require smooth design spaces, along with simplified controllers (e.g., open-loop torque control), resulting in relatively simple behaviors. Combining model-based gradients with model-free RL is an emerging trend to overcome these limitations [spielberg2019learning].
Table III: Summary of Single-Level Co-Design approaches.
Method Year Agent Type Application (Sim/Real) Controlling Brain Body Morphology Task Environment
Design Space Representation Design Space Representation Design Space Representation
EA-Based Methods
Auerbach et al. [auerbach2011evolving]
2011
Modular
Robot
Simulation
{ODE}
Learnable (CL) Individual Continuous (UF) Indirect Predefined (Full) Task‑Driven
CPPN-NEAT RNN CPPN-NEAT CPPN Navigation Task Reward
Hiller et al. [hiller2012automatic]
2012
Soft
Robot
Sim&Real
{Customized}
Predefined (OL) Universal Continuous (UF) Direct Predefined (Full) Task-Driven
None CPG EA GMM Locomotion Task Reward
Cheney et al. [cheney2013unshackling]
2013
[link]
Modular
Soft Robot
Simulation
{VoxCAD}
Predefined (OL) Universal Discrete (UF) Indirect Predefined (Full) Task-Driven
None CPG CPPN-NEAT CPPN Locomotion Task Reward
Cheney et al. [cheney2014electro]
2014
Modular
Soft Robot
Simulation
{VoxCAD}
Learnable (OL) Individual Discrete (UF) Indirect Predefined (Partial) Task-Driven
CPPN-NEAT CPG CPPN-NEAT CPPN Locomotion Task Reward
Cheney et al. [cheney2018scalable]
2018
[link]
Modular
Soft Robot
Simulation
{VoxCAD}
Learnable (OL) Individual Discrete (UF) Indirect Predefined (Full) Task-Driven
EA (ES) CPG EA (ES) CPPN Locomotion Task Reward
Kriegman et al. [kriegman2021fractals]
2021
[link]
Modular
Soft Robot
Sim&Real
{Voxcraft}
Learnable (OL) Universal Discrete (UF) Indirect Predefined (Full) Task-Driven
EA (GA) CPG EA (GA) CPPN Locomotion Task Reward
SGR [tanaka2022co]
2022
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Individual Discrete (UF) Indirect Predefined (Full) Task-Driven
HyperNEAT ANN HyperNEAT CPPN Loco.&Manip. Task Reward
Sathuluri et al. [sathuluri2023robust]
2023
Robot
Manipulator
Sim&Real
{Drake}
Learnable (CL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
EA (CMA-ES) ID Controller EA (CMA-ES) Vector Manipulation Task Reward
Davis et. al. [davis2023subtract]
2023
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Individual Discrete (UF) Direct Predefined (Full) Task-Driven
EA (GA) DNN EA (GA) Matrix Locomotion Task Reward
Bergonti et. al. [bergonti2024co]
2024
[link]
Winged
Drone
Simulation
{CasADi}
Learnable (OL) Individual Hybrid (UF) Direct Modifiable (Full) Task-Driven
EA (NSGA-II) Torques EA (NSGA-II) Vector Flight Task Reward
Ferigo et al. [ferigo2025totipotent]
2025
[link]
Modular
Soft Robot
Simulation
{2D-VSR-Sim}
Learnable (CL) Individual Discrete (UF) Indirect Predefined (Partial) Task-Driven
EA (ES) ANN EA (ES) CPPN Locomotion Task Reward
STGCN-GA [zhao2025concurrent]
2025
Modular
Robot
Simulation
{Webots}
Learnable (CL) Individual Discrete (UF) Direct Modifiable (Partial) Hybrid
EA (GA) DNN EA (GA) Graph Locomotion Task+Adapt
RL-Based Methods
Ha et al. [ha2019reinforcement]
2018
[link]
Legged
Robot
Simulation
{Gym&Bullet}
Learnable (CL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
REINFORCE DNN REINFORCE Vector Locomotion Task Reward
Schaff et al. [schaff2019jointly]
2019
[link]
Legged
Robot
Simulation
{Bullet}
Learnable (CL) Universal Continuous (F) Direct Predefined (Full) Task-Driven
RL (PPO) DNN Policy Grad. Vector (GMM) Locomotion Task Reward
Pathak et al. [pathak2019learning]
2019
[link]
Modular
Robot
Simulation
{Unity}
Learnable (CL) Universal Discrete (UF) Direct Modifiable (Partial) Task-Driven
RL (PPO) GNN RL (PPO) Graph (GNN) Locomotion Task Reward
HWASP [Chen2020HardwareAP]
2020
[link]
Robot
Manipulator
Sim&Real
{MuJoCo}
Learnable (CL) Individual Continuous (F) Direct Modifiable (Full) Task-Driven
RL (PPO) DNN RL (PPO) Vector Manipulation Task Reward
Whitman et al. [whitman2020modular2]
2020
[link]
Robot
Manipulator
Simulation
{Customized}
Predefined (OL) Universal Discrete (F) Indirect Modifiable (Full) Task-Driven
IK Solver Torques RL (Q-Learning) Vector (DQN) Target Reaching Task Reward
Transform2Act [yuan2022transformact]
2022
[link]
Rigid
Robot
Simulation
{MuJoCo}
Learnable (CL) Hybrid Hybrid (UF) Indirect Predefined (Partial) Task-Driven
RL (PPO) GNN RL (PPO) Graph (GNN) Locomotion Task Reward
N-LIMB [schaff2022n]
2022
[link]
Legged
Robot
Simulation
{Isaac Gym}
Learnable (CL) Universal Discrete (UF) Indirect Modifiable (Partial) Task-Driven
RL (PPO) Transformer Policy Grad. Grammar (Trans.) Locomotion Task Reward
Schaff et al. [schaff2022soft]
2022
[link]
Soft
Robot
Sim&Real
{SOFA}
Learnable (CL) Universal Discrete (UF) Direct Predefined (Partial) Task-Driven
RL (SAC) DNN Dist. Adapt. Vector (Gibbs.) Locomotion Task Reward
CuCo [wang2023curriculum]
2023
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Universal Discrete (UF) Indirect Predefined (Partial) Task-Driven
RL (PPO) Transformer RL (PPO) Vector (NCA) Loco.&Manip. Task Reward
PreCo [wang2023preco]
2023
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Universal Discrete (UF) Indirect Predefined (Partial) Task-Driven
RL (PPO) Transformer RL (PPO) Vector (Trans.) Loco.&Manip. Task Reward
SARD [dong2023symmetry]
2023
[link]
Rigid
Robot
Simulation
{MuJoCo}
Learnable (CL) Universal Hybrid (UF) Indirect Modifiable (Partial) Task-Driven
RL (PPO) GNN RL (PPO) Graph (GNN) Loco.&Manip. Task Reward
Sun et al. [sun2023co]
2023
Modular
Robot
Simulation
{Webots}
Learnable (CL) Universal Discrete (UF) Indirect Modifiable (Partial) Task-Driven
RL (TD3) DNN RL (Q-Learning) Matrix (DQN) Locomotion Task Reward
Liu et al. [liu2023learning]
2023
[link]
Robotic
Tools
Sim&Real
{Bullet&Box2D}
Learnable (CL) Universal Continuous (UF) Direct Modifiable (Full) Task-Driven
RL (PPO) GNN RL (PPO) Vector (GNN) Manipulation Task Reward
MORPH [he2024morph]
2024
[link]
Robot
Manipulator
Simulation
{MuJoCo}
Learnable (CL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
RL (PG) DNN EA&RL Vector Manipulation Task Reward
CompetEvo [huang2024competevo]
2024
[link]
Legged
Robot
Simulation
{MuJoCo}
Learnable (CL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
RL (PPO) DNN RL (PPO) Vector (DNN) Competition Task Reward
Li et al. [li2024reinforcement]
2024
[link]
Modular
Soft Robot
Simulation
{Customized}
Predefined (OL) Individual Discrete (UF) Indirect Predefined (Full) Task-Driven
None CPG RL (PPO) Tensor (DNN) Locomotion Task Reward
BodyGen [lu2025bodygen]
2024
[link]
Legged
Robot
Simulation
{MuJoCo}
Learnable (CL) Universal Hybrid (UF) Indirect Modifiable (Partial) Task-Driven
RL (PPO) Transformer RL (PPO) Vector (Trans.) Locomotion Task Reward
Physics-Based Methods
Spielberg et al. [spielberg2017functional]
2017
Legged
Robot
Sim&Real
{Drake}
Learnable (OL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
Traj. Opt. Torques NLP Solver Vector Locomotion Task Loss
Desai et al. [desai2018interactive]
2018
[link]
Legged
Robot
Simulation
{Customized}
Learnable (OL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
Grad. opt. Torques Adjoint Method Vector Locomotion Task Loss
Spielberg et al. [spielberg2019learning]
2019
Soft
Robot
Simulation
{ChainQueen}
Learnable (CL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
Grad. Opt. DNN Grad. Opt. Vector Locomotion Task Loss
Bravo et al. [bravo2020one]
2020
Rigid
Robot
Simulation
{Customized}
Learnable (OL) Individual Continuous (F) Direct Modifiable (Full) Task-Driven
Stoc. Prog. Torques Stoc. Prog. Vector Loco.&Manip. Task Loss
Spielberg et al. [spielberg2021co]
2021
Soft
Robot
Simulation
{ChainQueen}
Learnable (CL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
Grad Opt. DNN Grad. Opt. Particle Sets Loco.&Sensing Task Loss
DiffHand [xu2021end]
2021
[link]
Robot
Manipulator
Sim&Real
{Redmax}
Learnable (OL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
Grad. Opt. Torques Grad. Opt. Vector Manipulation Task Loss
DiffAcqua [ma2021diffaqua]
2021
[link]
Soft
Swimmers
Simulation
{Taichi }
Learnable (CL&OL) Individual Continuous (F) Indirect Predefined (Full) Task-Driven
Grad. Opt. DNN&CPG Grad. Opt. Wass. Bary. Swimming Task Loss
Sartore et al. [sartore2022optimization]
2022
Humanoid
Robot
Simulation
{CasADi}
Learnable (OL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
Grad. Opt. Torques Grad. Opt. Vector Payload Lifting Task Loss
Table IV: Continued from previous page.
Method Year Agent Type Application (Sim/Real) Controlling Brain Body Morphology Task Environment
Design Space Representation Design Space Representation Design Space Representation
Yuhn et al. [yuhn20234d]
2023
[link]
Soft
Robot
Simulation
{Taichi}
Learnable (OL) Individual Continuous (UF) Direct Predefined (Full) Task-Driven
Grad. Opt. Pulse Signals Grad. Opt. Density Locomotion Task Loss
Matthews et al. [matthews2023efficient]
2023
[link]
Soft
Robot
Sim&Real
{Customized}
Learnable (OL) Individual Continuous (UF) Direct Predefined (Full) Task-Driven
Grad. Opt. CPG Grad. Opt. Particle Sets Loco.&Manip. Task Loss
SoftZoo [wang2023softzoo]
2023
[link]
Soft
Robot
Simulation
{Taichi}
Learnable (CL) Individual Continuous (F) Hybrid Modifiable (Full) Task-Driven
Grad. Opt. DNN&CPG Grad. Opt. Customized Locomotion Task Loss
Sato et al. [sato2023topology]
2023
Soft
Robot
Simulation
{Taichi}
Learnable (OL) Individual Continuous (UF) Direct Predefined (Full) Task-Driven
Grad. Opt. CPG Grad. Opt. Density Locomotion Task Loss
Task2Morph [cai2023task2morph]
2023
[link]
Robot
Manipulator
Simulation
{Redmax}
Learnable (OL) Individual Continuous (F) Direct Modifiable (Full) Task-Driven
Grad. Opt. Torques Grad. Opt. Vector Manipulation Task Loss
Cochevelou et al. [cochevelou2023differentiable]
2023
Soft
Robot
Simulation
{Taichi}
Learnable (OL) Individual Continuous (UF) Direct Predefined (Full) Task-Driven
Grad. Opt. CPG Grad. Opt. Tensor Locomotion Task Loss
Ghansah et al. [ghansah2023humanoid]
2023
Humanoid
Robot
Sim&Real
{Customized}
Learnable (CL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
Traj. Opt. Torques Traj. Opt. Vector Locomotion Task Loss
Li et al. [lirobottool2023]
2023
Robot
Tools
Sim&Real
{DiffHand}
Predefined (OL) Universal Continuous (F) Direct Modifiable (Full) Task-Driven
None Torques Grad. Opt. Vector Manipulation Task Loss
Kobayashi et al. [kobayashi2024computational]
2024
Soft
Robot
Sim&Real
{Taichi}
Predefined (OL) Universal Continuous (UF) Direct Predefined (Full) Task-Driven
None Pneumatic Signals Grad. Opt. Density Locomotion Task Loss
Bravo et al. [bravo2024engineering]
2024
Legged
Robot
Sim&Real
{CasADi}
Learnable (CL) Hybrid Continuous (F) Direct Modifiable (Full) Task-Driven
Grad. Opt. Torques Grad. Opt. Vector Locomotion Task Loss
Vaish et al. [vaish2024co]
2024
Robot
Manipulator
Simulation
{Customized}
Learnable (OL) Individual Continuous (F) Direct Modifiable (Full) Task-Driven
Grad. Opt. Torques Grad. Opt. Vector Manipulation Task Loss
Sato et. al. [sato2025computational]
2025
Soft
Robot
Simulation
{Taichi}
Learnable (CL) Individual Continuous (UF) Direct Modifiable (Full) Task-Driven
Grad. Opt. DNN Grad. Opt. Density Locomotion Task Loss
Diversity-Based Methods
Krvcah et al. [krvcah2010solving]
2010
Virtual
Creature
Simulation
{ODE}
Learnable (CL) Individual Hybrid (UF) Indirect Predefined (Partial) Hybrid
EA (NEAT) ANN EA (NEAT) Nested Graph Loco.&Navi. Task+Novelty
Lehman et al. [lehman2011evolving]
2011
Virtual
Creature
Simulation
{ODE}
Learnable (CL) Individual Hybrid (UF) Indirect Predefined (Partial) Hybrid
EA (NSGA‑II) ANN EA (NSGA‑II) Nested Graph Loco.&Navi. Task+Novelty
Nordmoen et al. [nordmoen2020quality]
2020
Modular
Robot
Simulation
{Bullet}
Learnable (OL) Individual Hybrid (UF) Direct Predefined (Full) Hybrid
EA (QD) CPG EA (QD) Graph Locomotion Task+Diversity
Howard et al. [howard2020diversity]
2020
Legged
Robot
Simulation
{Bullet}
Learnable (OL) Individual Hybrid (UF) Direct Predefined (Full) Hybrid
EA (QD) CPG EA (QD) Vector Locomotion Task+Diversity
Nordmoen et al. [nordmoen2021map]
2021
Modular
Robot
Simulation
{Bullet}
Learnable (OL) Individual Hybrid (UF) Direct Predefined (Full) Hybrid
EA (QD) CPG EA (QD) Graph Locomotion Task+Diversity
Zardini et al. [zardini2021seeking]
2020
[link]
Modular
Soft Robot
Simulation
{Bullet}
Learnable (CL) Individual Hybrid (UF) Direct Predefined (Full) Hybrid
EA (NEAT) DNN EA (QD) Vector Locomotion Task+Diversity
Norstein et al. [norstein2023effects]
2023
[link]
Modular
Robot
Simulation
{Unity}
Learnable (OL) Individual Hybrid (UF) Direct Predefined (Full) Task‑Driven
EA (QD) CPG EA (QD) Tree Locomotion Task Reward
Xie et al. [xie2024map]
2024
Soft
Gripper
Sim&Real
{COMSOL}
Predefined (OL) Universal Continuous (UF) Direct Predefined (Full) Hybrid
None None EA (QD) Vector Manipulation Task+Diversity
No-Brainer [mertan2024no]
2024
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Predefined (OL) Universal Discrete (UF) Indirect Predefined (Full) Hybrid
None Stimuli EA (QD) CPPN Locomotion Task+Diversity
Mertan et al. [mertan2025controller]
2025
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Hybrid Discrete (UF) Indirect Predefined (Full) Hybrid
EA (QD) DNN EA (QD) CPPN Loco.&Manip. Task+Diversity
Mkhatshwa et al. [mkhatshwa2025body]
2025
[link]
Swarm
Robot
Simulation
{RoboGen}
Learnable (CL) Individual Continuous (F) Direct Modifiable (Partial) Hybrid
EA (QD) DNN EA (QD) Vector Object Transport Task+Diversity
3B-QD [nadizar2025enhancing]
2025
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Hybrid Discrete (UF) Direct Predefined (Partial) Hybrid
EA (QD) DNN EA (QD) Matrix Loco.&Manip. Task+Diversity
  • Note: The [link] in each row directs to the corresponding code or project website. The shaded section indicates the category of each item, while the white section below lists the specific methods and representations used. For RL-based methods, we further specify their design policy representations. CL: Closed-Loop, OL: Open-Loop, F: Fixed, UF: Unfixed, Loco.: Locomotion, Manip.: Manipulation, ID: Inverse Dynamics, GMM: Gaussian Mixture Model, Policy Grad.: Policy Gradient, Dist. Adapt.: Distribution Adaptation, DQN: Deep Q Network, Grad. Opt.: Gradient-Based Optimization, Trans.: Transformer, Gibbs.: Gibbs Distribution, Wass. Bary.: Wasserstein Barycenter, Stoc. Prog.: Stochastic Programming, Traj. Opt.: Trajectory Optimization, NLP:Nonlinear Programming, ANN: Artificial Neural Network
IV-B4 Diversity-Based Methods
A persistent challenge in single-level co-design is the premature convergence of both morphology and controller. Diversity-based methods address this issue by optimizing a repertoire of distinct, high-quality body–brain pairs instead of focusing on a single global optimum. Early work [krvcah2010solving, lehman2011evolving] demonstrated this principle by employing Novelty Search [doncieux2019novelty] with local competition within a three-dimensional morphology space (height, mass, and joint count), showing that localized selection pressures help preserve morphological diversity.
The subsequent line built upon Quality-Diversity (QD) algorithms [pugh2016quality, cully2017quality]. Canonical QD methods target behavioral diversity [batra2024quality, nilsson2021policy, flageat2024enhancing, pierrot2022multi, faldor2023map, pierrot2022diversity]. For example, MAP-Elites [mouret2015illuminating] discretizes a user-defined behavior descriptor (BD) space into cells and stores the best-performing solution within each. In the context of ECD, morphology descriptors (MDs) are introduced to index MAP-Elites archives. For instance, [nordmoen2020quality, nordmoen2021map] discretized a two-dimensional MD space (total voxels and active voxels) and stored the fastest walker in each cell, achieving greater robustness and diversity compared to objective-only EAs. Recent works extend this framework to legged robots and manipulators [howard2020diversity, xie2024map, norstein2023effects], or distill the archive into a universal controller to stabilize the co-design process [mertan2025controller]. Moreover, dual-map approaches separately represent morphology and controller spaces [zardini2021seeking], whereas 3B-QD [nadizar2025enhancing] maintains three independent archives for morphology, control, and behavior, further enhancing the illumination of the agent co-design space.
Diversity-based methods accelerate convergence by avoiding repeated controller optimization for each new morphology. Remaining challenges include the automatic discovery of informative descriptors, scaling to high-resolution material-property spaces [pigozzi2023factors], and providing theoretical guaranties for niche-optimal solutions. Comparative studies of QD approaches in ECD are available in [mkhatshwa2023impact, mkhatshwa2025body].
IV-C Generative Co-Design
Generative co-design leverages automated generative models to capture the relationship among morphology, task, and performance (evaluated by learning a controller), thereby enabling the optimization of morphology within latent or rule-defined spaces, as illustrated in Fig. 4 (c). We further categorize the generative co-design framework into three topics: 1) Rule-Based Generation, 2) Latent-Based Generation, and 3) Large-Model-Based Generation. Each class employs a distinct generative representation and search strategy to explore morphological design spaces, as shown in Table V.
IV-C1 Rule-Based Generation
Early work along this line typically utilized graph-based generative encodings [sims1994competition, sims1994evolving, pilat2008creature] and L-systems [hornby2001evolution, hornby2001evolving, hornby2003generative] to jointly generate morphology and controller parameters, which were then evolved to create task-oriented agents. More recent approaches combine grammars with advanced search strategies. Robogrammar [zhao2020robogrammar] and its extension [zhao2025multi], for example, encode robots as graphs, enforce valid assemblies through a graph grammar, and couple heuristic search with a model-based controller to efficiently identify high-performing designs. Zhao et al. [zhao2022automatic] extended this framework to aerial vehicles by combining discrete grammar exploration with a differentiable simulator, enabling both rediscovery of canonical drones as well as the discovery of novel configurations. Zharkov et al. [zharkov2024synergizing] further applied grammar-based linkage graphs to co-design underactuated, tendon-driven grippers.
Rules effectively reduce the search space to plausible designs, explicitly encode prior knowledge such as symmetry and connectivity, and enhance manufacturability and modularity. While they are flexible and interpretable, these rules depend on expert-crafted grammars, which can introduce bias into the search process and limit the exploration of novel or fine-grained morphological variations.
IV-C2 Latent-Based Generation
Latent-based methods leverage neural generative models to learn smooth mappings from low-dimensional latent vectors to full 3D or graph-based morphologies. These learned latent spaces support efficient gradient-based, evolutionary, or probabilistic search over complex design spaces and are increasingly influential in ECD.
Variational Autoencoders (VAEs) learn a continuous latent representation of morphology and, in some cases, control parameters, facilitating sample-efficient optimization. LABO [pan2021emergent] embeds high-dimensional hand morphologies and controller parameters into a VAE latent space and performs BO within that space, thereby reducing dimensionality and sample complexity for grasping tasks. GLSO [hu2023glso] extends this idea to graph-structured robots by training a graph-VAE on grammar-generated topologies and optimizing designs using BO. Liu et al. [liu2025terrain] proposed a Grammar-VAE that maps discrete morphologies into a continuous latent space and integrated differential evolution with a neural decoder, allowing the generative model to be reused across tasks. MorphVAE [song2024morphvae] further conditions latent sampling on task embeddings and introduces a “continuous natural selection” strategy to improve diversity and performance.
Generative Adversarial Networks (GANs) train a generator–discriminator pair to produce realistic morphologies by matching target data distributions. RoboGAN [hu2022modular] learns a one-to-many mapping from task descriptions (e.g., terrain type) to distributions over modular legged robots. Candidate bodies are sampled from the GAN manifold, while an evolutionary loop guides the search toward high-performing configurations, thereby enabling interpolation and diversity within the learned latent space.
Diffusion Models have recently emerged as powerful generative priors for embodied co-design. DiffuseBot [wang2023diffusebot] integrates differentiable physics into the denoising process, using simulation feedback at each step to bias generation toward high-performing designs. Dynamics-Guided Diffusion Models (DGDM) [xu2024dynamics] replace task-specific training data with physics-based guidance, enabling generalization to new manipulation tasks and novel geometries.
GFlowNets provide an alternative generative formulation in which morphology–controller pairs are sampled with probability proportional to their simulated performance. Recent works by [nagiredla2024robonet, nagiredlate, nagiredlacogent] demonstrated that GFlowNets can produce diverse, high-quality designs while learning a structured generative policy over complete embodied agents.
Latent generative models effectively capture complex, free-form geometries, exploit correlations among design variables, and typically offer significant gains in sample efficiency when compared to direct search methods. Once trained, a latent model can be reused across various tasks or paired with different controllers. However, training generative models typically requires large datasets and simulation efforts. Moreover, latent spaces may decode to invalid or unrealistic morphologies without appropriate constraints. Achieving the right balance between fidelity and diversity remains challenging.
IV-C3 Large-Model-Based Generation
Large pretrained models—including Large Language Models (LLMs) [naveed2025comprehensive] and Vision–Language Models (VLMs) [zhang2024vision]—have recently been introduced into ECD as generative priors or intelligent search operators. These models leverage broad knowledge and powerful generative capabilities to propose, mutate, or refine robot morphologies and control strategies.
A growing line of work uses LLMs as intelligent mutation operators or search heuristics. OpenELM [lehman2023evolution] demonstrates that a LLM trained on code can act as a high-quality mutation operator in genetic programming. In Sodarace experiments, the LLM repeatedly mutated code that generates robot walkers, producing thousands of novel and functional agents without prior robotics data; the resulting dataset was then used to train a conditional generative model that maps terrain descriptions to robot code. Building on this paradigm, LASeR [song2025laser] integrates a LLM into the optimization loop via a reflection mechanism that summarizes past search histories and prompts the LLM to generate improved design variants. RoboMoRe [fang2025robomore] further investigates LLM-driven co-design by asking the LLM to jointly propose candidate morphology–reward pairs during a coarse search, after which promising candidates undergo alternating refinement. RoboMorph [qiu2024robomorph] represents each robot with a structured grammar and relies on LLM reasoning to navigate vast combinatorial spaces, iteratively improving morphologies through feedback.
LLMs have also been utilized for human–AI collaborative design. Stella et al. [stella2023can] have shown that a LLM can provide high-level conceptual guidance and partial specifications for a tomato-harvesting gripper, illustrating its potential in early-stage design. Beyond text-based reasoning, vision models are beginning to appear in ECD. VLM-Engineer [gao2025vlmgineer] combines VLMs with evolutionary search to jointly design robotic tools and associated control policies. This line of work suggests future workflows in which a VLM first generates sketches or 2D concept designs from a task description, which are then lifted into full 3D morphologies.
Large models offer several advantages for ECD: they encapsulate extensive engineering and commonsense knowledge, generate creative or unconventional designs, support semantic prompting (e.g., “design a robot capable of climbing a ladder”), and diversify the search space. Their in-context learning abilities make them powerful assistants for bridging high-level human intent with concrete design proposals [makatura2023can, ma2024llm]. However, outputs from LLMs and VLMs are not guaranteed to be feasible, optimal, or physically consistent. Their behavior is often sensitive to prompt phrasing, and unguided LLM-driven searches can be highly sample-inefficient. Vision-based 3D generation is still in its infancy, and the reliance on proprietary foundation models raises additional concerns regarding reproducibility, transparency, and computational costs.
Table V: Summary of Generative Co-Design approaches.
Method Year Agent Type Application (Sim/Real) Controlling Brain Body Morphology Task Environment
Design Space Representation Design Space Representation Design Space Representation
Rule-Based
RoboGrammar [zhao2020robogrammar]
2020
[link]
Rigid
Robot
Simulation
{Bullet}
Learnable (CL) Individual Discrete (UF) Indirect Predefined (Full) Task‑Driven
MPPI MPC GHS Grammar Locomotion Task Reward
RoboGrammar [xu2021multi]
2021
[link]
Rigid
Robot
Simulation
{Bullet}
Learnable (CL) Individual Discrete (UF) Indirect Predefined (Full) Task‑Driven
MPPI MPC GHS Grammar Locomotion Task Reward
Zhao et al. [zhao2022automatic]
2022
Aerial
Vehicle
Simulation
{Customized}
Learnable (CL) Individual Hybrid (UF) Hybrid Predefined (Full) Task‑Driven
Grad. Opt. LQR Grad. Opt.& GHS Grammar&Vector Flight Task Reward
Zharkov et al. [zharkov2024synergizing]
2024
[link]
Robot
Manipulator
Sim&Real
{Rostok}
Predefined (OL) Individual Discrete (UF) Indirect Predefined (Full) Task‑Driven
None Torques MCTS Grammar Manipulation Task Reward
Latent-Based
Pan et al. [pan2021emergent]
2021
[link]
Robot
Manipulator
Simulation
{Bullet}
Learnable (OL) Individual Hybrid (UF) Indirect Modifiable (Full) Hybrid
Bayesian-Opt Torques Bayesian-Opt Latent (VAE) Manipulation Task+Complexity
RoboGAN [hu2022modular]
2022
Modular
Robot
Simulation
{Bullet}
Predefined (CL) Universal Discrete (F) Indirect Modifiable (Full) Task‑Driven
RL DNN EA (ES) Latent (GAN) Locomotion Task Reward
GLSO [hu2023glso]
2023
[link]
Modular
Robot
Simulation
{Bullet}
Learnable (CL) Individual Discrete (UF) Indirect Predefined (Full) Task‑Driven
MPPI MPC Bayesian-Opt Graph (VAE) Locomotion Task Reward
Diffusebot [wang2023diffusebot]
2023
[link]
Soft
Robot
Sim&Real
{SoftZoo}
Learnable(OL) Individual Continuous (UF) Indirect Predefined (Full) Task-Driven
Grad. Opt. CPG Grad. Opt. Latent (Diffusion) Loco.&Manip. Task Reward
MorphVAE [song2024morphvae]
2024
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Individual Discrete (UF) Indirect Predefined (Full) Task‑Driven
RL (PPO) DNN Selection Latent (VAE) Loco.&Manip. Task Reward
RoboNet [nagiredla2024robonet]
2024
Modular
Robot
Simulation
{MuJoCo}
Learnable (CL) Individual Discrete (UF) Indirect Predefined (Full) Task-Driven
RL DNN Grad. Opt. Graph (GFlowNet) Locomotion Task Reward
DGDM [xu2024dynamics]
2024
[link]
Robot
Gripper
Sim&Real
{MuJoCo}
Predefined (OL) Universal Continuous (UF) Indirect Modifiable (Full) Task-Driven
None Torques Grad. Opt. Latent (Diffusion) Manipulation Task Reward
TERoboNet [nagiredlate]
2025
Modular
Robot
Simulation
{MuJoCo}
Learnable (CL) Hybrid Discrete (UF) Indirect Predefined (Full) Task-Driven
RL (PPO) DNN Grad. Opt. Graph (GFlowNet) Locomotion Task Reward
COGNet [nagiredlacogent]
2025
Modular
Robot
Simulation
{MuJo&EvoG}
Learnable (CL) Individual Discrete (UF) Indirect Predefined (Full) Task-Driven
RL (PPO) DNN Grad. Opt. Graph (GFlowNet) Locomotion Task Reward
TAMS [liu2025terrain]
2025
[link]
Modular
Robot
Simulation
{MuJoCo}
Learnable (CL) Individual Discrete (UF) Indirect Modifiable (Full) Task-Driven
MPPI MPC EA (DE) Latent (VAE) Locomotion Task Reward
Large-Model-Based
OpenELM [lehman2023evolution]
2025
[link]
Sodarace
Robot
Simulation
{Sodarace}
Learnable (OL) Individual Hybrid (UF) Indirect Modifiable (Full) Hybrid
LLM with GP CPG LLM with GP Python Program Locomotion Task+Diversity
LASeR [song2025laser]
2025
[link]
Modular
Soft Robot
Simulation
{EvoGym}
Learnable (CL) Individual Discrete (UF) Direct Predefined (Full) Hybrid
RL (PPO) DNN EA (LLM) Matrix Loco.&Manip. Task+Diversity
RoboMoRe [fang2025robomore]
2025
[link]
Rigid
Robot
Simulation
{MuJoCo}
Learnable (CL) Individual Continuous (F) Direct Modifiable (Full) Task-Driven
RL (SAC) DNN LLM XML Locomotion LLM-Generated
RoboMorph [qiu2024robomorph]
2025
[link]
Modular
Robot
Simulation
{MuJoCo}
Learnable (CL) Individual Discrete (UF) Indirect Predefined (Full) Task-Driven
RL (SAC) DNN EA (LLM) Grammar Locomotion Task Reward
VLMGINEER [gao2025vlmgineer]
2025
[link]
End Effector
Tools
Simulation
{Bullet}
Learnable (OL) Individual Continuous (UF) Direct Predefined (Full) Task-Driven
EA (VLM) Way Points EA (VLM) URDF Manipulation Task Reward
  • Note: The [link] in each row directs to the corresponding code or project website. The shaded section indicates the category of each item, while the white section below lists the specific methods and representations used. For Latent-based methods, we further specify their generative models. CL: Closed-Loop, OL: Open-Loop, F: Fixed, UF: Unfixed, Loco.: Locomotion, Grad. Opt.: Gradient-Based Optimization, Manip.: Manipulation, MCTS: Monte‑Carlo Tree Search, GHS: Graph Heuristic Search, MPPI: Model Predictive Path Integral, DE: Differential Evolution, GP: Genetic Programming
IV-D Open-Ended Co-Design
Open-ended co-design aims for the continual or unbounded improvement of agents by enabling their bodies, controllers, and environments to co-evolve. These systems promote continuous innovation by enabling agents and environments to mutually generate new challenges and adaptations, as shown in Fig. 4 (d). We categorize open-ended co-design into two primary subtopics: 1) Brain–Body–Environment Co-Evolution and 2) Developmental Evolution, as summarized in Table VI.
IV-D1 Brain-Body-Environment Co-Evolution
This approach treats the task environment as a co-optimized component alongside morphology and control. By evolving environments (or task curricula) together with agents, it creates continuously shifting challenges that prevent premature convergence and promote both behavioral and morphological improvement. The environment thus functions simultaneously as a generator of new selective pressures and as a filter for successful morphology–controller adaptations.
Foundational work in this direction is the Paired Open-Ended Trailblazer (POET) framework [wang2019poet, wang2020enhanced]. POET maintains a population of environment–controller pairs, simultaneously mutating environments and evolving agents. Policies that succeed in one environment may be transferred to other contexts. Stensby et al. [stensby2021co] extended POET to jointly evolve morphologies and controllers in tandem with their environments. Their results indicate that learned environment progressions yield greater morphological diversity than handcrafted curricula and reduce premature convergence [eguiarte2024premature].
Recent work has introduced RL-driven co-evolution of the brain, body, and environment. Ao et al. proposed MECE [ao2023curriculum], which formulates a curriculum in which both morphology and the environment evolve according to learned RL policies. This produces progressively challenging and structured tasks that drive sustained adaptation. Another emerging trend augments environment evolution with generative models. For example, LLM-POET [aki2024llm] replaces manual environment mutation with a LLM that generates increasingly complex environments for a population of soft-robot agents. As agents improve, the LLM proposes novel tasks that escalate difficulty, realizing an open-ended loop of continual challenge and innovation.
IV-D2 Developmental Evolution
Developmental co-evolution draws inspiration from evolutionary developmental biology (Evo-Devo) and developmental robotics, where genotypes encode growth processes that shape both morphology and control over an agent’s lifetime (ontogeny) [asada2001cognitive, bongard2003evolving, lungarella2003developmental, kriegman2018morphological, van2021comparing, ivanchenko2003developmental]. Open-endedness emerges from developmental plasticity [higashinaka2024evolution]: agents are evaluated across multiple developmental stages or allowed to modify their bodies during their lifetime, while evolution optimizes the developmental rules themselves.
Early research has explored developmental encodings, including L-systems [rieffel2014growing], grammar rules, and neural networks [pontes2022single, joachimczak2016artificial]. Bongard [bongard2003evolving] demonstrated that artificial ontogeny can generate hierarchical, repeated structures and controllers that emerge as bodies grow, yielding more structured solutions than direct genotype-phenotype encodings. The other research line focuses on morphological development. Here, developmental rules adjust parameters such as leg angles or actuator amplitudes during learning [bongard2011morphological, bongard2010utility, naya2020experiment], or progressively “unfreeze” degrees of freedom to facilitate skill acquisition [lungarella2002adaptivity, berthouze2004motor, savastano2012incremental]. Later work introduced mechanisms for growing new body parts during an agent’s lifetime [arita2016alife, benureau2022morphological, jaafar2025eco], providing additional adaptive flexibility. For real-world deployment, virtual growth is typically halted once the simulated agent reaches the physical dimensions of the hardware robot, after which the learned controller is transferred to the physical platform [vujovic2017evolutionary, brodbeck2015morphological]. For comprehensive overviews of developmental evolution, we refer readers to [naya2021morphological, wang2021survey]. Overall, open-ended co-design is not yet a unified methodology but rather a collection of emerging ideas that leverage ontogeny, plasticity, and staged learning, and it awaits further exploration.
Table VI: Summary of Open-Ended Co-Design approaches.
Method Year Agent Type Application (Sim/Real) Controlling Brain Body Morphology Task Environment
Design Space Representation Design Space Representation Design Space Representation
Co-Evo
Stensby et al. [stensby2021co]
2021
[link]
Legged
Robot
Simulation
{Gym}
Learnable (CL) Individual Continuous (F) Direct Modifiable (Full) Hybrid
EA (GA) DNN EA (GA) Vector Locomotion Task+Diversity
MECE [ao2023curriculum]
2023
Rigid
Robot
Simulation
{MuJoCo}
Learnable (CL) Universal Hybrid (UF) Indirect Modifiable (Full) Task-Driven
RL (PPO) GNN RL (PPO) Graph (GNN) Locomotion Task Reward
Developmental Methods
Rieffel et al. [rieffel2014growing]
2014
Soft
Robot
Simulation
{PhysX}
Learnable (OL) Individual Hybrid (UF) Indirect Predefined (Full) Task-Driven
EA (GA) CPG EA (GA) L-System Locomotion Task Reward
Brodbeck et al. [brodbeck2015morphological]
2015
Modular
Robot
Real-World
Learnable (OL) Individual Discrete (UF) Direct Predefined (Full) Task-Driven
EA (GA) CPG EA (GA) Vector Locomotion Task Reward
Joachimczak et al. [joachimczak2016artificial]
2016
Virtual
Creature
Simulation
{Customized}
Learnable (OL) Universal Hybrid (UF) Indirect Predefined (Full) Task-Driven
EA (NEAT) ANN&CPG EA (NEAT) ANN Locomotion Task Reward
Vujovic et al. [vujovic2017evolutionary]
2017
Legged
Robot
Real-World
Learnable (OL) Individual Hybrid (F) Direct Predefined (Full) Task-Driven
EA PID EA Vector Locomotion Task Reward
Naya et al. [naya2020experiment]
2020
Quadrupedal
Robot
Simulation
{V-REP}
Learnable (OL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
EA (NEAT) ANN Development Vector Locomotion Task Reward
Benureau et al. [benureau2022morphological]
2022
Virtual
Creature
Simulation
{Customized}
Learnable (OL) Individual Continuous (F) Direct Predefined (Full) Task-Driven
EA (ES) CPG Development Vector Locomotion Task Reward
NCRS [pontes2022single]
2022
[link]
Modular
Robot
Simulation
{Box2d}
Learnable (CL) Individual Discrete (UF) Indirect Predefined (Partial) Task-Driven
EA (CMA-ES) NCA EA (CMA-ES) NCA Navigation Task Reward
Higashinaka et al. [higashinaka2024evolution]
2024
Modular
Soft Robot
Simulation
{EvoGym}
Predefined (OL) Individual Discrete (UF) Direct Predefined (Full) Task-Driven
None CPG EA (GA) Matrix Locomotion Task Reward
Jaafar et al. [jaafar2025eco]
2025
Virtual
Creature
Simulation
{Bullet}
Learnable (CL) Individual Hybrid (UF) Indirect Modifiable (Partial) Task-Driven
HyperNEAT CPPN&DNN HyperNEAT CPPN Locomotion Task Reward
  • Note: The [link] in each row directs to the corresponding code or project website. The shaded section indicates the category of each item, while the white section below lists the specific methods and representations used. CL: Closed-Loop, OL: Open-Loop, F: Fixed, UF: Unfixed, Loco.: Locomotion, Manip.: Manipulation, PID: Proportional Integral Derivative, NCA: Neural Cellular Automata
V Testbeds: Applying Embodied Co-Design
Embodied co-design has evolved rapidly due to the availability of high-fidelity simulated platforms and increasingly advanced real-world fabrication techniques. These resources provide standardized environments, reproducible tasks, and practical feedback loops that facilitate the automated design and validation of embodied agents [long2025survey]. This section surveys representative benchmarks, datasets, and applications of ECD across both simulated and real-world domains.
V-A Simulated Platforms
V-A1 Rigid-Body Simulators
Widely used robot simulators, including MuJoCo [todorov2012mujoco], Bullet [panerati2021learning], Unity [juliani2018unity, Veenstra2023], Gazebo [koenig2004design], Webots [michel2004cyberbotics], SAPIEN [xiang2020sapien], DART [lee2018dart], Rostok [borisov2022reconfigurable], and V-Rep [rohmer2013v], provide native support for standard robot description formats such as XML, MJCF, URDF, and SDF. This compatibility enables the rapid modification of morphologies and establishes these simulators as the foundation for numerous works in ECD, as summarized in Tables IVI. The convergence of robotics, reinforcement learning, and photorealistic rendering has further increased the demand for scalable, GPU-accelerated simulation. Platforms such as NVIDIA Isaac Gym and its extensions [makoviychuk2021isaac, kulkarni2023aerial, gu2024humanoid] enable massively parallel environments that support the simultaneous training of large sets of morphological variants with controllers [cheng2024structural, wang2024evolution, chen2023evolving, ringel2025text2robot, jin2025evolutionary, chen2024pretraining, du2025efficient, schaff2022n], significantly accelerating ECD pipelines.
While much of the existing ECD literature relies on non-differentiable simulators that only provide forward dynamics, a growing number of differentiable platforms supporting inverse dynamics have emerged. These tools—Dojo [howell2022dojo], Brax [freeman2021brax], Neuralsim [heiden2021neuralsim], Nimble [werling2021fast] (the differentiable version of DART), DiffRedMax [xu2021accelerated], and SDRS [ye2024sdrs]—facilitate gradient-based co-design. Based on these efforts, frameworks such as Genesis [Genesis] and MuJoCo Playground [mujoco_playground_2025] aim to unify differentiable physics, GPU acceleration, and integrated rendering within a single API, enabling end-to-end pipelines for the co-design of complex embodied agents with neural network–based controllers, as shown in Table VII.
V-A2 Soft-Body Simulators
For soft-bodied agents, early platforms such as SOFA [coevoet2017software, deimel2017automated], VoxCAD [cheney2013unshackling, cheney2014electro, cheney2018scalable], and VoxCraft [kriegman2021fractals, li2024reinforcement] remain influential baselines due to their detailed physical models of deformable materials. Lightweight two-dimensional frameworks, such as 2D-VSR-Sim [ferigo2025totipotent], further enable rapid algorithmic prototyping. Recently, EvolutionGym [bhatia2021evolution] introduces a comprehensive suite of tasks along with a diverse material palette for voxel-based soft robots, which has since been widely adopted in ECD research [katayama2023ga, dongleveraging2024, harada2024lamarckian, zhao2024morphological, wang2023curriculum, wang2023preco, xie2025accelerating, zhao2025multi, liu2023rapidly, saito2024effective, zhao2025cross, davis2023subtract, tanaka2022co, mertan2024no, mertan2025controller, nadizar2025enhancing, song2025laser, song2024morphvae, liu2025morphology]. Complementary to voxel-based systems, several platforms integrate differentiable soft-body physics, including DiffTaichi [hu2019difftaichi], ChainQueen [hu2019chainqueen], and DiffAqua [ma2021diffaqua]. SoftZoo [wang2023softzoo] extends this direction by providing a differentiable benchmark across various terrains (e.g., flat ground, snow, and water) and task objectives (e.g., speed, turning, and path following). In parallel, Sorotoki [caasenbrood2024sorotoki] provides an integrated MATLAB toolkit for co-designing soft robots.
V-A3 Embodied Tasks Simulators
Notably, recent specialized simulators for embodied intelligence, such as iGibson [li2022igibson], AI2-THOR [kolve2017ai2], Habitat-Sim [szot2021habitat], and LIBERO [liu2023libero], provide realistic and interactive 3D environments that integrate rich visual, physical, and semantic information. These platforms support the training and evaluation of increasingly complex embodied agents, including those with advanced control paradigms (e.g., Vision-Language-Action models) [zitkovich2023rt, kim2024openvla, bjorck2025gr00t] and various morphologies (e.g., humanoids). Furthermore, they provide challenging benchmarks for long-horizon tasks, which makes them well-suited foundations for advancing ECD research in the future.
V-B Datasets
In addition to the aforementioned simulation platforms, several recent ECD datasets have been developed to support the systematic analysis of co-design across diverse embodied agents. RoboCrafter-QA [chen2025large] provides comparative data for evaluating whether LLMs can link high-level task descriptions to low-level morphological and material choices in soft robotics. RoboDesign1M [le2025robodesign1m] offers a large-scale dataset of one million robotic design samples across multiple domains, enabling tasks such as design–image generation, design-centric visual question answering, and design–image retrieval.
Table VII: A Summary of the Past Decade’s Simulators and Physics Engines Supporting Embodied Co-Design.
Platform Year Agent Differentiable Link
SOFA 2015 Soft No [link]
VoxCAD 2015 Soft No [link]
ChainQueen 2018 Soft Yes [link]
DART 2018 Rigid No [link]
Webots 2018 Mixed No [link]
DiffTaichi 2019 Mixed Yes [link]
Revolve 2019 Soft No [link]
Habitat-Sim 2019 Mixed No [link]
SAPIEN 2020 Rigid No [link]
2D-VSR-Sim 2020 Soft No [link]
TDS 2020 Rigid Yes [link]
Brax 2021 Rigid Yes [link]
MuJoCo 2021 Mixed No [link]
PyBullet 2021 Mixed No [link]
Isaac Gym 2021 Mixed No [link]
Evolution Gym 2021 Soft No [link]
Nimble Physics 2021 Rigid Yes [link]
DiffHand 2021 Rigid Yes [link]
DiffAqua 2021 Soft Yes [link]
AI2-THOR 2021 Mixed No [link]
iGibson 2021 Mixed No [link]
Modular EvoGym 2022 Soft No [link]
2D-MR-Sim 2022 Soft No [link]
2D-Robot-Evolution 2022 Soft No [link]
Rostok 2022 Mixed No [link]
Dojo 2022 Rigid Yes [link]
EMR 2023 Modular No [link]
SoftZoo 2023 Soft Yes [link]
LIBERO 2023 Rigid No [link]
RoboGen 2024 Mixed No [link]
Revolve2 2024 Mixed No [link]
Genesis 2024 Mixed Yes [link]
Sorotoki 2024 Soft No [link]
SDRS 2024 Soft Yes [link]
V-C Real-World Applications
While simulation facilitates rapid exploration, physical validation remains the ultimate test of ECD. Recent advances in fabrication, multi-material 3D printing, and embedded sensing have made it feasible to realize co-designed morphologies. In this section, we summarize recent real-world applications.
V-C1 Rigid Robots
These efforts primarily focus on developing agents with enhanced stability, agility, and energy efficiency. Many systems adopt bi-level or single-level co-design frameworks to jointly optimize morphology and control. Recent real-world demonstrations span bipedal [cheng2024structural], quadrupedal [bjelonic2023learning, bravo2024engineering, nygaard2021environmental, bravo2024engineering, ringel2025text2robot, ha2018computational2, ha2017joint], micro-scale [spielberg2017functional], and wheeled robots [park2021computational], resulting in highly adaptable agents that demonstrate robust performance.
V-C2 Soft Robots
Researchers have also investigated the embodied co-design of soft robots made from highly deformable materials [caasenbrood2024sorotoki]. A pioneering example is the shape-changing crawling robot introduced in [schaff2022soft], which demonstrates how deformable morphologies can be co-optimized with control for adaptive locomotion. In parallel, the VoxCraft platform [kriegman2021fractals] enables the rapid physical realization of voxel-based soft robots through modular assembly and disassembly, thereby supporting large-scale embodied evolution in hardware-rich environments. Among these applications, we also observe emerging trends in leveraging differentiable simulation alongside advanced actuation methods to narrow the co-design gap between simulation and real-world deployment [matthews2023efficient, bacher2021design].
V-C3 Modular Robot
Open frameworks, such as Revolve [revolve2, lan2021learning] and RoboGen [auerbach2014robogen, jelisavcic2017real], provide C++ and Python libraries that facilitate the creation, simulation, and management of robots within the Gazebo platform. These frameworks also enable virtual designs to be 3D-printed and subsequently deployed in the real world. More recently, the Granulobot system [saintyves2024self] demonstrates how gear-like modular units can self-assemble into aggregates with tunable rigidity, offering new insights into ECD. Similarly, EMARGE [moreno2021emerge] introduces a platform for full-cycle morphological evolution in reconfigurable hardware. We refer readers to [liang2025decoding] for a comprehensive review along this line.
V-C4 Aerial and Aquatic Vehicles
For aerial agents, ECD targets aerodynamic efficiency, control efficiency, and payload trade-offs, which are typically validated through flight testing or wind-tunnel experiments [bergonti2024co, xu2019learning]. Aquatic agents—both underwater and surface—require joint optimization of geometry, propulsion, and control to reduce drag and improve maneuverability and sensing in fluid environments, with evaluations conducted in pools or through field deployments [katzschmann2018exploration, sariman2025ur].
V-C5 Surgical Robots
ECD is also gaining increasing attention in surgical robotics and wearable device design, where co-optimization must simultaneously satisfy strict safety, sterility, and human-compatibility requirements [berger2022design, alcaraz2024using, alcaraz2025designing].
V-C6 Humanoids
Embodied co-design for humanoids focuses on anthropomorphic morphology, joint ranges, balance, and perception–action integration, all under strict safety and ergonomic constraints [ghansah2023humanoid, vanteddu2024cad, liu2025embracing].
V-C7 Manipulators and End-Effectors
Because full-body real-world ECD remains highly challenging, much practical progress has been made in co-designing manipulators and specialized end-effectors [chen2021co]. Approaches include generative methods based on diffusion models [wang2023diffusebot, xu2024dynamics], Grammar-based search [zharkov2024synergizing], bi-level co-design with surrogate models [bai2025learning, islam2024task, dong2024cagecoopt], and single-level co-design leveraging learned dynamics models [yi2025co]. These methods yield manipulators and tools with improved efficiency, robustness, and generalization in real-world manipulation tasks.
VI Challenges and Future Trends
In this section, we identify some open challenges and future directions in ECD, based on both retrospectives of the methods discussed in this survey and outlooks to the emerging trends of embodied intelligence.
VI-A Open Challenges
VI-A1 Efficiency and Scalability
The co-design space of embodied agents is prohibitively vast. As discussed in Section IV-A, a wide range of ECD methods typically require expensive inner-loop policy optimization processes to support the morphological search. This computational burden makes the development of efficient and accurate performance predictors essential for scaling ECD to complex, high-dimensional scenarios. In addition, although techniques such as curriculum learning [wang2023curriculum, ao2023curriculum], representation learning [spielberg2019learning, song2024morphvae], and transfer learning [zhao2024morphological, harada2024lamarckian, zhao2025cross, chen2018hardware] have been introduced to promote co-design knowledge reuse in ECD, scalability remains a critical challenge. Co-design approaches need to generalize across agents of varying sizes and structural complexities, rather than being limited to narrow task domains or specific morphological scales.
VI-A2 Generalization
Most ECD techniques follow a “one-robot–one-task” paradigm, optimizing morphology and control specifically for a single task. This contrasts with biological organisms—such as humans—whose morphologies enable a broad repertoire of behaviors without task-specific redesign. A key challenge is the explicit co-design of generalist agents capable of performing multiple tasks, rather than narrowly tuned specialists [wang2023preco]. Advancing toward such generality raises complementary challenges, including the design of agents with intrinsic structural resilience, self-assessment, and self-repair capabilities [horibe2021regenerating, bilodeau2017self, islam2023review, xie2024soft], and methods that support the online co-adaptation of morphology and control in response to changing environmental conditions.
VI-A3 Sim-to-Real
As introduced in Section V-C, despite significant advances in simulation, closing the sim-to-real gap remains a major challenge for embodied co-design [eiben2021real, kriegman2020scalable, van2021influence]. Morphologies that perform well under idealized physics often fail in reality due to unmodeled factors such as frictional variability, material nonlinearity, complex contact dynamics, heterogeneity, sensor noise, actuation delays, and manufacturing imperfections. A widely adopted strategy is to perform the morphological search on cheap surrogates [he2024morph, du2025efficient], followed by validation and refinement in high-fidelity simulation or on hardware, incorporating human manufacturability checks before fabrication. Emerging techniques, including domain randomization [fadini2022simulation, schaff2023sim, lirobottool2023] and differentiable simulation with realistic material models [bacher2021design], aim to further mitigate this gap. From the fabrication perspective, realizing evolved morphologies—especially multi-material, non-planar, embedded-sensor assemblies, and highly deformable structures—remains challenging. Promising avenues for narrowing the gap include digital twins for tighter sim–real integration and reconfigurable modular hardware that enables rapid physical validation and iterative co-design.
VI-A4 Theories
Although numerous ECD approaches have emerged, few works have explored the theoretical foundations underlying this approach. A solid theoretical foundation can not only guide the design of more effective embodied agents but also clarify fundamental concepts, such as the brain-body trade-off. Furthermore, more practical metrics grounded in morphological computation could offer a holistic assessment of agent performance and prevent the unprincipled replacement of components with incremental technologies in common ECD pipelines. Hence, a more unified theory is indispensable to take the development of ECD to the next stage.
VI-A5 Benchmarks and Datasets
The ECD field still lacks widely adopted benchmarks, though platforms such as EvolutionGym [bhatia2021evolution] and SoftZoo [wang2023softzoo] represent important initial steps. Unlike traditional control optimization problems, benchmarking embodied co-design poses unique challenges, requiring diverse metrics to reflect the coupled nature of morphology and control. Expanding the infrastructure and establishing standardized ECD datasets could facilitate more rigorous and reproducible evaluations in ECD research.
VI-B Future Trends
Despite the challenges mentioned above in ECD, many promising directions remain worth exploring. From our perspective, future trends in ECD may include, but are not limited to, the following three main topics:
VI-B1 Human-AI Collaborative ECD
Rather than replacing humans, generative AI tools—such as LLMs and VLMs—can act as co-designers by proposing morphologies, aligning semantic descriptions with agent behaviors [matthews2020crowd], explaining trade-offs, drafting fabrication steps, and generating initial controllers based on high-level objectives [stella2023can, chen2025robots]. Meanwhile, human designers contribute domain knowledge, safety judgments, and ethical priors. Combining both approaches reduces design cycle time, increases the diversity of candidate designs, and makes ECD accessible to non-specialists. Recent works [song2025laser, ringel2025text2robot, fang2025robomore, gao2025vlmgineer, kulz2025design] have demonstrated preliminary pipelines that can transform natural language prompts into robot designs within minutes to hours. However, the full potential of these tools in reasoning and explainability to inform human designers has yet to be explored. Therefore, we believe that investigating such a human-ai pipeline is a promising direction.
VI-B2 Multi-Agent Collaborative ECD
This direction envisions a framework in which multiple autonomous machines—such as software agents, specialized LLMs, or learned policies—collaboratively design an embodied robot. Each agent contributes distinct capabilities, enabling interaction modes such as coordination, collaboration, or even competition. For example, a “morphology agent” proposes geometry, an “actuation agent” selects motors or stiffness parameters, a “control agent” designs controllers, and a “fabrication agent” ensures manufacturability. Together, these agents coordinate to generate coherent robot designs. Recent agentic frameworks have provided practical blueprints for such distributed co-design [shen2023hugginggpt, yang2024llm, qiu2024llm]. By decomposing the ECD problem into specialized roles, multi-agent systems reduce per-agent complexity, enable parallel exploration, and align naturally with modular hardware development pipelines.
VI-B3 Open-Ended and Lifelong ECD
As discussed in Section IV-D, open-ended co-design enables systems to continuously generate new tasks while concurrently optimizing morphologies and control policies to address them. This suggests that, when combined with lifelong-learning techniques [liu2021lifelong, mendez2023embodied, liu2023libero] that accumulate, reuse, and compose skills over long horizons, such systems can produce an ongoing stream of increasingly capable embodied machines. ECD, therefore, has the potential to evolve into a persistent creative ecology rather than remaining a bounded optimization problem.
VII Conclusion
The interplay among morphology, control, and task environment is fundamental not only for biological organisms but also for the development of intelligent embodied agents. In this survey, we review recent advances in embodied co-design (ECD) and organize them into a hierarchical taxonomy. This taxonomy provides a structured framework that enables researchers to contextualize current progress and identify opportunities for cross-disciplinary collaboration. We also review notable applications in both simulated and physical environments. Although current results indicate that co-designed agents outperform fixed-body baselines, many open questions remain unanswered. We highlight these challenges and discuss promising directions for future research. We hope that this survey serves as both a foundational reference for newcomers and a comprehensive resource for experts seeking deeper insights into ECD.
[Uncaptioned image] Yuxing Wang (Member, IEEE) received the B.Eng. degree in communication engineering from Southwest Minzu University (SMU) in 2020, and the M.Eng. degree in electronic and information engineering from Tsinghua University (THU) in 2023. He is currently pursuing a Ph.D. degree in control science and engineering at THU. Before that, he was a research assistant with THU and a research intern with Tencent AI Lab. His research interests include evolutionary learning and embodied intelligence, with a particular focus on evolutionary reinforcement learning for control optimization, and embodied co-design for modular robotic systems.

### v2

@@
 Brain-body co-evolution enables animals to develop complex behaviors in their environments.
-Inspired by this biological synergy, embodied co-design (ECD) has emerged as a transformative paradigm for creating intelligent agents—from virtual creatures to physical robots—by jointly optimizing their morphologies and controllers rather than treating control in isolation.
-This integrated approach facilitates richer environmental interactions and robust task performance.
+Inspired by this biological synergy, embodied co-design (ECD) has emerged as a transformative paradigm for creating intelligent agents—from virtual creatures to physical robots—through the joint optimization of morphology and control.
+This integrated approach facilitates richer environmental interactions and enhances task performance.
 In this survey, we provide a systematic overview of recent advances in ECD.
 We first formalize the concept of ECD and position it within related fields.
-We then introduce a hierarchical taxonomy: a lower layer that breaks down agent design into three fundamental components—controlling brain, body morphology, and task environment—and an upper layer that integrates these components into four major ECD frameworks: bi-level, single-level, generative, and open-ended.
-This taxonomy allows us to synthesize insights from more than one hundred recent studies.
+We then introduce a hierarchical taxonomy in which a lower layer breaks down agent design into three fundamental components—controlling brain, body morphology, and task environment—while an upper layer integrates these components into four major ECD frameworks: bi-level, single-level, generative, and open-ended.
+This taxonomy provides a unified lens through which we distill insights from over one hundred recent studies.
 We further review notable benchmarks, datasets, and applications in both simulated and real-world scenarios.
-Finally, we identify significant challenges and offer insights into promising future research directions.
-A project associated with this survey has been created at https://github.com/Yuxing-Wang-THU/SurveyBrainBody.
+Finally, we discuss significant challenges and offer insights into promising future research directions.
+A project associated with this survey has been established at https://github.com/Yuxing-Wang-THU/SurveyBrainBody.
 I Introduction
 Embodied intelligence (EI) seeks to integrate cognitive capabilities into autonomous machines, enabling them to learn from and adapt to complex environments [howard2019evolving].
-Situated at the intersection of artificial intelligence and robotics, EI has recently made significant advancements in fusing perception, decision-making, and actuation.
-This progress has empowered various robots to tackle previously intractable challenges, such as dynamic locomotion [lee2020learning, zhang2024whole], end-to-end navigation [zheng2024towards, schumann2024velma], and dexterous manipulation [chi2023diffusion, huang2023voxposer, wang2024dexcap], to name just a few.
-In traditional EI, an agent’s morphology (body) is typically hand-designed and treated as a fixed part of the environment, while only its controller (brain) is optimized for task-oriented behaviors.
-However, this separation has clear limitations.
-Since an agent’s performance is intertwined with both its morphology and control, a pre-specified morphology can be suboptimal: it constrains sensing and actuation, thereby narrowing the space of achievable behaviors.
+Situated at the intersection of artificial intelligence and robotics, EI has achieved significant advancements in fusing perception, decision-making, and actuation, empowering various robots to tackle previously intractable challenges, such as dynamic locomotion [lee2020learning, zhang2024whole], end-to-end navigation [zheng2024towards, schumann2024velma], and dexterous manipulation [chi2023diffusion, huang2023voxposer, wang2024dexcap], to name just a few.
+In traditional EI, an agent’s morphology (body) is typically hand-designed and treated as a fixed component of the environment, while only its controller (brain) is optimized for task execution.
+However, this separation presents clear limitations.
+A pre-specified morphology can constrain sensing and actuation, thereby narrowing the space of achievable behaviors.
 In other words, when morphological design prioritizes mechanical optimality without considering how it will be controlled, the burden of achieving coordinated behavior shifts almost entirely to the controller.
-This imbalance often becomes a bottleneck for task performance and raises a fundamental question: How should we design the optimal form of an embodied agent to effectively perform tasks in its environment?
-Figure 1: Differences between control design only and embodied co-design (ECD).
-The concept of morphological computation [paul2006morphological, muller2017morphological, pfeifer2009morphological, pfeifer2006body, mertan2024no] posits that a well-designed body can offload or simplify computations typically assigned to the brain, thereby facilitating more efficient behaviors.
+This imbalance often creates a bottleneck in task performance and raises a fundamental question: How should we design the optimal form of an embodied agent to effectively perform tasks in its environment?
+Fig. 1: Differences between control design only and embodied co-design.
+The concept of morphological computation [paul2006morphological, muller2017morphological, pfeifer2009morphological, pfeifer2006body, mertan2024no] posits that a well-designed body can simplify computations typically assigned to the brain, thereby facilitating more efficient behaviors.
 This perspective introduces a brain-body dynamic: changes in morphology necessitate corresponding adaptations in control to exploit newly acquired capabilities, while environmental feedback can, in turn, guide further refinement of morphology.
-These synergies have motivated various efforts to improve the current design procedure for embodied agents.
-As a result, embodied co-design (ECD), an integrated design framework that co-optimizes an agent’s morphology with its control mechanism to enhance task performance, becomes a critical topic in EI.
-ECD was originally studied in the Evolutionary Robotics (ER) community [harvey2005evolutionary, silva2016open], where both robotic behaviors and morphologies are concurrently optimized using Evolutionary Algorithms (EAs).
-However, this topic remains emergent for EI, particularly as agents and their tasks continue to grow in complexity and diversity.
-In contrast to controller-only optimization, ECD introduces a vast and often combinatorial design space for morphology-control configurations, accompanied by fragile interdependencies.
-Specifically, evaluating a candidate morphology typically requires re-training the controller from scratch, since even minor morphological changes can invalidate previously learned controllers.
-Moreover, variations in environmental conditions, design constraints, and task objectives produce heterogeneous feedback signals, further complicating the co-design process, as illustrated in Fig.
-1.
-Consequently, ECD is highly non-convex and computationally expensive, highlighting the increasing necessity for automated ECD techniques over traditional manual engineering.
-Figure 2: Co-designed embodied agents in both virtual and real environments.
-Example images numbered from (a) to (p) are adapted from [sims1994competition, hornby2001evolving, shim2003generating, Veenstra2023, ma2021diffaqua, ha2019reinforcement, luck2020data, zhao2020robogrammar, gupta2021embodied, xu2021accelerated, cheney2013unshackling, zardini2021seeking, medvet2021biodiversity, wang2023curriculum, wang2023softzoo, yue2025toward], respectively.
-Example images numbered from (q) to (y) are adapted from [nygaard2021environmental, revolve2, Strgar-RSS-24, xie2024map, schaff2022soft, liu_voxcraft_2020, kobayashi2024computational, matthews2023efficient, cheng2024structural], respectively.
-Motivation: Recent breakthroughs in domains such as evolutionary algorithms [davis2023subtract, bergonti2024co, ferigo2025totipotent], reinforcement learning (RL) [yuan2022transformact, wang2023curriculum, dongleveraging2024, lu2025bodygen], differentiable simulation [xu2021accelerated, ma2021diffaqua, Strgar-RSS-24], and large models [gao2025vlmgineer, fang2025robomore, song2025laser] have led to a surge of novel ECD approaches, rapidly reshaping the capabilities of co-design systems.
-However, their assumptions, representations, computational pipelines, evaluation benchmarks, and limitations have not been systematically reviewed, which makes it difficult for researchers to develop a coherent understanding of the ECD landscape.
-A comprehensive survey is increasingly needed.
-Our goal is to consolidate these frontier developments, clarify conceptual foundations, highlight emerging trends, and identify open challenges that will shape the next generation of embodied intelligence research.
+These synergies have motivated various efforts to improve the current agent design procedure.
+As a result, embodied co-design (ECD), an integrated framework that co-optimizes an agent’s morphology with its control mechanism to enhance task performance, becomes a critical topic in EI.
+ECD was initially explored in the Evolutionary Robotics (ER) community [harvey2005evolutionary, silva2016open], where both robot morphology and behavior are jointly optimized using Evolutionary Algorithms (EAs) [harvey2005evolutionary].
+However, this topic remains emergent for EI, particularly as agents and their tasks continue to increase in complexity and diversity.
+Compared to controller-only optimization, ECD introduces a vast, often combinatorial design space for morphology-control configurations, accompanied by fragile interdependencies.
+Specifically, evaluating a candidate morphology typically requires re-training the controller from scratch, as even minor morphological changes can invalidate previously learned controllers.
+Variations in environmental conditions, design constraints, and task objectives produce heterogeneous feedback signals, further complicating the co-design process (Fig.
+1).
+Consequently, ECD problems are highly non-convex and computationally expensive, underscoring the growing need for automated methodologies instead of traditional manual engineering.
+Motivation: Recent breakthroughs in domains such as evolutionary algorithms [davis2023subtract, bergonti2024co, ferigo2025totipotent], reinforcement learning (RL) [yuan2022transformact, wang2023curriculum, dongleveraging2024, lu2025bodygen], differentiable simulation [xu2021accelerated, ma2021diffaqua, Strgar-RSS-24], and large models [gao2025vlmgineer, fang2025robomore, song2025laser] have led to a surge of novel ECD approaches that rapidly reshape the capabilities of co-design systems.
+However, their assumptions, representations, computational pipelines, evaluation benchmarks, and limitations have not been systematically reviewed, making it difficult for researchers to develop a coherent understanding of the ECD landscape.
+Our goal is to clarify the conceptual foundations, consolidate frontier developments, highlight emerging trends, and identify open challenges that will shape the next generation of research in EI.
 Taxonomy: To enhance readability for researchers across disciplines, we propose a two-layer hierarchical taxonomy in Section III.
@@
 Building upon this foundation, the upper layer categorizes existing ECD methods into four major frameworks: bi-level (Section IV-A), single-level (Section IV-B), generative (Section IV-C), and open-ended (Section IV-D), with further subcategories identified within each.
-This taxonomy establishes a common vocabulary and provides a conceptual map for embodied co-design, facilitating cross-domain understanding.
+This taxonomy establishes a common vocabulary and provides a conceptual map for ECD, facilitating cross-domain understanding.
 It also encourages researchers to explore beyond their primary areas.
-For instance, specialists in RL for control optimization may discover how their methodologies align with a specific ECD framework, in conjunction with previously unconsidered morphological evolution, thereby opening pathways for new interdisciplinary collaborations.
-Scope: ECD bridges robotics, artificial intelligence, control theory, and cognitive science.
-Guided by our proposed taxonomy, we focus on analyzing the design spaces and representation methods of morphology, control, and task environments.
-More importantly, we investigate how these components are jointly coordinated within different co-design frameworks.
-Our aim is to provide an organizational perspective that clarifies the conceptual and technical interrelations among various ECD works, rather than exhaustively reviewing any particular algorithmic subfield (e.g., RL for control optimization) or engineering detail (e.g., 3D printing for morphological fabrication).
-Related Works: Most prior surveys either provide broad overviews of embodied intelligence [long2025survey, duan2022survey, sun2024comprehensive, liu2025embodied] or focus on specific subtopics, such as embodied learning [liang2025large], manipulation [li2025survey], and navigation [xiong2025sensing].
-These works primarily emphasize sensorimotor learning, while ECD has received limited attention.
-Partial reviews of ECD exist within specific domains, including soft robotics [zhao2024exploring, chen2020design, stroppa2024optimizing], humanoid robots [liu2025embracing, yue2025toward], evolutionary robotics [doncieux2014beyond, prabhu2018survey, alattas2019evolutionary, pandey2022accessible, strgar2024evolution], and artificial life [lai2021virtual, casadei2023artificial, ha2022collective].
-However, they lack a structured analytical framework for systematically comparing co-design methodologies and do not incorporate recent advances, benchmarks, and emerging applications.
-In contrast, our survey provides a comprehensive and up-to-date analysis of ECD, covering a wide range of agent types and offering in-depth discussions on key challenges and future directions.
-Contributions: To the best of our knowledge, this paper presents the first comprehensive survey of ECD. The main contributions are summarized below:
+For instance, specialists in RL for control optimization may discover how their methodologies align with a specific ECD framework, in conjunction with previously unconsidered morphological evolution, thereby opening pathways for new collaborations.
+Scope: Embodied co-design bridges robotics, artificial intelligence, control theory, and cognitive science.
+Guided by our proposed taxonomy, we analyze the design spaces and representation methods for morphology, control, and task environments, with a particular emphasis on how these components are coordinated within various co-design frameworks.
+Rather than exhaustively reviewing any specific algorithmic subfield (e.g., RL for control optimization) or engineering detail (e.g., 3D printing for morphological fabrication), this work provides an organizational perspective that clarifies the conceptual and technical interrelations among different ECD works.
+Related Works: Most prior surveys either provide broad overviews of embodied intelligence [long2025survey, duan2022survey, sun2024comprehensive, liu2025embodied] or focus on isolated subfields such as embodied learning [liang2025large], manipulation [li2025survey], and navigation [xiong2025sensing].
+These works emphasize sensorimotor integration, while attention to ECD has been largely limited.
+To our knowledge, partial reviews of ECD exist in specific domains, including soft robotics [zhao2024exploring, chen2020design, stroppa2024optimizing], humanoid robots [liu2025embracing, yue2025toward], evolutionary robotics [doncieux2014beyond, prabhu2018survey, alattas2019evolutionary, pandey2022accessible, strgar2024evolution], and artificial life [lai2021virtual, casadei2023artificial, ha2022collective].
+However, these reviews lack a structured taxonomy for systematically comparing ECD methods; they do not incorporate recent advances, benchmarks, and emerging applications.
+Features: Unlike previous works, our survey is the first to provide a systematic, method-centric analysis of ECD across various agent types, optimization paradigms, and application domains.
+Rather than treating morphology as a secondary concern or focusing on isolated subfields, we, for the first time, introduce a two-layer hierarchical taxonomy that explicitly unifies the design process of embodied agents, thus enabling principled comparisons across various ECD methods.
+We further incorporate recent advances—including differentiable simulation, quality-diversity search, latent and foundation-model-based generation—together with benchmarks, datasets, and real-world testbeds that have been largely absent from related reviews.
+By combining conceptual unification with practical evaluation criteria and open challenges, this survey offers both a coherent theoretical framework and an up-to-date technical reference for the ECD community.
+Contributions: The main contributions of our survey are summarized below:
   1. 1.
-    We conceptualize modeling approaches to ECD from three complementary perspectives: bi-level optimization, sequential decision-making, and model-based optimization (Section II).
-Together, this set of lenses provides a multifaceted view of the ECD landscape.
+    We conceptualize ECD from three complementary perspectives: bi-level optimization, sequential decision-making, and model-based optimization (Section II), providing a multifaceted view of the ECD problem.
   2. 2.
-    We contribute a hierarchical taxonomy (Section III) that organizes co-design spaces and identifies key design choices.
-The proposed taxonomy aims to deepen researchers’ understanding of the fundamental components involved in embodied co-design.
+    We contribute a hierarchical taxonomy (Section III) that organizes co-design spaces and identifies key design choices, offering a unified framework for understanding the fundamental components of ECD.
   3. 3.
-    We present a systematic technical review (Section IV) of major advances in ECD over the past five years, analyzing representative methodologies and their core principles.
+    We present a systematic technical review (Section IV) of major advances in ECD over the past five years, analyzing representative methods and their core principles.
   4. 4.
-    We survey simulation platforms and real-world testbeds commonly used in ECD research (Section V), highlighting their capabilities, representative demonstrations, and application domains.
+    We survey simulation platforms, datasets, and real-world testbeds in ECD research (Section V), highlighting their capabilities and representative demonstrations.
   5. 5.
-    We identify current bottlenecks and existing gaps in ECD, outlining promising research directions (Section VI) that may benefit a broad range of communities and help advance the state of the art.
-The remainder of this survey is organized as follows.
-We begin with the background of ECD in Section II, which includes an introduction to its fundamental elements, mathematical definitions, and related topics.
-In Section III, we introduce our hierarchical taxonomy.
-In Section IV, we elaborate on four co-design frameworks based on the proposed taxonomy.
-We discuss notable benchmarks and contemporary applications of ECD in Section V.
-Finally, we outline the challenges of modern ECD and provide future perspectives in Section VI.
+    We identify current bottlenecks in ECD and outline promising research directions (Section VI) that may benefit a broad range of communities and help advance the state of the art.
 II Background: Formalizing Embodied Co-Design
@@
 An embodied agent, such as a robot, can be characterized by its body morphology and controlling brain.
-The body morphology $\omega\in\Omega$ defines the agent’s physical interface with the outer surroundings, including its structural topology, material properties, and the spatial arrangement of sensors and actuators.
+The body morphology $\omega\in\Omega$ defines the agent’s physical interface with its environment, encompassing structural topology, material properties, and the spatial configuration of sensors and actuators.
 The controlling brain $\pi\in\Pi$ maps observed states to action commands, driving the body to exert forces that interact with the task environment.
@@
 Following the sequential workflow commonly adopted in engineering practice, we divide the overall co-design process into two main stages: the design stage $\mathcal{D}$ and the control stage $\mathcal{C}$.
-As a result, the ECD problem can be represented as a $9$-tuple $\left\langle\mathcal{S_{D}},\mathcal{S_{C}},\mathcal{A_{D}},\mathcal{A_{C}},\mathcal{F_{D}},\mathcal{F_{C}},\mathcal{R_{D}},\mathcal{R_{C}},\gamma\right\rangle$, where $\gamma\in(0,1]$ denotes the reward discount factor, and the remaining elements correspond to the relevant design spaces.
+As a result, the ECD problem can be represented as a $9$-tuple $\left\langle\mathcal{S_{D}},\mathcal{S_{C}},\mathcal{A_{D}},\mathcal{A_{C}},\mathcal{F_{D}},\mathcal{F_{C}},\mathcal{R_{D}},\mathcal{R_{C}},\gamma\right\rangle$, where $\gamma\in(0,1]$ denotes the reward discount factor, with the remaining elements corresponding to the relevant design spaces.
 For clarity, explicit representations of constraints—such as morphological or task-specific constraints—are omitted and incorporated implicitly within their respective design spaces.
@@
 In the subsequent control stage, a policy $\pi_{c}(a_{t}^{c}|s_{t}^{c},\omega)$ conditioned on $\omega$ directs the generated morphology to perform tasks, following the environmental transition dynamics $F_{c}(s^{c}_{t+1}|s^{c}_{t},a^{c}_{t},\omega)$.
-Based on the above notations, the single-level objective $R$ of ECD is now expressed as:
+Based on the aforementioned notations, the single-level objective $R$ of ECD is now expressed as:
 $\begin{split}R=\mathbb{E}_{\pi_{d},\pi_{c}}\left[\sum_{t=0}^{T_{d}}\gamma^{t}r^{d}_{t}+\sum_{t=T_{d}+1}^{\infty}\gamma^{t}r^{c}_{t}\right]\end{split}$ (2)
 where $r^{d}_{t}\in\mathcal{R_{D}}$ and $r^{c}_{t}\in\mathcal{R_{C}}$ denote the rewards obtained during the design and control stages, respectively.
-Leveraging effective gradient approximation techniques in RL [schulman2017proximal, silver2014deterministic, ciosek2020expected], both the design policy and the control policy can be updated iteratively through trial-and-error interactions.
+By leveraging effective gradient approximation techniques in RL [schulman2017proximal], both the design policy and the control policy can be iteratively updated through trial-and-error interactions.
 II-A4 Model-Based Optimization View
 It is noteworthy that, in the perspectives discussed above, the environmental transition model $F_{c}(s^{c}_{t+1}|s^{c}_{t},a^{c}_{t},\omega)$ is typically treated as a black box.
-However, as discussed in Section IV-B3, many ECD approaches construct $F_{c}$ as either a differentiable simulator [hu2019chainqueen, du2021diffpd, hu2019difftaichi] or a learned dynamics model [navez2025modeling].
+However, as discussed in Section IV-B3, many ECD approaches construct $F_{c}$ either as a differentiable simulator [hu2019chainqueen, du2021diffpd, hu2019difftaichi] or as a learned dynamics model [navez2025modeling].
 Conceptually, the transition can be written as an abstract function $s^{c}_{t+1}=F_{c}(s^{c}_{t},a^{c}_{t},\omega)$, where $a^{c}_{t}$ denotes a vector of control signals, such as the joint torques applied during that time-step.
@@
 $\partialderivative{\mathcal{L}}{\omega}=\left(\partialderivative{\mathcal{L}}{s^{c}_{t+1}}\right)\left(\partialderivative{F_{c}}{\omega}\right)$ (5)
-This model-based formulation facilitates end-to-end, gradient-based optimization for inverse design in ECD and provides substantial advantages in sample efficiency.
+This model-based formulation facilitates end-to-end, gradient-based optimization for inverse design in ECD, providing substantial advantages in sample efficiency.
 II-B Philosophical Foundations
@@
 II-C Related Topics
-Evolutionary Robotics (ER) originally focused on evolving robot behaviors [bongard2013evolutionary, doncieux2015evolutionary, medvet2023evolutionary].
-For example, behavioral trees have been evolved using genetic programming [scheper2016behavior].
+Evolutionary Robotics (ER) originally focused on the evolution of robot behaviors [bongard2013evolutionary, scheper2016behavior].
 Over time, the scope expanded to include the evolutionary design of both robotic hardware and control policies.
 A seminal contribution is Sims’ work [sims1994evolving], which evolves 3D virtual creatures by jointly optimizing their graph-based morphologies and neural controllers.
-Contemporary ECD builds on these insights but extends beyond traditional evolutionary robotics, incorporating a wider range of optimization and learning techniques to improve scalability, generalization, and applicability.
+Contemporary ECD builds upon these insights while extending beyond traditional ER by incorporating a broader array of representation and learning techniques to enhance scalability, generalization, and applicability.
 Artificial Life (A-Life) investigates the co-evolution of morphology and control, aiming to recreate life-like processes in virtual environments [moreno2005agency, aguilar2014past].
-While A-Life shares the central concept of brain–body–environment coupling with ECD, it prioritizes the exploration of biological phenomena such as self-organization [gershenson2020self], adaptation [bedau2003artificial], and evolutionary dynamics [ito2016population, pilat2012evolution] in nature, rather than engineering practical embodied agents with specific functionalities.
+While A-Life shares the central concept of brain–body–environment coupling with ECD, it prioritizes the exploration of biological phenomena such as self-organization [gershenson2020self], adaptation [bedau2003artificial], and evolutionary dynamics [ito2016population] in nature, rather than engineering practical embodied agents with specific functionalities.
 Computational Robot Design leverages algorithmic tools to automate the synthesis of robots, typically following a sequential design process.
-By formulating complex constrained optimization problems, this approach employs techniques such as topology optimization [sigmund2000topology, sigmund2013topology] and CAD integration [tian2022assemble, willis2022joinable] to address structural feasibility, material selection, electronics integration, fabrication constraints, and dynamic modeling.
-Control, however, is often simplified or treated as a secondary component.
-In contrast, ECD is a broader concept that tightly couples morphology and control within a unified optimization loop, aiming not only to generate feasible body designs but also to produce adaptive, task-capable behaviors.
+This approach formulates complex constrained optimization problems and employs techniques such as topology optimization [sigmund2000topology, sigmund2013topology] and CAD integration [tian2022assemble, willis2022joinable] to address issues related to structural feasibility, material selection, electronics integration, and fabrication constraints.
+Control, however, is often simplified or regarded as a secondary component.
+In contrast, ECD is a broader concept that tightly integrates morphology and control within a unified optimization loop, aiming to generate not only feasible body designs but also adaptive, task-capable behaviors.
 Robot Learning aims to equip robots with the ability to autonomously acquire perceptual, decision-making, and action skills, rather than relying solely on hand-engineered solutions [kroemer2021review].
-It is closely connected to ECD, as learning-based control optimization serves as a central mechanism for co-design in tasks such as locomotion, manipulation, and navigation.
-Recent advances in robot learning have integrated model-based approaches [grandia2023perceptive], reinforcement learning [matsuo2022deep, singh2022reinforcement], large-model-based reasoning [xiao2025robot], imitation learning [hussein2017imitation, fang2019survey], and self-supervised methods [nair2020contextual, jeong2020self, kahn2021badgr] to enhance high-dimensional sensorimotor loops.
+It is closely connected to ECD, as learning-based control optimization serves as a central mechanism for solving tasks such as locomotion, manipulation, and navigation.
+Recent advances in robot learning have integrated model-based approaches [grandia2023perceptive], reinforcement learning [matsuo2022deep], large-model-based reasoning [xiao2025robot], imitation learning [hussein2017imitation], and self-supervised methods [jeong2020self] to enhance high-dimensional sensorimotor loops.
 However, most robot learning frameworks assume a fixed morphology.
@@
 III Taxonomy: Decomposing Embodied Co-Design
-Figure 3: An overview of the embodied co-design landscape including foundational elements and co-design frameworks.
+Fig. 2: An overview of the embodied co-design landscape including foundational elements and co-design frameworks.
 ECD operates within an expansive workspace shaped by the interplay of control, morphology, and task environment.
 As shown in Fig.
-3, the lower layer of our taxonomy organizes this space by complexity, while the upper layer defines four ECD frameworks, each with sub-categories that guide the co-design process.
-In the following sections, we detail each component, highlight recent advances, discuss common constraints, summarize key evaluation metrics, and illustrate a representative ECD pipeline using a prototype.
+2, our taxonomy organizes these fundamental elements into a lower layer, while the upper layer defines four ECD frameworks, each with sub-categories that guide the co-design process.
+In the following sections, we detail each component, highlight recent advances, discuss common constraints, summarize key evaluation metrics, then illustrate a representative ECD pipeline using a prototype.
 III-A Controlling Brain
@@
 III-A1 Design Space
-The design space encompasses both the ways controllers are parameterized and the modes in which they operate.
-Controller Parameters can be predefined—such as action primitives [iii2021taskagnostic], fixed waveform generators [cheney2013unshackling], and hand-crafted gaits [kobayashi2024computational]—or be fully learnable, optimized through methods such as trajectory optimization [spielberg2017functional, bergonti2024co], reinforcement learning, and imitation learning.
-Modern ECD approaches typically employ learnable controllers, which enable adaptation to varying morphologies and support flexible, real-time improvements in behavior.
+The design space of controlling brain encompasses both the ways by which controllers are parameterized and the modes in which they operate.
+Controller Parameters can be predefined—such as action primitives [iii2021taskagnostic], fixed waveform generators [cheney2013unshackling], and hand-crafted gaits [kobayashi2024computational]—or they can be fully learnable, optimized through methods such as trajectory optimization, reinforcement learning, and imitation learning.
+Modern ECD approaches typically employ learnable controllers that enable adaptation to varying morphologies and support flexible, real-time improvements in behavior.
 Operating Mode defines the fundamental strategy by which a control system executes its commands and manages its actions.
@@
 The controller representation defines the mapping between the control policy and the agent’s physical structure.
-Classical approaches adopt either individual or universal representations, with recent research exploring hybrid variants that combine their strengths.
+Classical approaches adopt either individual or universal representations, while recent research explores hybrid variants that combine their strengths.
 Individual representations directly bind the controller architecture or dimensionality to a specific morphology.
-Examples include PID controllers with per-joint gains that account for the inertia and stiffness of specific body segments [ang2005pid, borase2021review], as well as Model Predictive Control (MPC) when an accurate dynamic model is available [schwenzer2021review, carron2019data].
-Another example is Central Pattern Generators (CPGs), which mirror a robot’s limb topology to generate coordinated rhythmic motion [ijspeert2008central, bliss2012central, deshpande2023deepcpg].
+Examples include low-level PID controllers with per-joint gains that account for the inertia and stiffness of specific body segments [ang2005pid, borase2021review], as well as Model Predictive Control (MPC) when an accurate dynamic model is available [schwenzer2021review, carron2019data].
+Another example is Central Pattern Generators (CPGs), which mirror a robot’s limb topology to produce coordinated rhythmic motion [bliss2012central, deshpande2023deepcpg].
 Higher-level neural controllers, such as Multi-Layer Perceptrons (MLPs), adopt input/output sizes tied to a robot’s sensors and actuators.
-Individualized controllers are appropriate for topology-fixed agents that share the same morphological structures but differ solely in their kinematic or dynamic parameters.
-This specialization, however, reduces transferability: even minor morphological changes—such as the addition of a joint or modifications to limb geometry—typically require controller restructuring or retraining.
-Consequently, per-agent representations struggle to generalize across heterogeneous state and action spaces, restricting scalability.
+Individualized controllers are suitable for topology-fixed agents that share identical morphological structures but differ solely in their kinematic or dynamic parameters.
+This specialization, however, reduces transferability; even minor morphological changes—such as the addition of a joint or modifications to limb geometry—typically necessitate restructuring the controller.
+Consequently, per-agent representations struggle to generalize across heterogeneous state and action spaces.
 Universal representations aim to generalize across diverse morphologies.
-Representing a body as a kinematic graph naturally invites Graph Neural Networks (GNNs) [munikoti2023challenges]: their message-passing mechanisms enable a single policy to generalize across graphs with varying node counts, thereby supporting scalable control across different robotic embodiments.
+Representing a body as a kinematic graph naturally invites the use of Graph Neural Networks (GNNs) [munikoti2023challenges]; their message-passing mechanisms enable a single policy to generalize across graphs with varying node counts, thereby supporting scalable control across different robotic embodiments.
 Representative GNN-based controllers include Nervernet [wang2018nervenet], DGN [pathak2019learning], Whitman et al.
 [whitman2023learning], SMP [huang2020one], MAT [li2024mat], MG [pan2025morphology], and GCNT [ijcai2025p972].
-Subsequent research in cognitive science has further shown that modularity can enhance control generalization [schilling2024modularity, nadizar2022collective, pigozzi2022evolving, pigozzi2022evolving2, dong2022low, zhang2024composer, tjandrasuwita2024memo].
-To capture the long-range dependencies that GNNs inherently struggle with, transformer-based controllers incorporate global attention over all body parts.
-This idea has been explored in numerous embodied AI systems, such as Amorpheus [kurin2020my], MetaMorph [gupta2022metamorph], SWAT [hong2021structure], AnyMorph [trabucco2022anymorph], Xiong et al.
+Subsequent research in cognitive science has further demonstrated that modularity can enhance control generalization [schilling2024modularity].
+To capture the long-range dependencies that GNNs inherently struggle to address, transformer-based controllers incorporate global attention over all body parts.
+This concept has been explored in numerous embodied AI studies, such as Amorpheus [kurin2020my], MetaMorph [gupta2022metamorph], SWAT [hong2021structure], AnyMorph [trabucco2022anymorph], Xiong et al.
 [xiong2023universal], Body Transformer [sferrazza2024body], HeteroMorpheus [hao2024heteromorpheus], and URMA [bohlinger2024one].
 Beyond ECD, universal controllers have also shown strong utility in transfer learning [chen2018hardware, peng2018sim, cho2024meta], meta-learning [nagabandi2018learning, ghadirzadeh2021bayesian], and cross-embodiment learning [yang2024pushing, shafiee2024manyquadrupeds, wang2024sparse, chen2024mirage, patel2025get, mertan2024towards, feng2023genloco, luo2024moral].
-This underscores their potential as a unifying representation for embodied AI.
-Hybrid representations seek to balance specialization and generality by decomposing the controller or control mechanism into a shared, morphology-agnostic module that encodes universal motor and perceptual priors [bjelonic2023learning], and a morphology-specific module that adapts outputs to specific embodiments [chen2023evolving, luck2020data, luck2021robot, sliacka2023co, sun2023co].
-Hybrid representations can be seamlessly integrated into various pipelines, such as pretraining-finetuning [yue2025toward, jin2025evolutionary] and meta-learning [belmonte2022meta], to facilitate cross-morphology generalization and faster adaptation during the ECD process.
+Hybrid representation seeks to balance specialization and generalization by decomposing the controller or control mechanism into a shared, morphology-agnostic module that encodes universal motor and perceptual priors, and a morphology-specific module that adapts outputs to specific embodiments [bjelonic2023learning, chen2023evolving, luck2020data, luck2021robot, sliacka2023co, sun2023co].
+Hybrid representations can be seamlessly integrated into various pipelines, such as pretraining-finetuning [yue2025toward, jin2025evolutionary] and meta-learning [belmonte2022meta], to facilitate cross-morphology generalization and accelerate adaptation during the ECD process.
 III-A3 Discussion
-Training a dedicated controller for each morphology yields an accurate estimate of that body’s potential by fully exploiting its passive dynamics and affordances.
-However, this approach is computationally expensive and biases the search toward morphologies that are easier to learn.
-Empirical results from [kvalsund2022centralized, mertan2023modular, chiappa2022dmap, luo2023comparison] have demonstrated that universal controllers can promote better synchronization, greater morphological diversity, and reduced premature convergence.
-However, they also require more comprehensive evaluation metrics to “diagnose” whether performance gains arise from enhanced controller capabilities or advantageous morphological adaptations.
-These observations raise an important question: must one optimize a controller to evaluate morphology?
+Training a dedicated controller can provide an accurate estimate of a morphology’s potential by fully exploiting its passive dynamics and affordances.
+Empirical results from [kvalsund2022centralized, mertan2023modular, chiappa2022dmap, luo2023comparison] have demonstrated that well-designed controllers can enhance morphological diversity and mitigate premature convergence.
+However, they require more comprehensive evaluation efforts to “diagnose” whether performance gains arise from enhanced controller capabilities or advantageous morphological adaptations.
+This raises an important question: must one optimize a controller to evaluate morphology?
 Recent work in ECD suggests that the answer may be no.
@@
 Discrete parameters typically arise in modular systems where agents are composed of a finite set of building blocks, topologies, or design configurations.
-Continuous parameters govern real-valued attributes such as link lengths, shape, and material properties.
+Continuous parameters govern real-valued attributes such as link lengths, shapes, and material properties.
 More general embodied systems adopt a hybrid scheme in which certain choices, such as whether to attach a limb, are discrete, whereas attributes like limb length are continuous.
-This formulation balances structural flexibility with optimization tractability, enabling scalable search over high-dimensional morphologies.
+This formulation balances structural flexibility with optimization tractability, enabling scalable searches over high-dimensional morphologies.
 Topological Structure further defines the scope of morphological exploration.
 In fixed topology settings, such as those involving bipedal [ha2019reinforcement], quadrupedal [belmonte2022meta], or manipulator skeletons [cai2023task2morph], optimization is constrained to parametric variations within a predefined template.
-This approach ensures stability and facilitates hardware deployment, but it restricts the discovery of novel morphologies.
-In contrast, unfixed topology settings allow for arbitrary structural modifications, ranging from voxel-based [bhatia2021evolution] to graph-based agents [yuan2022transformact] capable of growing, pruning, or reconfiguring their skeletons.
-While enabling open-ended exploration and emergent specialization, these approaches also raise concerns regarding physical plausibility, degeneracy, and manufacturability.
+This approach ensures stability and facilitates hardware deployment; however, it restricts the discovery of novel morphologies.
+In contrast, unfixed topology settings permit arbitrary structural modifications, ranging from voxel-based [bhatia2021evolution] to graph-based agents [yuan2022transformact] capable of growing, pruning, or reconfiguring their skeletons.
+While enabling open-ended exploration, these approaches also raise concerns regarding manufacturability.
 III-B2 Representation
@@
 Indirect encodings compress the agent’s morphological information into compact descriptors or generative rules, which are then expanded into complete morphologies through developmental or decision-making processes.
-Examples include L-Systems [veenstra2022effects], generative grammars [zhao2020robogrammar, zhao2022automatic, zharkov2024synergizing], Neural Cellular Automata (NCA) [veenstra2020different, wang2023curriculum], Compositional Pattern-Producing Networks (CPPNs) [cheney2013unshackling], and reinforcement learning-based decision-making [yuan2022transformact, wang2023curriculum].
+Examples include L-systems [veenstra2022effects], generative grammars [zhao2020robogrammar], Neural Cellular Automata (NCA) [veenstra2020different], Compositional Pattern-Producing Networks (CPPNs) [cheney2013unshackling], and reinforcement learning-based decision-making [yuan2022transformact, wang2023curriculum].
 Indirect encodings typically yield smoother fitness landscapes [thomson2024understanding], facilitate the escape from morphological local optima, and produce modular, symmetric morphologies with enhanced transferability.
-Hybrid encodings aim to combine the strengths of both approaches by explicitly parameterizing part of the morphology while generating the remainder procedurally or via learned models [dongleveraging2024, ringel2025text2robot, saito2024effective, le2024improving, wang2023softzoo, mertan2025controller].
+Hybrid encodings aim to combine the strengths of both approaches by explicitly parameterizing part of the morphology while generating the remainder procedurally or via learned models [dongleveraging2024, ringel2025text2robot, saito2024effective, le2024improving, mertan2025controller].
 This design facilitates both global exploration and local refinement.
@@
 III-B3 Discussion
-The trade-offs between direct and indirect morphological encodings remain an open question in ECD.
+The trade-off between direct and indirect morphological encodings remains an open question in ECD.
 Prior work [pigozzi2023morphology] has reported that direct encodings often yield faster learning and more consistent convergence.
-In contrast, indirect encodings tend to generate more diverse, spatially coherent, and globally varying morphologies, although the corresponding learning improvements can be less predictable.
-Comparative studies [komosinski2001comparison, veenstra2017evolution] further demonstrated that indirect encodings are advantageous during early exploration due to their compactness and genetic reuse, whereas direct encodings may outperform them in later optimization stages.
-These insights underscore the potential of employing hybrid morphology representation strategies that begin with indirect encodings for broad exploration, followed by a transition to direct encodings for fine-grained refinement.
+In contrast, indirect encodings generate more diverse, spatially coherent, and globally varying morphologies; however, their impact on learning performance is less predictable.
+Comparative studies [komosinski2001comparison, veenstra2017evolution] further demonstrate that indirect encodings provide advantages during early exploration due to their compactness and genetic reuse, while direct encodings may excel during later optimization stages.
+These insights underscore the potential of hybrid strategies that begin with indirect encodings for broad exploration and subsequently transition to direct encodings for fine-grained refinement.
 III-C Task Environment
@@
 III-C1 Design Space
-The design space of the task environment specifies whether environmental parameters remain fixed or vary during co-design, and how feedback signals are delivered.
+The design space of task environment specifies whether environmental parameters remain fixed or vary during co-design and how feedback signals are delivered.
 Environmental Parameters are typically categorized into predefined and modifiable types.
 In predefined settings, all environmental attributes, such as terrain roughness and object placement, are fixed, resulting in stable and repeatable dynamics.
-Such static environments are widely utilized to isolate environmental factors, thereby enabling the targeted specialization of morphology and control.
-However, this specialization often results in brittle solutions that degrade under unmodeled perturbations or deployment beyond training conditions.
-In contrast, treating the environment as modifiable—for example, by sampling from a distribution of scenario parameters, introducing curriculum-based difficulty [wang2019poet], or employing non-stationary task sequences—encourages the emergence of agents that generalize more effectively [bossens2020qed, auerbach2014environmental, urbina2024brain].
+However, this specialization often results in brittle solutions that degrade under unmodeled perturbations or when deployed beyond training conditions.
+In contrast, treating the environment as modifiable—for example, by sampling from a distribution of scenario parameters, introducing curriculum-based difficulty, or employing non-stationary task sequences—encourages agents to generalize more effectively.
 Observability is a fundamental property of the task environment, as it determines the fidelity and completeness of the information an agent can perceive about the underlying system state.
 In a fully observable setting, the agent’s observations constitute a sufficient statistic of the true environment state.
-This condition simplifies learning and planning, as seen in classical control and various reinforcement learning tasks.
-Conversely, partial observability arises when agents are provided with incomplete [schaff2022soft], noisy [bjelonic2023learning], or local per-body/per-joint information [wang2023curriculum, schaff2022n, yuan2022transformact], which decouples immediate observations from the global state.
-Under such conditions, the agent must maintain an internal belief state or memory to integrate temporal observations.
-The degree of observability, therefore, shapes an agent’s perceptual and control mechanisms, ranging from simple reactive policies to sophisticated memory architectures or active perception strategies.
-It is also a critical axis for benchmarking an agent’s robustness and generalization in real-world scenarios.
+This condition simplifies learning and planning, as observed in classical control and various reinforcement learning tasks.
+Conversely, partial observability occurs when agents receive incomplete [schaff2022soft], noisy [bjelonic2023learning], or local per-body/per-joint information [wang2023curriculum, schaff2022n, yuan2022transformact], which decouples immediate observations from the global state.
+In such cases, the agent must maintain an internal belief state or memory to integrate temporal observations.
+The degree of observability, therefore, inherently influences an agent’s perceptual and control mechanisms; it is also a critical axis for benchmarking an agent’s robustness and generalization in real-world scenarios.
 III-C2 Representation
-This defines the objectives, constraints, and evaluative (reward) signals that guide both the learning dynamics and the emergent morphology–control solutions in ECD.
+The environmental representation defines the objectives, constraints, and evaluative (reward) signals that guide the optimization dynamics of ECD.
 We broadly categorize these representations into three types: task-driven, task-agnostic, and hybrid.
@@
 The task environment shapes, and is shaped by, the agent’s morphology and control.
-Prior studies have demonstrated that robots co-designed on uneven terrain tend to develop more stable, balanced, and limb-rich morphologies, whereas those created on flat terrain favor simpler forms [miras2020environmental, miras2019effects, spanellis2021environment, corucci2018evolving, ito2013alife].
-Similarly, morphologies that evolved in complex, multi-objective environments transfer more effectively to novel tasks than those from simpler settings [gupta2021embodied].
-Moreover, multi-task scenarios in ECD introduce additional brain–body trade-offs, where improvements on one task may degrade performance on others [wang2023preco].
-As a result, multi-task co-design aims to cultivate broadly capable brain–body pairs by leveraging shared structures across tasks rather than optimizing each task independently.
+Prior studies indicate that robots co-designed on uneven terrain tend to develop more stable, balanced, and limb-rich morphologies, whereas flat terrains favor simpler forms [miras2020environmental, miras2019effects, spanellis2021environment, corucci2018evolving].
+Moreover, morphologies designed in complex, multi-task environments exhibit stronger transfer to novel tasks than those developed in simpler settings [gupta2021embodied, wang2023preco].
+Note that multi-task scenarios introduce additional brain–body trade-offs, as morphological adaptations that improve one task may degrade performance on others.
+Consequently, multi-task ECD seeks to cultivate broadly capable brain–body pairs by leveraging the shared structures across tasks.
 III-D Co-Design Constraints
-Constraints are intrinsic to the design space of embodied agents.
-Unconstrained ECD often leads to overfitting: the morphological search may yield designs that are unnecessarily complex for a single task or, conversely, result in trivial solutions with poor transferability.
+Despite their significance, many ECD approaches still operate in largely unconstrained settings, which can lead to overfitting.
+Specifically, morphological search may yield designs that are unnecessarily complex for a single task or, conversely, result in trivial solutions with poor transferability.
 Therefore, principled regularization is essential for guiding the search toward balanced brain-body pairs.
-Morphological constraints, such as structural stability, material availability, joint range-of-motion, and actuation limits, need to be considered during optimization.
-Common strategies include enforcing symmetry priors [dong2023symmetry, gupta2021embodied] and incorporating penalties for manufacturing time, cost, and complexity into objective functions, as demonstrated in monotone co-design [censi2015mathematical, zardini2021co, wilhelm2023constraint].
-Behavioral constraints, such as safety considerations, actuator saturation, and collision avoidance, can further limit the feasible solution space and hinder the exploitation of simulators.
-Despite their significance, many ECD approaches still operate in largely unconstrained settings.
-Integrating practical engineering and fabrication constraints remains a challenge for the deployment of real-world applications.
+Morphological constraints, such as structural stability, material availability, joint range-of-motion, and actuation limits, can be explicitly incorporated into the design process.
+Common strategies include enforcing symmetry priors [dong2023symmetry, gupta2021embodied] and penalizing manufacturing time, cost, and complexity in the objective function, as in monotone co-design [censi2015mathematical, zardini2021co, wilhelm2023constraint].
+Behavioral constraints, such as safety, actuator saturation, and collision avoidance, can further limit the feasible behavior space and hinder the overuse of simulators.
 III-E Co-Design Frameworks
-Based on the fundamental elements mentioned above, we further categorize ECD approaches into four frameworks: (1) bi-level co-design, (2) single-level co-design, (3) generative co-design, and (4) open-ended co-design.
-These categories reflect how algorithms optimize morphology, control, and, in some cases, the task environment.
-In Section IV, we elaborate on these frameworks and analyze the existing literature through the lens of our new taxonomy.
+Based on the fundamental elements mentioned above, we categorize ECD approaches into four frameworks—(1) bi-level co-design, (2) single-level co-design, (3) generative co-design, and (4) open-ended co-design—because these categories capture the most significant differences in how algorithms coordinate morphology, control, and, optionally, the task environment during optimization (Section IV).
+The bi-level/single-level split distinguishes between hierarchical (nested) and unified (joint) optimization loops, thereby separating methods based on evaluation cost, convergence behavior, and morphological diversity.
+The generative category isolates approaches that rely on learned or rule-based priors to generate candidate bodies, emphasizing the trade-offs between sample efficiency and design freedom.
+Open-ended co-design, finally, groups methods that promote continual innovation through the co-evolution of task environments and agents.
+Together, these axes establish a design space for ECD methods that can guide the development of new embodied agents and their customization for specific applications.
 III-F Evaluation Metrics
@@
 III-F1 Task Performance
-Task performance remains the primary application-level metric, directly measuring goal achievement.
-Evaluating this aspect is expensive during morphology search, because it typically requires complete policy optimization and extensive rollout-based testing.
+Task performance remains the primary application-level metric, as it directly measures goal achievement.
+Evaluating this aspect is costly during morphology search because it typically requires complete policy optimization and extensive rollout-based testing.
 III-F2 Morphological Computation
@@
 which measures how informative the previous control state $s^{c}_{t}$ is about the next state $s^{c}_{t+1}$ when the action $a^{c}_{t}$ is known.
-A large MC indicates that morphology and environment strongly drive the state transition.
-While recent studies have directly quantified and optimized MC in physical robots [chandiramani2024improving, ghazi2016evaluating], practical applications often rely on simplified or surrogate estimators [zahedi2013quantifying, ghazi2015quantifying, fuchslin2013morphological, muller2017morphological].
+A large MC indicates that morphology and environment strongly drive state transitions.
+Although recent studies have directly quantified and optimized MC in physical robots [chandiramani2024improving, ghazi2016evaluating], most practical applications rely on simplified or surrogate estimators [zahedi2013quantifying, muller2017morphological].
 III-F3 Empowerment
-Similar to MC, empowerment measures the maximum influence an agent can exert on future states through its actions [klyubin2005empowerment, de2018unified, leibfried2019unified, jung2011empowerment], and is widely used in reinforcement learning exploration [houthooft2016vime, kumar2018empowerment, choi2021variational].
+Similar to MC, empowerment measures the maximum influence an agent can exert on future states through its actions [klyubin2005empowerment, leibfried2019unified], and it is widely used in reinforcement learning exploration [houthooft2016vime].
 Formally, the empowerment value for a control state $s^{c}_{t}$ is defined as the channel capacity between the action $a^{c}_{t}$ and the next state $s^{c}_{t+1}$:
@@
 In ECD, high empowerment indicates that a morphology can effectively translate diverse control signals into predictable outcomes, making downstream control easier [iii2021taskagnostic].
-III-F4 Quality-Diversity (QD) Score
-As discussed in Section IV-B4, diversity-based ECD approaches focus on discovering a wide repertoire of high-performing morphologies and behaviors.
-The QD score quantifies coverage and performance within a discretized Morphological Descriptor (MD) space, where each cell represents a morphology niche.
+III-F4 Quality-Diversity Score
+As discussed in Section IV-B4, diversity-based ECD methods aim to discover a wide repertoire of high-performing morphologies and behaviors.
+Accordingly, the Quality-Diversity (QD) score measures both coverage and performance over a discretized Morphological Descriptor (MD) space, where each cell corresponds to a distinct morphological niche.
 For an archive with $N$ cells, the QD score is computed as the sum of the task returns $\mathcal{R}_{i}$ of the best individual in each occupied cell:
@@
 III-F5 Complexity
-Complexity measures the cost of both the controller and the morphology.
-Controller complexity can be quantified using architecture-agnostic metrics such as FLOPs, which estimate computational cost, latency, and energy consumption [xie2025morphology].
-Morphological complexity can be quantified through component counts (e.g., links, joints, voxels, and degrees of freedom) or structural descriptors (e.g., connectivity, actuator distribution, material heterogeneity, and symmetry).
+Complexity captures the cost associated with both the controller and the morphology.
+Controller complexity can be measured using architecture-agnostic metrics, such as FLOPs, which estimate computational load, latency, and energy consumption [xie2025morphology].
+Morphological complexity can be quantified via component counts (e.g., links, joints, degrees of freedom) or structural descriptors (e.g., connectivity, material heterogeneity, and symmetry).
 III-F6 Robustness
-Robustness evaluates an agent’s resilience to perturbations, including morphological damage (e.g., locked joints, sensor faults) and environmental variations.
-Robust agents maintain high performance and enable rapid adaptation, reflecting their inherent fault tolerance [powers2021good].
+Robustness evaluates an agent’s resilience to perturbations, including morphological damage (e.g., locked joints, sensor faults) and environmental variations [powers2021good].
+Robust agents maintain high performance and enable rapid adaptation, reflecting their inherent fault tolerance.
 In summary, these metrics provide a multidimensional evaluation framework for ECD approaches, helping researchers enhance their understanding of how physical structures and controllers computationally contribute to behavior.
@@
 At the bottom of Fig.
-4, we use BipedalWalker agent  [^1] https://gymnasium.farama.org/environments/box2d/bipedal_walker/ in the OpenAI Gymnasium platform as a working example to summarize the common pipeline of ECD.
+3, we use BipedalWalker agent  [^1] https://gymnasium.farama.org/environments/box2d/bipedal_walker/ in the OpenAI Gymnasium platform as a working example to summarize the common pipeline of ECD.
   • Controlling Brain: The controller is a learnable multilayer perceptron (MLP) that receives joint angles, velocities, and ground-contact signals as inputs and outputs joint torques (individual representation), forming a standard closed-loop controller architecture.
@@
 IV Methods: Synthesizing Embodied Co-Design
-Figure 4: An overview of four embodied co-design frameworks.
+Fig.
+3: An overview of four embodied co-design frameworks.
 Example images of co-designed agents from (a) to (e) are adapted from DERL [gupta2021embodied], Softzoo [wang2023softzoo], EvolutionGym [bhatia2021evolution], RoboGrammar [zhao2020robogrammar], and [ha2019reinforcement].
 In this section, we elaborate on four ECD frameworks and organize them into sub-categories following the taxonomy outlined in Section III.
-For each sub-category, we analyze its core methodological components and highlight recent approaches.
+For each sub-category, we analyze its core methodological components and highlight recent advances.
 An overview of these frameworks is presented in Fig.
-4.
+3.
 IV-A Bi-Level Co-Design
@@
 IV-A1 Evolutionary Reinforcement Learning
-In this method, the inner loop for control policy optimization typically employs reinforcement learning (RL) [singh2022reinforcement, wang2022deep], whereas the outer loop for morphology evolution relies on various evolutionary algorithms (EAs).
+In this method, the inner loop for control policy optimization typically employs reinforcement learning (RL) [singh2022reinforcement, wang2022deep], while the outer loop for morphology evolution utilizes various evolutionary algorithms (EAs).
 Two major classes of RL methods are used for policy optimization: model-based and model-free approaches.
-Model-free RL learns a policy directly from interactions, without explicitly modeling transition dynamics or reward functions.
-Representative algorithms include proximal policy optimization (PPO) [schulman2017proximal], trust region policy optimization (TRPO) [schulman2015trust], deep deterministic policy gradient (DDPG) [silver2014deterministic, lillicrap2015continuous], twin delayed DDPG [fujimoto2018addressing], and soft actor-critic (SAC) [haarnoja2018soft].
-Model-based RL, in contrast, leverages a provided or learned dynamics model to refine the control policy.
+Model-free RL learns a policy directly from interactions without explicitly modeling transition dynamics.
+Representative algorithms include proximal policy optimization (PPO) [schulman2017proximal], trust region policy optimization (TRPO) [schulman2015trust], deep deterministic policy gradient (DDPG) [lillicrap2015continuous], twin delayed DDPG [fujimoto2018addressing], and soft actor-critic (SAC) [haarnoja2018soft].
+In contrast, Model-based RL leverages either a provided or learned dynamics model to refine the control policy.
 Representative frameworks include Mb-Mf [nagabandi2018neural], PETS [chua2018deep], GPS [levine2013guided], and PILCO [deisenroth2011pilco].
-Model-based approaches are more sample-efficient but can struggle in complex domains where accurate dynamics modeling is harder than direct policy learning.
+Model-based approaches are more sample-efficient but may struggle in complex domains where accurate dynamics modeling is harder than direct policy learning.
 EAs are effective for exploring high-dimensional morphology spaces.
 In the outer loop, evolutionary operators determine how new morphologies are generated, thereby shaping both the breadth and depth of exploration [carvalho2024role].
-In discrete design spaces, the mutation operator typically adds or removes structural elements—e.g., joints, limbs, modules, kinematic branches—or modifies material properties.
+In discrete design spaces, the mutation operator typically adds or removes structural elements—e.g., joints, limbs, modules, and kinematic branches—or modifies material properties.
 In continuous design spaces—including link lengths, joint stiffness, and mass distribution—Gaussian perturbations are commonly used.
@@
 Notably, generative models, such as Large Language Models (LLMs), have been used as high-level mutation operators to propose structured, human-interpretable edits based on prompts [song2025laser], further expanding the operator toolbox for morphology search.
-By alternating between individual-level controller learning and population-level morphology evolution, ERL methods mirror Darwinian evolution augmented with a Baldwinian learning mechanism [luo2022effects, li2023evaluation].
-A representative example is Neural Graph Evolution (NGE) [wang2019neural], which evolves graph-structured, rigid-body morphologies while training universal GNN-based controllers via PPO.
-NGE demonstrates that GNN controllers can generalize across related body structures, thereby reducing the cost of controller retraining during morphology search.
+By alternating between individual-level controller learning and population-level morphology evolution, ERL methods reflect Darwinian evolution, enhanced by a Baldwinian learning mechanism [luo2022effects, li2023evaluation].
+A representative example of this line is Neural Graph Evolution (NGE) [wang2019neural], which evolves graph-structured morphologies while training universal GNN-based controllers via PPO.
+NGE demonstrates that GNN controllers can generalize across various body structures, thereby reducing the cost of controller retraining during morphology search.
 The scalability of ERL is further illustrated by Deep Evolutionary Reinforcement Learning (DERL) [gupta2021embodied].
 In contrast to NGE, DERL employs individual controllers that are trained from scratch for each morphology.
-Coupled with increasingly challenging environments—ranging from flat-ground locomotion to uneven terrain and object manipulation—this framework enables the emergence of morphological intelligence, resulting in agents that exhibit increased robustness, adaptability, and behavioral diversity.
-Similar strategies appear in later works [park2021computational, bhatia2021evolution, chen2023evolving, wang2024learn, yue2025toward].
-As summarized in Table I, recent research has also expanded the classical ERL workflow along several dimensions.
-First, richer morphological representations have been introduced, including hyperbolic embeddings [dongleveraging2024] and text-to-mesh generation [ringel2025text2robot].
-Second, advanced control-learning techniques, such as meta-RL [belmonte2022meta], transfer learning [harada2024lamarckian], imitation learning [hernandezfollowing], and policy-reuse [xie2025accelerating], have improved sample efficiency and generalization.
-A parallel line of work introduces novel morphological selection mechanisms, such as successive halving [katayama2023ga], multi-fidelity archiving [zhao2024morphological, nagiredla2024ecode], and multi-objective balancing [zhao2025multi], which can prevent premature convergence in morphology space [doncieux2014beyond, mertan2024investigating].
+Coupled with increasingly challenging environments, this framework enables the emergence of morphological intelligence, resulting in agents that exhibit increased robustness, adaptability, and behavioral diversity.
+Similar strategies appear in subsequent works [park2021computational, bhatia2021evolution, chen2023evolving, wang2024learn, yue2025toward, cheng2024structural, harada2024lamarckian, ringel2025text2robot, jin2025evolutionary].
+As summarized in Table I, recent research has expanded the classical ERL workflow along several dimensions.
+First, richer morphological representations have been introduced, including hyperbolic embeddings [dongleveraging2024], text-to-mesh generation [ringel2025text2robot], and latent space embedding [jeon2025convergent].
+Second, advanced control-learning techniques, such as meta-RL [belmonte2022meta], transfer learning [harada2024lamarckian], imitation learning [hernandezfollowing], and policy-sharing [xie2025accelerating, li2024generating, wang2024evolution], have improved sample efficiency and generalization.
+A parallel line of work introduces novel morphological selection mechanisms, such as paired competition [banarse2019body], successive halving [katayama2023ga], multi-fidelity archiving [zhao2024morphological, nagiredla2024ecode], and multi-objective balancing [zhao2025multi], which can prevent premature convergence in morphology space [doncieux2014beyond, mertan2024investigating].
 Notably, the combination of EA and RL has also been widely explored in control optimization.
-These hybrid methods maintain a population of policies that evolve across generations, enabling broad exploration of the policy space while refining individual controllers via RL’s gradient-based or value-based updates.
+These hybrid methods maintain a population of policies that evolve across generations, allowing for broad exploration of the policy space while refining individual controllers via RL’s gradient-based or value-based updates.
 Representative approaches include ERL [khadka2018evolution], CERL [khadka2019collaborative], PDERL [bodnar2020proximal], SERL [wang2022surrogate], and ERL-Re2 [hao2023erl].
 Comprehensive reviews of this line of work can be found in [drugan2019reinforcement, sigaud2023combining, bai2023evolutionary, zhu2023survey, lin2025evolutionary, li2024bridging].
-IV-A2 Surrogate-Assisted Methods
-A central challenge in bi-level co-design arises from its hierarchical structure, where the inner-loop optimization is computationally intensive and often requires days of training for complex tasks.
-Surrogate-assisted methods address this issue by employing predictive models to approximate task performance, thereby reducing the need for repeated full-scale policy evaluations.
-Bayesian Optimization (BO) [shahriari2015taking] has emerged as a prominent framework for modeling the morphology–performance relationship in ECD.
-As a black-box, zeroth-order optimizer, BO efficiently handles expensive objective evaluations.
-At each iteration, BO updates a surrogate model using previously evaluated morphologies and their performance, then selects new candidates through acquisition functions such as Probability of Improvement (PI), Expected Improvement (EI), Entropy Search (ES), or Upper Confidence Bound (UCB), thereby balancing exploration and exploitation [wilson2018maximizing].
-The selected candidates are evaluated, and their results are used to update the surrogate for the subsequent iteration.
-Gaussian Processes (GPs) are among the most commonly used surrogates in BO.
-For instance, HPC-BBO [liao2019data] introduced a hierarchical framework that facilitates parallel hardware validation.
-Similarly, [bjelonic2023learning] pre-trained a universal controller in simulation and used BO to optimize parallel–elastic spring parameters for a real quadruped, demonstrating effective sim-to-real transfer.
-Variants of BO have been further explored in [bhatia2021evolution, hu2021neural, saito2024effective, pan2021emergent].
-To enhance generalization, [chen2024pretraining] combined domain randomization with a pretraining–finetuning pipeline.
-Extending beyond single-task BO, [dong2024cagecoopt] proposed a multi-task BO framework for robot manipulators.
-More recently, [schneider2025task] combined BO with Hyperband to generate novel manipulator designs that improve performance across both seen and unseen tasks.
-In addition to traditional surrogates used in BO, recent approaches have explored alternative choices.
-For example, Q-value functions (Q-Surro) from RL have been used to evaluate candidate morphologies [luck2020data, luck2021robot, du2025efficient].
-Multi-dynamics models can further enhance generalization across different agent designs [sliacka2023co].
-Alternatively, when historical interaction data is available, near-optimal control actions (A-Surro) can be partially or fully transferred from previously optimized designs to accelerate evaluation [liu2023rapidly].
-More recent studies by [zhao2025cross, strgar2025accelerated] developed universal policies that span multiple tasks and can be efficiently used as a zero-shot performance evaluator (A-Surro), thereby improving efficiency.
-In [bai2025learning], manipulator morphologies are evaluated and optimized using pre-collected trajectories and a learned reward model (R-Model), with the reliance on reward model evaluations gradually increasing throughout the training process.
 Table I: Summary of bi-level co-design approaches.
@@
 {MuJoCo}
-Learnable (CL) Universal Discrete (UF) Direct Predefined (Partial) Task-Driven
+Learnable (CL) Universal Hybrid (UF) Direct Predefined (Partial) Task-Driven
 RL (PPO) GNN EA (Mutation) Graph Locomotion Task Reward
@@
 {MuJoCo}
-Learnable (CL) Individual Discrete (UF) Direct Modifiable (Full) Task-Driven
+Learnable (CL) Individual Hybrid (UF) Direct Modifiable (Full) Task-Driven
 RL (PPO) DNN EA (Mutation) Tree Loco.&Manip. Task Reward
@@
 RL (SAC) DNN EA Toolbox Vector Fall Recovery Task Reward
+LOKI [jeon2025convergent]
+2025
+[link]
+Legged
+Robot
+Simulation
+{MuJoCo}
+Learnable (CL) Universal Hybrid (UF) Indirect Predefined (Partial) Hybrid
+RL (PPO) Transformer EA (Search) Latent Locomotion Task+Diversity
 Surrogate-Assisted Methods
@@
 RL (N2M2) DNN BO&HyperBand Vector Manipulation Task Reward
-Table II: Continued from previous page.
-Method Year Agent Type Application (Sim/Real) Controlling Brain Body Morphology Task Environment
-Design Space Representation Design Space Representation Design Space Representation
 Du et al. [du2025efficient]
@@
 Grad. Opt. DNN&CPG EA (A-Surro) Matrix Locomotion Task Reward
+Table II: Continued from previous page.
+Method Year Agent Type Application (Sim/Real) Controlling Brain Body Morphology Task Environment
+Design Space Representation Design Space Representation Design Space Representation
 Bai et. al. [bai2025learning]
@@
 Opt.: Gradient-Based Optimization, Nonlinear-Opt: Nonlinear Optimization, DDP: Differential Dynamic Programming, LQR: Linear Quadratic Regulators, AME: Asynchronous Morpho-Evolution,Hom.: Homeokinesis
+IV-A2 Surrogate-Assisted Methods
+A central challenge in bi-level co-design arises from its hierarchical structure, where the inner-loop optimization is computationally expensive and often requires days of training for complex tasks.
+Surrogate-assisted methods address this issue by employing predictive models to approximate agent task performance, thereby reducing the need for repeated full-scale policy evaluations.
+Bayesian Optimization (BO) [shahriari2015taking] has emerged as a prominent framework for modeling the morphology–performance relationship in ECD.
+As a black-box, zeroth-order optimizer, BO efficiently handles expensive objective evaluations.
+At each iteration, BO updates a surrogate model using previously evaluated morphologies and their performance, then selects new candidates through acquisition functions, such as Probability of Improvement (PI), Expected Improvement (EI), Entropy Search (ES), or Upper Confidence Bound (UCB), thereby balancing exploration and exploitation [wilson2018maximizing].
+The selected candidates are evaluated, and their results are used to update the surrogate model for the subsequent iteration.
+Gaussian Processes (GPs) are among the most commonly used surrogates in BO.
+For instance, HPC-BBO [liao2019data] introduced a hierarchical framework that facilitates parallel hardware validation.
+Similarly, [bjelonic2023learning] pre-trained a universal controller in simulation and used BO to optimize parallel–elastic spring parameters for a real quadruped, demonstrating effective sim-to-real transfer.
+Variants of BO with GPs have been further explored in [bhatia2021evolution, hu2021neural, saito2024effective, pan2021emergent, rajani2023co].
+To enhance generalization, [chen2024pretraining] combined domain randomization with a pretraining–finetuning pipeline.
+Extending beyond single-task BO, [dong2024cagecoopt] proposed a multi-task BO framework for robot manipulators.
+More recently, [schneider2025task] combined BO with Hyperband to generate novel manipulator designs that improve performance across both seen and unseen tasks.
+In addition to traditional surrogates used in BO, recent approaches have explored alternative choices.
+For example, Q-value functions (Q-Surro) from RL have been used to evaluate candidate morphologies [luck2020data, luck2021robot, du2025efficient].
+Multi-dynamics models can further enhance generalization across different agent designs [sliacka2023co].
+Alternatively, when historical interaction data is available, near-optimal control actions (A-Surro) can be partially or fully transferred from previously optimized designs to accelerate evaluation [liu2023rapidly].
+More recent studies by [zhao2025cross, strgar2025accelerated] developed universal policies that span multiple tasks and can be efficiently employed as a zero-shot performance evaluator (A-Surro), thereby improving efficiency.
+Moreover, in [bai2025learning], manipulator morphologies are evaluated and optimized using pre-collected trajectories and a learned reward model (R-Model), with the reliance on reward model evaluations gradually increasing throughout the training process.
 IV-A3 Other Methods
-The inner-loop controller optimization can be implemented using alternatives to RL, including trajectory optimization [fadini2022simulation], evolutionary algorithms [jelisavcic2019lamarckian, le2022morpho, le2025efficient], numerical solvers [sartore2023codesign, ha2018computational2], and information-theoretic classifiers [iii2021taskagnostic].
+Notably, the inner-loop controller optimization can be implemented using alternatives to RL, including trajectory optimization [fadini2022simulation], evolutionary algorithms [jelisavcic2019lamarckian, le2022morpho, le2025efficient], numerical solvers [sartore2023codesign, ha2018computational2], and information-theoretic classifiers [iii2021taskagnostic].
 To accelerate bi-level co-design, Goff et al.
-proposed A-MEL [goff2024efficient], which integrates asynchronous learning and evolution to reduce both time and computational costs.
+proposed A-MEL [le2024improving], which integrates asynchronous learning and evolution to reduce both time and computational costs.
 Strgar et al.
-[Strgar-RSS-24] further leveraged massively parallel differentiable simulations to efficiently optimize neural controllers for large populations of candidate morphologies.
+[Strgar-RSS-24] further leveraged massively parallel differentiable simulations to efficiently optimize neural controllers for large morphology populations.
 IV-B Single-Level Co-Design
 As shown in Fig.
-4 (b), single-level co-design treats both morphology and control as variables within a unified optimization process, unlike bi-level co-design, which employs nested loops for these aspects.
-We organize recent work along this line into four subcategories: 1) EA-Based Methods, 2) RL-Based Methods, 3) Physics-Based Methods, and 4) Diversity-Based Methods, which are summarized in Table III.
+3 (b), single-level co-design treats both morphology and control as variables within a unified optimization process.
+We categorize recent work along this line into four subcategories: 1) EA-Based Methods, 2) RL-Based Methods, 3) Physics-Based Methods, and 4) Diversity-Based Methods, which are summarized in Table III.
 IV-B1 EA-Based Methods
 EAs have a long-standing history of jointly evolving robot morphology and control.
-In these methods, each individual in the population encodes a complete agent, encompassing both the body structure and the controller parameters.
+In these methods, each individual in the population encodes either a complete agent or only its morphology, along with a simplified controller.
 The evolutionary loop evaluates fitness and applies operators such as selection, mutation, and crossover to explore the combined design space.
@@
 Morphological innovation protection [cheney2018scalable] mitigates premature convergence by preserving newly evolved traits, while other studies incorporate heuristic search [ha2018computational], biologically inspired strategies [davis2023subtract], and neuroscience-inspired approaches [ferigo2025totipotent].
-Multi-objective EAs, such as NSGA-II [deb2002fast], have been utilized to balance competing design objectives and constraints.
+Moreover, multi-objective EAs, such as NSGA-II [deb2002fast], have been utilized to balance competing design objectives and constraints.
 For example, minimizing energy consumption and mission time jointly in morphing aerial vehicles [bergonti2024co].
-EA-based methods excel at discovering creative, unexpected morphologies due to their global and relatively unconstrained search.
-In soft robotics, they have been widely used to automatically generate complex geometries and behaviors [cheney2013unshackling, hiller2010evolving, cheney2014electro, hiller2012automatic, kriegman2021fractals].
+EA-based methods excel at discovering novel morphologies due to their global and relatively unconstrained search capabilities.
+These methods have been widely used to generate complex geometries and behaviors [cheney2013unshackling, hiller2010evolving, cheney2014electro, hiller2012automatic, kriegman2021fractals, sathuluri2023robust].
 However, scalability remains a major limitation: population-based evaluation is computationally expensive—often requiring thousands of simulations per generation—and convergence is typically slow, mirroring the efficiency bottlenecks observed in bi-level co-design.
@@
 Early methods incorporated morphology directly into the policy representation.
-For example, [ha2019reinforcement] used a population-based policy-gradient method (REINFORCE) to update control and morphology simultaneously.
-Similarly, [chen2018hardware] embedded body parameters into the computational graph of an RL policy, enabling joint optimization of physical parameters and neural network weights.
+For example, [ha2019reinforcement] used a population-based policy-gradient method (REINFORCE) to update both control and morphology simultaneously.
+Similarly, [Chen2020HardwareAP] embedded body parameters into the computational graph of a RL policy.
 [schaff2019jointly] co-optimized a differentiable distribution over morphology parameters alongside a universal controller; this framework was later extended by [schaff2022n] for more complex embodiments.
-Other works, such as [whitman2020modular2], employed Q-learning to guide modular manipulator design, while [luck2020data, luck2021robot] combined RL with a learned Q-value surrogate to predict the agent’s task performance.
-A complementary paradigm treats design modifications themselves as actions within the RL framework.
+Other works, such as [whitman2020modular2], employed Q-learning to guide modular manipulator design.
+A complementary paradigm treats design modifications as actions within the RL framework.
 Under this formulation, the agent concurrently learns to control and perform morphological “design actions”.
-A representative work is Transform2Act [yuan2022transformact], which introduces a transform-and-control paradigm that formulates design optimization as a conditional policy over both control and morphology, with policy-gradient updates and experience sharing across designs for improved sample efficiency.
-Subsequent extensions incorporated curriculum learning [wang2023curriculum], multi-task pretraining [wang2023preco], symmetry priors [dong2023symmetry], hierarchical modeling [sun2023co], competition mechanisms [huang2024competevo], and temporal credit assignment [lu2025bodygen], with applications ranging from embodied manipulation [liu2023learning, kulz2025design] to freeform modular robot design [li2024reinforcement].
-Moreover, real-world considerations have also motivated sim-to-real approaches.
+A representative work is Transform2Act [yuan2022transformact], which introduces a transform-and-control paradigm that formulates design optimization as a conditional policy over both control and morphology.
+This approach incorporates policy-gradient updates and experience sharing across designs to improve sample efficiency.
+Subsequent extensions have incorporated curriculum learning [wang2023curriculum], multi-task pretraining [wang2023preco], symmetry priors [dong2023symmetry], hierarchical modeling [sun2023co], competition mechanisms [huang2024competevo], and temporal credit assignment [lu2025bodygen], with applications ranging from embodied manipulation [liu2023learning, kulz2025design] to freeform modular robot design [li2024reinforcement].
+Moreover, real-world considerations have motivated sim-to-real approaches.
 [schaff2022soft] showed that co-designed soft robots can effectively navigate complex terrains.
@@
 If the simulator, analytic model, or learned dynamics is differentiable, gradients can be computed with respect to both the controller and body parameters.
-Task losses can then be backpropagated through time using standard optimizers (Grad.
+Task losses can then be backpropagated through time using gradient-based optimizers (Grad.
 Opt.), such as ADAM [adam] or L-BFGS [zhu1997algorithm].
@@
 Early efforts relied on parametric analytic models [spielberg2017functional, ha2018computational] and adjoint methods [desai2018interactive] to model the complex relationships between morphology and control parameters.
-Subsequent research introduced stochastic programming combined with trajectory optimization [bravo2020one, bravo2024engineering], improving scalability across tasks and environments.
+Subsequent research introduced stochastic programming in combination with trajectory optimization [bravo2020one, bravo2024engineering], thereby improving scalability across tasks and environments.
 More recent methods incorporate hardware and environmental constraints [ghansah2023humanoid, vaish2024co, sartore2022optimization].
@@
 These processes enable gradients to propagate through more expressive design spaces.
-Built on this foundation, topology-agnostic parameterizations, such as deformation-based [xu2021end, cai2023task2morph, liu2023learning, lirobottool2023] and density-based schemes [yuhn20234d, matthews2023efficient, sato2023topology, sato2025computational], have been proposed for agents with complex geometries.
+Built on this foundation, deformation-based [xu2021end, cai2023task2morph, liu2023learning, lirobottool2023] and density-based schemes [yuhn20234d, matthews2023efficient, sato2023topology, sato2025computational] have been proposed for agents with complex geometries.
 Additional mechanisms include particle sets [spielberg2021co, matthews2023efficient], Wasserstein barycenters [ma2021diffaqua, wang2023softzoo], and CPPN-based encodings [cochevelou2023differentiable].
@@
 Grad. Opt. Torques Grad. Opt. Vector Payload Lifting Task Loss
-Table IV: Continued from previous page.
-Method Year Agent Type Application (Sim/Real) Controlling Brain Body Morphology Task Environment
-Design Space Representation Design Space Representation Design Space Representation
 Yuhn et al. [yuhn20234d]
@@
 Grad. Opt. Torques Grad. Opt. Vector Manipulation Task Loss
+Table IV: Continued from previous page.
+Method Year Agent Type Application (Sim/Real) Controlling Brain Body Morphology Task Environment
+Design Space Representation Design Space Representation Design Space Representation
 Cochevelou et al. [cochevelou2023differentiable]
@@
 Diversity-Based Methods
-Krvcah et al. [krvcah2010solving]
-2010
-Virtual
-Creature
-Simulation
-{ODE}
-Learnable (CL) Individual Hybrid (UF) Indirect Predefined (Partial) Hybrid
-EA (NEAT) ANN EA (NEAT) Nested Graph Loco.&Navi. Task+Novelty
-Lehman et al. [lehman2011evolving]
-2011
-Virtual
-Creature
-Simulation
-{ODE}
-Learnable (CL) Individual Hybrid (UF) Indirect Predefined (Partial) Hybrid
-EA (NSGA‑II) ANN EA (NSGA‑II) Nested Graph Loco.&Navi. Task+Novelty
 Nordmoen et al. [nordmoen2020quality]
@@
 IV-B4 Diversity-Based Methods
-A persistent challenge in single-level co-design is the premature convergence of both morphology and controller.
-Diversity-based methods address this issue by optimizing a repertoire of distinct, high-quality body–brain pairs instead of focusing on a single global optimum.
-Early work [krvcah2010solving, lehman2011evolving] demonstrated this principle by employing Novelty Search [doncieux2019novelty] with local competition within a three-dimensional morphology space (height, mass, and joint count), showing that localized selection pressures help preserve morphological diversity.
-The subsequent line built upon Quality-Diversity (QD) algorithms [pugh2016quality, cully2017quality].
-Canonical QD methods target behavioral diversity [batra2024quality, nilsson2021policy, flageat2024enhancing, pierrot2022multi, faldor2023map, pierrot2022diversity].
-For example, MAP-Elites [mouret2015illuminating] discretizes a user-defined behavior descriptor (BD) space into cells and stores the best-performing solution within each.
+A persistent challenge in single-level ECD is the premature convergence of both morphology and controller.
+Diversity-based methods address this issue by optimizing a repertoire of high-quality body–brain pairs, rather than concentrating on a single global optimum.
+Early work by [krvcah2010solving, lehman2011evolving] demonstrated this principle by employing Novelty Search [doncieux2019novelty] with local competition in a three-dimensional morphology space (height, mass, and joint count), showing that localized selection pressures help preserve morphological diversity.
+The subsequent work built upon Quality-Diversity (QD) algorithms [pugh2016quality, cully2017quality].
+Canonical QD methods target only behavioral diversity.
+For example, MAP-Elites [mouret2015illuminating] discretizes a user-defined behavior descriptor (BD) space into cells and stores the best-performing solution within each cell.
 In the context of ECD, morphology descriptors (MDs) are introduced to index MAP-Elites archives.
@@
 Recent works extend this framework to legged robots and manipulators [howard2020diversity, xie2024map, norstein2023effects], or distill the archive into a universal controller to stabilize the co-design process [mertan2025controller].
-Moreover, dual-map approaches separately represent morphology and controller spaces [zardini2021seeking], whereas 3B-QD [nadizar2025enhancing] maintains three independent archives for morphology, control, and behavior, further enhancing the illumination of the agent co-design space.
-Diversity-based methods accelerate convergence by avoiding repeated controller optimization for each new morphology.
-Remaining challenges include the automatic discovery of informative descriptors, scaling to high-resolution material-property spaces [pigozzi2023factors], and providing theoretical guaranties for niche-optimal solutions.
-Comparative studies of QD approaches in ECD are available in [mkhatshwa2023impact, mkhatshwa2025body].
+Moreover, dual-map approaches separately represent morphology and controller spaces [zardini2021seeking], while 3B-QD [nadizar2025enhancing] maintains three independent archives for morphology, control, and behavior, further enhancing the illumination of the agent co-design space.
+Despite their effectiveness, diversity-based ECD faces several challenges, including the automatic discovery of informative descriptors, scalability to high-dimensional material and property spaces [pigozzi2023factors, mertan2024no], and the lack of theoretical guarantees for niche-optimal solutions.
+Comparative evaluations of QD-based approaches in ECD can be found in [mkhatshwa2023impact, mkhatshwa2025body].
 IV-C Generative Co-Design
-Generative co-design leverages automated generative models to capture the relationship among morphology, task, and performance (evaluated by learning a controller), thereby enabling the optimization of morphology within latent or rule-defined spaces, as illustrated in Fig.
+Generative co-design leverages automated generative models to capture the relationships among morphology, tasks, and performance (evaluated by learning a controller).
+This approach enables optimization within latent or rule-defined spaces, as illustrated in Fig.
 3 (c).
-We further categorize the generative co-design framework into three topics: 1) Rule-Based Generation, 2) Latent-Based Generation, and 3) Large-Model-Based Generation.
-Each class employs a distinct generative representation and search strategy to explore morphological design spaces, as shown in Table V.
+We further categorize the generative co-design framework into three sub-topics: 1) Rule-Based Generation, 2) Latent-Based Generation, and 3) Large-Model-Based Generation.
+Each class employs a distinct generative representation and search strategy to explore morphology spaces, as summarized in Table V.
 IV-C1 Rule-Based Generation
@@
 More recent approaches combine grammars with advanced search strategies.
-Robogrammar [zhao2020robogrammar] and its extension [zhao2025multi], for example, encode robots as graphs, enforce valid assemblies through a graph grammar, and couple heuristic search with a model-based controller to efficiently identify high-performing designs.
+Robogrammar [zhao2020robogrammar] and its extension [xu2021multi] encode robots as graphs, enforce valid assemblies through a graph grammar, and couple heuristic search with a model-based controller to efficiently identify high-performing designs.
 Zhao et al.
@@
 [zharkov2024synergizing] further applied grammar-based linkage graphs to co-design underactuated, tendon-driven grippers.
-Rules effectively reduce the search space to plausible designs, explicitly encode prior knowledge such as symmetry and connectivity, and enhance manufacturability and modularity.
-While they are flexible and interpretable, these rules depend on expert-crafted grammars, which can introduce bias into the search process and limit the exploration of novel or fine-grained morphological variations.
-IV-C2 Latent-Based Generation
-Latent-based methods leverage neural generative models to learn smooth mappings from low-dimensional latent vectors to full 3D or graph-based morphologies.
-These learned latent spaces support efficient gradient-based, evolutionary, or probabilistic search over complex design spaces and are increasingly influential in ECD.
-Variational Autoencoders (VAEs) learn a continuous latent representation of morphology and, in some cases, control parameters, facilitating sample-efficient optimization.
-LABO [pan2021emergent] embeds high-dimensional hand morphologies and controller parameters into a VAE latent space and performs BO within that space, thereby reducing dimensionality and sample complexity for grasping tasks.
-GLSO [hu2023glso] extends this idea to graph-structured robots by training a graph-VAE on grammar-generated topologies and optimizing designs using BO.
-Liu et al.
-[liu2025terrain] proposed a Grammar-VAE that maps discrete morphologies into a continuous latent space and integrated differential evolution with a neural decoder, allowing the generative model to be reused across tasks.
-MorphVAE [song2024morphvae] further conditions latent sampling on task embeddings and introduces a “continuous natural selection” strategy to improve diversity and performance.
-Generative Adversarial Networks (GANs) train a generator–discriminator pair to produce realistic morphologies by matching target data distributions.
-RoboGAN [hu2022modular] learns a one-to-many mapping from task descriptions (e.g., terrain type) to distributions over modular legged robots.
-Candidate bodies are sampled from the GAN manifold, while an evolutionary loop guides the search toward high-performing configurations, thereby enabling interpolation and diversity within the learned latent space.
-Diffusion Models have recently emerged as powerful generative priors for embodied co-design.
-DiffuseBot [wang2023diffusebot] integrates differentiable physics into the denoising process, using simulation feedback at each step to bias generation toward high-performing designs.
-Dynamics-Guided Diffusion Models (DGDM) [xu2024dynamics] replace task-specific training data with physics-based guidance, enabling generalization to new manipulation tasks and novel geometries.
-GFlowNets provide an alternative generative formulation in which morphology–controller pairs are sampled with probability proportional to their simulated performance.
-Recent works by [nagiredla2024robonet, nagiredlate, nagiredlacogent] demonstrated that GFlowNets can produce diverse, high-quality designs while learning a structured generative policy over complete embodied agents.
-Latent generative models effectively capture complex, free-form geometries, exploit correlations among design variables, and typically offer significant gains in sample efficiency when compared to direct search methods.
-Once trained, a latent model can be reused across various tasks or paired with different controllers.
-However, training generative models typically requires large datasets and simulation efforts.
-Moreover, latent spaces may decode to invalid or unrealistic morphologies without appropriate constraints.
-Achieving the right balance between fidelity and diversity remains challenging.
-IV-C3 Large-Model-Based Generation
-Large pretrained models—including Large Language Models (LLMs) [naveed2025comprehensive] and Vision–Language Models (VLMs) [zhang2024vision]—have recently been introduced into ECD as generative priors or intelligent search operators.
-These models leverage broad knowledge and powerful generative capabilities to propose, mutate, or refine robot morphologies and control strategies.
-A growing line of work uses LLMs as intelligent mutation operators or search heuristics.
-OpenELM [lehman2023evolution] demonstrates that a LLM trained on code can act as a high-quality mutation operator in genetic programming.
-In Sodarace experiments, the LLM repeatedly mutated code that generates robot walkers, producing thousands of novel and functional agents without prior robotics data; the resulting dataset was then used to train a conditional generative model that maps terrain descriptions to robot code.
-Building on this paradigm, LASeR [song2025laser] integrates a LLM into the optimization loop via a reflection mechanism that summarizes past search histories and prompts the LLM to generate improved design variants.
-RoboMoRe [fang2025robomore] further investigates LLM-driven co-design by asking the LLM to jointly propose candidate morphology–reward pairs during a coarse search, after which promising candidates undergo alternating refinement.
-RoboMorph [qiu2024robomorph] represents each robot with a structured grammar and relies on LLM reasoning to navigate vast combinatorial spaces, iteratively improving morphologies through feedback.
-LLMs have also been utilized for human–AI collaborative design.
-Stella et al.
-[stella2023can] have shown that a LLM can provide high-level conceptual guidance and partial specifications for a tomato-harvesting gripper, illustrating its potential in early-stage design.
-Beyond text-based reasoning, vision models are beginning to appear in ECD.
-VLM-Engineer [gao2025vlmgineer] combines VLMs with evolutionary search to jointly design robotic tools and associated control policies.
-This line of work suggests future workflows in which a VLM first generates sketches or 2D concept designs from a task description, which are then lifted into full 3D morphologies.
-Large models offer several advantages for ECD: they encapsulate extensive engineering and commonsense knowledge, generate creative or unconventional designs, support semantic prompting (e.g., “design a robot capable of climbing a ladder”), and diversify the search space.
-Their in-context learning abilities make them powerful assistants for bridging high-level human intent with concrete design proposals [makatura2023can, ma2024llm].
-However, outputs from LLMs and VLMs are not guaranteed to be feasible, optimal, or physically consistent.
-Their behavior is often sensitive to prompt phrasing, and unguided LLM-driven searches can be highly sample-inefficient.
-Vision-based 3D generation is still in its infancy, and the reliance on proprietary foundation models raises additional concerns regarding reproducibility, transparency, and computational costs.
+Rule-based approaches effectively constrain the search space to plausible designs, explicitly encode prior knowledge (e.g., symmetry and connectivity), and improve manufacturability.
+However, their reliance on expert-crafted grammars can introduce bias and restrict the exploration of novel or fine-grained morphological variations.
 Table V: Summary of Generative Co-Design approaches.
@@
 MPPI MPC GHS Grammar Locomotion Task Reward
-RoboGrammar [xu2021multi]
+MOGHS [xu2021multi]
 2021
@@
 2025
-[link]
 Modular
@@
 Opt.: Gradient-Based Optimization, Manip.: Manipulation, MCTS: Monte‑Carlo Tree Search, GHS: Graph Heuristic Search, MPPI: Model Predictive Path Integral, DE: Differential Evolution, GP: Genetic Programming
-IV-D Open-Ended Co-Design
-Open-ended co-design aims for the continual or unbounded improvement of agents by enabling their bodies, controllers, and environments to co-evolve.
-These systems promote continuous innovation by enabling agents and environments to mutually generate new challenges and adaptations, as shown in Fig.
-4 (d).
-We categorize open-ended co-design into two primary subtopics: 1) Brain–Body–Environment Co-Evolution and 2) Developmental Evolution, as summarized in Table VI.
-IV-D1 Brain-Body-Environment Co-Evolution
-This approach treats the task environment as a co-optimized component alongside morphology and control.
-By evolving environments (or task curricula) together with agents, it creates continuously shifting challenges that prevent premature convergence and promote both behavioral and morphological improvement.
-The environment thus functions simultaneously as a generator of new selective pressures and as a filter for successful morphology–controller adaptations.
-Foundational work in this direction is the Paired Open-Ended Trailblazer (POET) framework [wang2019poet, wang2020enhanced].
-POET maintains a population of environment–controller pairs, simultaneously mutating environments and evolving agents.
-Policies that succeed in one environment may be transferred to other contexts.
-Stensby et al.
-[stensby2021co] extended POET to jointly evolve morphologies and controllers in tandem with their environments.
-Their results indicate that learned environment progressions yield greater morphological diversity than handcrafted curricula and reduce premature convergence [eguiarte2024premature].
-Recent work has introduced RL-driven co-evolution of the brain, body, and environment.
-Ao et al.
-proposed MECE [ao2023curriculum], which formulates a curriculum in which both morphology and the environment evolve according to learned RL policies.
-This produces progressively challenging and structured tasks that drive sustained adaptation.
-Another emerging trend augments environment evolution with generative models.
-For example, LLM-POET [aki2024llm] replaces manual environment mutation with a LLM that generates increasingly complex environments for a population of soft-robot agents.
-As agents improve, the LLM proposes novel tasks that escalate difficulty, realizing an open-ended loop of continual challenge and innovation.
-IV-D2 Developmental Evolution
-Developmental co-evolution draws inspiration from evolutionary developmental biology (Evo-Devo) and developmental robotics, where genotypes encode growth processes that shape both morphology and control over an agent’s lifetime (ontogeny) [asada2001cognitive, bongard2003evolving, lungarella2003developmental, kriegman2018morphological, van2021comparing, ivanchenko2003developmental].
-Open-endedness emerges from developmental plasticity [higashinaka2024evolution]: agents are evaluated across multiple developmental stages or allowed to modify their bodies during their lifetime, while evolution optimizes the developmental rules themselves.
-Early research has explored developmental encodings, including L-systems [rieffel2014growing], grammar rules, and neural networks [pontes2022single, joachimczak2016artificial].
-Bongard [bongard2003evolving] demonstrated that artificial ontogeny can generate hierarchical, repeated structures and controllers that emerge as bodies grow, yielding more structured solutions than direct genotype-phenotype encodings.
-The other research line focuses on morphological development.
-Here, developmental rules adjust parameters such as leg angles or actuator amplitudes during learning [bongard2011morphological, bongard2010utility, naya2020experiment], or progressively “unfreeze” degrees of freedom to facilitate skill acquisition [lungarella2002adaptivity, berthouze2004motor, savastano2012incremental].
-Later work introduced mechanisms for growing new body parts during an agent’s lifetime [arita2016alife, benureau2022morphological, jaafar2025eco], providing additional adaptive flexibility.
-For real-world deployment, virtual growth is typically halted once the simulated agent reaches the physical dimensions of the hardware robot, after which the learned controller is transferred to the physical platform [vujovic2017evolutionary, brodbeck2015morphological].
-For comprehensive overviews of developmental evolution, we refer readers to [naya2021morphological, wang2021survey].
-Overall, open-ended co-design is not yet a unified methodology but rather a collection of emerging ideas that leverage ontogeny, plasticity, and staged learning, and it awaits further exploration.
+IV-C2 Latent-Based Generation
+Latent-based methods leverage neural generative models to learn smooth mappings from low-dimensional latent vectors to full 3D or graph-based morphologies.
+These learned latent spaces support efficient gradient-based, evolutionary, or probabilistic search over complex design spaces and are increasingly influential in ECD.
+Variational Autoencoders (VAEs) learn continuous latent representations of morphology and, in some cases, control parameters, facilitating sample-efficient optimization.
+For example, LABO [pan2021emergent] embeds high-dimensional hand morphologies and controller parameters into a VAE latent space and performs BO within that space, thereby reducing dimensionality and sample complexity for grasping tasks.
+GLSO [hu2023glso] extends this concept to graph-structured robots by training a graph-VAE on grammar-generated topologies and optimizing designs using BO.
+Liu et al.
+[liu2025terrain] proposed a Grammar-VAE and integrated differential evolution with a neural decoder, allowing the generative model to be reused across tasks.
+MorphVAE [song2024morphvae] further conditions latent sampling on task embeddings and introduces a continuous natural selection strategy to improve diversity and performance.
+Generative Adversarial Networks (GANs) train a generator–discriminator pair to produce realistic morphologies by matching target data distributions.
+RoboGAN [hu2022modular] learns a one-to-many mapping from task descriptions (e.g., terrain type) to distributions over modular legged robots.
+Candidate bodies are sampled from the GAN manifold, while an evolutionary loop guides the search toward high-performing configurations, thereby enabling interpolation and diversity within the learned latent space.
+Diffusion Models have recently emerged as powerful generative priors for embodied co-design.
+DiffuseBot [wang2023diffusebot] integrates differentiable physics into the denoising process, using simulation feedback at each step to bias generation toward high-performing designs.
+Another recent work is Dynamics-Guided Diffusion Models (DGDM) [xu2024dynamics], which replaces task-specific training data with physics-based guidance, enabling generalization to new manipulation tasks and novel geometries.
+GFlowNets offer an alternative generative formulation in which morphology–controller pairs are sampled with a probability proportional to their simulated performance.
+Recent works by [nagiredla2024robonet], [nagiredlate], and [nagiredlacogent] have demonstrated that GFlowNets can generate diverse, high-quality designs while learning a structured policy over complete embodied agents.
+Generative models can capture the inter-correlations among design variables.
+Once trained, a latent model can be reused across various tasks or paired with different controllers.
+However, training generative models typically requires large datasets and substantial simulation efforts.
+Moreover, latent spaces may decode to invalid or unrealistic morphologies without appropriate constraints.
+Achieving the right balance between fidelity and diversity remains a challenge.
+IV-C3 Large-Model-Based Generation
+Large pretrained models—including Large Language Models (LLMs) [naveed2025comprehensive] and Vision–Language Models (VLMs) [zhang2024vision]—have recently been introduced into ECD, by leveraging broad knowledge and powerful generative capabilities to propose, mutate, or refine morphologies and control strategies.
+A growing body of work utilizes LLMs as intelligent mutation operators or search heuristics.
+A pioneering work is OpenELM [lehman2023evolution], which demonstrates that a LLM trained on code can act as a high-quality mutation operator in genetic programming.
+In the Sodarace experiments, the LLM consistently mutated code to generate robot walkers, and the resulting dataset was then used to train a conditional generative model that maps terrain descriptions to robot code.
+Building on this paradigm, LASeR [song2025laser] integrates a LLM into the optimization loop via a reflection mechanism that summarizes past search histories, prompting the LLM to generate improved design variants.
+RoboMoRe [fang2025robomore] further investigates LLM-driven co-design by instructing the LLM to propose candidate morphology–reward pairs jointly during a coarse search, after which promising candidates undergo alternating refinement.
+RoboMorph [qiu2024robomorph] represents each robot with a structured grammar and relies on LLM reasoning to iteratively improve morphologies through feedback.
+LLMs have also been utilized for human–AI collaborative design.
+Stella et al.
+[stella2023can] have shown that a LLM can provide high-level conceptual guidance and partial specifications for a tomato-harvesting gripper, illustrating its potential in early-stage design.
+Beyond text-based reasoning, vision models are beginning to emerge in ECD.
+VLM-Engineer [gao2025vlmgineer] combines VLMs with evolutionary search to jointly design robotic tools and associated control policies.
+This line of work suggests future workflows in which a VLM first generates sketches or 2D concept designs from a task description, which are then translated into complete 3D morphologies.
+Large models offer several advantages for ECD: they encapsulate extensive engineering and commonsense knowledge, enable creative and unconventional designs, support semantic prompting (e.g., “design a robot capable of climbing a ladder”), and diversify the search space.
+Their in-context learning capabilities make them effective interfaces between high-level human intent and concrete design proposals [makatura2023can, ma2024llm].
+However, the outputs of LLMs and VLMs are not guaranteed to be feasible, optimal, or physically consistent, highlighting the need for stronger grounding and validation mechanisms.
 Table VI: Summary of Open-Ended Co-Design approaches.
@@
 EA (GA) CPG EA (GA) L-System Locomotion Task Reward
-Brodbeck et al. [brodbeck2015morphological]
-2015
-Modular
-Robot
-Real-World
-Learnable (OL) Individual Discrete (UF) Direct Predefined (Full) Task-Driven
-EA (GA) CPG EA (GA) Vector Locomotion Task Reward
 Joachimczak et al. [joachimczak2016artificial]
@@
 CL: Closed-Loop, OL: Open-Loop, F: Fixed, UF: Unfixed, Loco.: Locomotion, Manip.: Manipulation, PID: Proportional Integral Derivative, NCA: Neural Cellular Automata
+IV-D Open-Ended Co-Design
+Open-ended co-design aims to achieve continual improvement by allowing an agent’s morphology, controller, and environment to co-evolve.
+By removing fixed terminal objectives, these systems foster sustained innovation, as agents and environments mutually generate new challenges and adaptations, as shown in Fig.
+3 (d).
+We categorize open-ended co-design into two primary subtopics: 1) Brain–Body–Environment Co-Evolution and 2) Developmental Evolution.
+IV-D1 Brain-Body-Environment Co-Evolution
+This approach treats the task environment as a co-optimized component alongside morphology and control.
+Foundational work in this direction is the Paired Open-Ended Trailblazer (POET) framework [wang2019poet, wang2020enhanced].
+POET maintains a population of environment–controller pairs, simultaneously mutating environments and evolving agents.
+Policies that succeed in one environment may be transferred to other contexts.
+Stensby et al.
+[stensby2021co] extended POET to jointly evolve morphologies and controllers in tandem with their environments.
+Their results indicate that learned environmental progressions yield greater morphological diversity than handcrafted curricula and reduce the risk of premature convergence.
+Recent work has introduced RL-driven co-evolution of the brain, body, and environment.
+Ao et al.
+proposed MECE [ao2023curriculum], which formulates a curriculum where both morphology and the environment evolve according to learned RL policies.
+This approach produces progressively challenging and structured tasks that promote sustained adaptation.
+Another emerging trend enhances environment evolution through the use of generative models.
+For instance, LLM-POET [aki2024llm] replaces manual environment mutation with a LLM that generates increasingly complex environments for a population of soft-robot agents.
+As agents improve, the LLM proposes novel tasks that escalate in difficulty, creating an open-ended loop of continual challenge and innovation.
+IV-D2 Developmental Evolution
+Developmental co-evolution draws inspiration from evolutionary developmental biology (Evo-Devo) and developmental robotics, where genotypes encode growth processes that shape both morphology and control over an agent’s lifetime [kriegman2018morphological, van2021comparing, arita2016alife, bongard2011morphological, bongard2010utility].
+Early research has investigated developmental encodings, which include L-systems [rieffel2014growing], grammar rules, and neural networks [pontes2022single, joachimczak2016artificial].
+Bongard [bongard2003evolving] demonstrated that artificial ontogeny can generate hierarchical, repeated structures and controllers that emerge as bodies grow, yielding more structured solutions.
+The other research focus is on morphological development.
+Here, developmental rules adjust parameters, such as leg angles or actuator amplitudes, during learning [naya2020experiment] or progressively “unfreeze” degrees of freedom to facilitate skill acquisition [lungarella2002adaptivity, berthouze2004motor, savastano2012incremental].
+Later work introduced mechanisms for growing new body parts during an agent’s lifetime [benureau2022morphological, jaafar2025eco, vujovic2017evolutionary], providing additional adaptive flexibility.
+For comprehensive overviews of developmental evolution, we refer readers to [naya2021morphological, wang2021survey].
+Overall, open-ended co-design is not yet a unified methodology but rather a collection of emerging ideas that leverage ontogeny, plasticity, and staged learning, and it awaits further exploration.
 V Testbeds: Applying Embodied Co-Design
-Embodied co-design has evolved rapidly due to the availability of high-fidelity simulated platforms and increasingly advanced real-world fabrication techniques.
-These resources provide standardized environments, reproducible tasks, and practical feedback loops that facilitate the automated design and validation of embodied agents [long2025survey].
+Embodied co-design has evolved rapidly due to the availability of high-fidelity simulation platforms and increasingly advanced real-world fabrication techniques.
+These resources provide standardized environments, reproducible tasks, and practical feedback loops that facilitate the automated design and validation of embodied agents.
 This section surveys representative benchmarks, datasets, and applications of ECD across both simulated and real-world domains.
@@
 V-A1 Rigid-Body Simulators
-Widely used robot simulators, including MuJoCo [todorov2012mujoco], Bullet [panerati2021learning], Unity [juliani2018unity, Veenstra2023], Gazebo [koenig2004design], Webots [michel2004cyberbotics], SAPIEN [xiang2020sapien], DART [lee2018dart], Rostok [borisov2022reconfigurable], and V-Rep [rohmer2013v], provide native support for standard robot description formats such as XML, MJCF, URDF, and SDF.
-This compatibility enables the rapid modification of morphologies and establishes these simulators as the foundation for numerous works in ECD, as summarized in Tables IVI.
-The convergence of robotics, reinforcement learning, and photorealistic rendering has further increased the demand for scalable, GPU-accelerated simulation.
-Platforms such as NVIDIA Isaac Gym and its extensions [makoviychuk2021isaac, kulkarni2023aerial, gu2024humanoid] enable massively parallel environments that support the simultaneous training of large sets of morphological variants with controllers [cheng2024structural, wang2024evolution, chen2023evolving, ringel2025text2robot, jin2025evolutionary, chen2024pretraining, du2025efficient, schaff2022n], significantly accelerating ECD pipelines.
+As summarized in Table VII, widely used robot simulators, including MuJoCo [todorov2012mujoco], Bullet [panerati2021learning], Unity [juliani2018unity, Veenstra2023], Gazebo [koenig2004design], Webots [michel2004cyberbotics], SAPIEN [xiang2020sapien], DART [lee2018dart], Rostok [borisov2022reconfigurable], and V-Rep [rohmer2013v], provide native support for standard robot description formats such as XML, MJCF, URDF, and SDF.
+This compatibility facilitates rapid modifications to morphologies and establishes these simulators as foundational tools for numerous works in ECD.
+The convergence of robotics and AI has further increased the demand for scalable, GPU-accelerated simulations.
+Platforms such as NVIDIA Isaac Gym and its extensions [makoviychuk2021isaac, kulkarni2023aerial, gu2024humanoid] enable massively parallel environments that support simultaneous training of large sets of morphological variants with controllers [cheng2024structural, wang2024evolution, chen2023evolving].
 While much of the existing ECD literature relies on non-differentiable simulators that only provide forward dynamics, a growing number of differentiable platforms supporting inverse dynamics have emerged.
-These tools—Dojo [howell2022dojo], Brax [freeman2021brax], Neuralsim [heiden2021neuralsim], Nimble [werling2021fast] (the differentiable version of DART), DiffRedMax [xu2021accelerated], and SDRS [ye2024sdrs]—facilitate gradient-based co-design.
-Based on these efforts, frameworks such as Genesis [Genesis] and MuJoCo Playground [mujoco_playground_2025] aim to unify differentiable physics, GPU acceleration, and integrated rendering within a single API, enabling end-to-end pipelines for the co-design of complex embodied agents with neural network–based controllers, as shown in Table VII.
+These tools—Dojo [howell2022dojo], Brax [freeman2021brax], Neuralsim [heiden2021neuralsim], Nimble [werling2021fast] (the differentiable version of DART), DiffRedMax [xu2021accelerated], and SDRS [ye2024sdrs]—facilitate physics-based co-design.
+Building on these efforts, frameworks such as Genesis [Genesis] and MuJoCo Playground [mujoco_playground_2025] aim to unify differentiable physics, GPU acceleration, and integrated rendering within a single API.
 V-A2 Soft-Body Simulators
-For soft-bodied agents, early platforms such as SOFA [coevoet2017software, deimel2017automated], VoxCAD [cheney2013unshackling, cheney2014electro, cheney2018scalable], and VoxCraft [kriegman2021fractals, li2024reinforcement] remain influential baselines due to their detailed physical models of deformable materials.
+For soft-bodied agents, early platforms such as SOFA [coevoet2017software, deimel2017automated], VoxCAD [cheney2013unshackling, cheney2014electro, cheney2018scalable], and VoxCraft [kriegman2021fractals, li2024reinforcement] remain influential due to their detailed physical models of deformable materials.
 Lightweight two-dimensional frameworks, such as 2D-VSR-Sim [ferigo2025totipotent], further enable rapid algorithmic prototyping.
-Recently, EvolutionGym [bhatia2021evolution] introduces a comprehensive suite of tasks along with a diverse material palette for voxel-based soft robots, which has since been widely adopted in ECD research [katayama2023ga, dongleveraging2024, harada2024lamarckian, zhao2024morphological, wang2023curriculum, wang2023preco, xie2025accelerating, zhao2025multi, liu2023rapidly, saito2024effective, zhao2025cross, davis2023subtract, tanaka2022co, mertan2024no, mertan2025controller, nadizar2025enhancing, song2025laser, song2024morphvae, liu2025morphology].
+Recently, EvolutionGym [bhatia2021evolution] introduces a comprehensive suite of tasks along with a diverse material palette for voxel-based soft robots, which has since been widely adopted in ECD research [katayama2023ga, dongleveraging2024, harada2024lamarckian, zhao2024morphological, xie2025accelerating, zhao2025multi, liu2023rapidly, saito2024effective, zhao2025cross, davis2023subtract, tanaka2022co, mertan2024no, mertan2025controller, nadizar2025enhancing, song2025laser, song2024morphvae, liu2025morphology].
 Complementary to voxel-based systems, several platforms integrate differentiable soft-body physics, including DiffTaichi [hu2019difftaichi], ChainQueen [hu2019chainqueen], and DiffAqua [ma2021diffaqua].
-SoftZoo [wang2023softzoo] extends this direction by providing a differentiable benchmark across various terrains (e.g., flat ground, snow, and water) and task objectives (e.g., speed, turning, and path following).
-In parallel, Sorotoki [caasenbrood2024sorotoki] provides an integrated MATLAB toolkit for co-designing soft robots.
+SoftZoo [wang2023softzoo] advances this direction by providing a benchmark across various terrains (e.g., flat ground, snow, and water) and task objectives (e.g., speed, turning, and path following).
+In parallel, Sorotoki [caasenbrood2024sorotoki] provides an integrated MATLAB toolkit.
 V-A3 Embodied Tasks Simulators
 Notably, recent specialized simulators for embodied intelligence, such as iGibson [li2022igibson], AI2-THOR [kolve2017ai2], Habitat-Sim [szot2021habitat], and LIBERO [liu2023libero], provide realistic and interactive 3D environments that integrate rich visual, physical, and semantic information.
-These platforms support the training and evaluation of increasingly complex embodied agents, including those with advanced control paradigms (e.g., Vision-Language-Action models) [zitkovich2023rt, kim2024openvla, bjorck2025gr00t] and various morphologies (e.g., humanoids).
-Furthermore, they provide challenging benchmarks for long-horizon tasks, which makes them well-suited foundations for advancing ECD research in the future.
-V-B Datasets
-In addition to the aforementioned simulation platforms, several recent ECD datasets have been developed to support the systematic analysis of co-design across diverse embodied agents.
-RoboCrafter-QA [chen2025large] provides comparative data for evaluating whether LLMs can link high-level task descriptions to low-level morphological and material choices in soft robotics.
-RoboDesign1M [le2025robodesign1m] offers a large-scale dataset of one million robotic design samples across multiple domains, enabling tasks such as design–image generation, design-centric visual question answering, and design–image retrieval.
+These platforms support the training and evaluation of increasingly complex embodied agents, including those with advanced control paradigms (e.g., Vision-Language-Action models [zitkovich2023rt, kim2024openvla, bjorck2025gr00t]) and various morphologies (e.g., humanoids).
+Furthermore, they provide challenging benchmarks for long-horizon tasks involving navigation, manipulation, and multi-step planning, which makes them well-suited foundations for advancing ECD research in the future.
 Table VII: A Summary of the Past Decade’s Simulators and Physics Engines Supporting Embodied Co-Design.
@@
 SDRS 2024 Soft Yes [link]
+V-B Datasets
+In addition to the aforementioned simulation platforms, several ECD datasets have been developed to support the systematic analysis of agent design.
+RoboCrafter-QA [chen2025large] provides comparative data to evaluate whether LLMs can effectively link high-level task descriptions to low-level morphological and material choices in soft robotics.
+RoboDesign1M [le2025robodesign1m] offers a large-scale dataset of one million robotic design samples across multiple domains, enabling tasks such as design–image generation, design-centric visual question answering, and design–image retrieval.
+Fig.
+4: Co-designed embodied agents in both virtual and real environments.
+Example images numbered from (a) to (p) are adapted from [sims1994competition, hornby2001evolving, shim2003generating, Veenstra2023, ma2021diffaqua, ha2019reinforcement, luck2020data, zhao2020robogrammar, gupta2021embodied, xu2021accelerated, cheney2013unshackling, zardini2021seeking, medvet2021biodiversity, wang2023curriculum, wang2023softzoo, yue2025toward], respectively.
+Example images numbered from (q) to (y) are adapted from [nygaard2021environmental, revolve2, Strgar-RSS-24, xie2024map, schaff2022soft, liu_voxcraft_2020, kobayashi2024computational, matthews2023efficient, cheng2024structural], respectively.
 V-C Real-World Applications
 While simulation facilitates rapid exploration, physical validation remains the ultimate test of ECD.
-Recent advances in fabrication, multi-material 3D printing, and embedded sensing have made it feasible to realize co-designed morphologies.
-In this section, we summarize recent real-world applications.
+Recent advances in fabrication, 3D printing, and embedded sensing have made it feasible to realize co-designed morphologies.
+In this section, we summarize recent real-world applications (Fig.
+4).
 V-C1 Rigid Robots
@@
 We also review notable applications in both simulated and physical environments.
-Although current results indicate that co-designed agents outperform fixed-body baselines, many open questions remain unanswered.
+Although the current results indicate that co-designed agents outperform fixed-body baselines, many questions remain unanswered.
 We highlight these challenges and discuss promising directions for future research.
-We hope that this survey serves as both a foundational reference for newcomers and a comprehensive resource for experts seeking deeper insights into ECD.
+We hope this survey serves as both a foundational reference for newcomers and a comprehensive resource for experts seeking deeper insights into ECD.
 [Uncaptioned image] Yuxing Wang (Member, IEEE) received the B.Eng.
@@
 degree in control science and engineering at THU.
-Before that, he was a research assistant with THU and a research intern with Tencent AI Lab.
+Before that, he worked as a research assistant at THU and as a research intern at Tencent AI Lab.
 His research interests include evolutionary learning and embodied intelligence, with a particular focus on evolutionary reinforcement learning for control optimization, and embodied co-design for modular robotic systems.
+[Uncaptioned image] Zhiyu Chen received the B.Sc.
+degree in electronic engineering from Tsinghua University, Beijing, China, in 2025.
+He is currently pursuing a M.Eng.
+degree in artificial intelligence at Tsinghua University.
+His research interests are in reinforcement learning theory, embodied intelligence, and large language models.
+[Uncaptioned image] Tiantian Zhang received the B.Sc.
+degree in automation from the Department of Information Science and Technology, Central South University, Changsha, China, in 2015, and the M.Sc.
+and Ph.D.
+degrees in control science and engineering from the Department of Automation, Tsinghua University, Beijing, China, in 2018 and 2024, respectively.
+She is currently a Post-Doctoral Researcher with the Center for Artificial Intelligence and Robotics, Shenzhen International Graduate School, Tsinghua University.
+Her research interests include data science, intelligent decision-making, and reinforcement learning.
+She has been serving as reviewers for Nature Communications, TNNLS, TMLR and AAAI.
+[Uncaptioned image] Qiyue Yin received the Ph.D.
+degree from the National Laboratory of Pattern Recognition, Institute of Automation, Chinese Academy of Sciences (CASIA), Beijing, China, in 2017.
+He is currently an Associate Professor at CASIA.
+His major research interests include machine learning, pattern recognition and artificial intelligence on games.
+[Uncaptioned image] Yongzhe Chang received the Ph.D.
+degree in machine learning&data analysis from University of New South Wales, Australia, in 2020.
+He was a post-doctoral research fellow with the AI&Robot laboratory, Tsinghua University, from 2021 to 2023.
+Then he joined Tsinghua University as a associate researcher since 2023.
+His research interests include machine learning, reinforcement learning, large model and intelligent robot, etc.
+[Uncaptioned image] Zhiheng Li (Member, IEEE) received the Ph.D.
+degree in control science and engineering from Tsinghua University, Beijing, China, in 2009.
+He is currently an Associate Professor with the Department of Automation, Tsinghua University, and with the Graduate School at Shenzhen, Tsinghua University, Shenzhen, China.
+His research interests include traffic operation, advanced traffic management system, urban traffic planning, and intelligent transportation systems.
+Dr.
+Li is an Associate Editor for IEEE TRANSACTIONS ON INTELLIGENT TRANSPORTATION SYSTEMS.
+[Uncaptioned image] Liang Wang (Fellow, IEEE) received the B.Eng.
+and M.Eng.
+degrees from Anhui University in 1997 and 2000, respectively, and the Ph.D.
+degree from the Institute of Automation, Chinese Academy of Sciences (CASIA) in 2004.
+From 2004 to 2010, he was a research assistant with Imperial College London, United Kingdom, and Monash University, Australia, a research fellow with the University of Melbourne, Australia, and a lecturer with the University of Bath, United Kingdom, respectively.
+He is currently a full professor of the Hundred Talents Program with the National Lab of Pattern Recognition, CASIA.
+He has widely authored or coauthored in highly ranked international journals, such as TPAMI and TIP, and leading international conferences, such as CVPR, ICCV, and ICDM.
+His research interests include machine learning, pattern recognition, and computer vision.
+He is an IAPR fellow.
+[Uncaptioned image] Xueqian Wang (Member, IEEE) received the M.Sc.
+and Ph.D.
+degrees in control science and engineering from the Harbin Institute of Technology (HIT), Harbin, China, in 2005 and 2010, respectively.
+From June 2010 to February 2014, he was a Post-Doctoral Researcher with HIT.
+From March 2014 to November 2019, he was an Associate Professor with the Division of Informatics, Shenzhen International Graduate School, Tsinghua University, Shenzhen, China.
+He is currently a Professor and the Leader of the Center for Artificial Intelligence and Robotics, Shenzhen International Graduate School, Tsinghua University.
+His research interests include robot dynamics and control, teleoperation, intelligent decision-making and game playing, and fault diagnosis.